### v1

# E-values, Multiple Testing and Beyond

Abstract
We discover a connection between the Benjamini-Hochberg (BH) procedure and the recently proposed e-BH procedure Wang and Ramdas (2022) with a suitably defined set of e-values. This insight extends to a generalized version of the BH procedure and the model-free multiple testing procedure in Barber and Candès (2015) (BC) with a general form of rejection rules. The connection provides an effective way of developing new multiple testing procedures by aggregating or assembling e-values resulting from the BH and BC procedures and their use in different subsets of the data. In particular, we propose new multiple testing methodologies in three applications, including a hybrid approach that integrates the BH and BC procedures, a multiple testing procedure aimed at ensuring a new notion of fairness by controlling both the group-wise and overall false discovery rates (FDR), and a structure adaptive multiple testing procedure that can incorporate external covariate information to boost detection power. One notable feature of the proposed methods is that we use a data-dependent approach for assigning weights to e-values, significantly enhancing the efficiency of the resulting e-BH procedure. The construction of the weights is non-trivial and is motivated by the leave-one-out analysis for the BH and BC procedures. In theory, we prove that the proposed e-BH procedures with data-dependent weights in the three applications ensure finite sample FDR control. Furthermore, we demonstrate the efficiency of the proposed methods through numerical studies in the three applications.
Keywords: Benjamini-Hochberg procedure, Cross-fitting, E-values, False discovery rate, Leave-one-out analysis, Multiple testing
1 Introduction
When working with high-dimensional data in modern scientific fields, the problem of multiple testing often arises when we explore a vast number of hypotheses with the goal of detecting signals while also controlling some error measures, such as the false discovery rate (FDR). The Benjamini-Hochberg (BH) procedure (Benjamini and Hochberg, 1995) is perhaps the most widely used FDR-controlling procedure that rejects a hypothesis whenever its p-value is less than or equal to an adaptive rejection threshold determined by the whole set of p-values. Barber and Candès (2015) proposed a model-free FDR-controlling (BC) procedure that estimates the number of false rejections by leveraging the symmetry of p-values under the null and compares each p-value with an adaptive threshold. More recently, there is a growing literature on utilizing e-values for statistical inference under different contexts, see, e.g., (Grünwald et al., 2020; Shafer, 2021; Vovk and Wang, 2021; Xu et al., 2021; Ignatiadis et al., 2022; Dunn et al., 2023; Xu and Ramdas, 2023). In particular, Wang and Ramdas (2022) proposed a multiple testing procedure (named e-BH procedure) by applying the BH procedure to e-values, which was shown to control the FDR even when the e-values exhibit arbitrary dependence.
In this work, we establish a connection between the BH and e-BH procedures with a suitably defined set of e-values, proving that they yield identical rejection sets. This connection extends to a generalized version of the BH and BC procedures, which can have a more general form for the rejection rule. Based on these findings, we propose new multiple testing procedures by aggregating e-values from the BH and BC procedures or assembling e-values from different subsets of the data. Specifically, we consider three concrete applications, which we will illustrate in detail below.
Our first application concerns the development of a robust and efficient multiple testing procedure by leveraging the strength of the BH and BC procedures. The BH and BC procedures employ different strategies to estimate the number of false rejections, leading to distinct performances depending on the underlying signal density and strength. Empirical results in the literature suggest that neither one dominates the other. For instance, the BH procedure is usually more powerful when the underlying signal is sparse, while the BC procedure excels in the case of dense signal (Arias-Castro and Chen, 2017). It is often impossible to know in advance which method will perform better in real-world applications. Reporting results from the better-performing method can lead to an inflated FDR and is considered as a type of data snooping. To develop a more efficient procedure with guaranteed FDR control, we propose a hybrid method that utilizes e-values as a bridge to integrate the testing results from the BH and BC procedures. Specifically, we construct a set of e-values by properly weighting the e-values from both the BH and BC procedures. We then use the constructed e-values as the input for the e-BH procedure. We show that the resulting hybrid procedure controls the FDR and can significantly improve the worse-performing method in finite sample.
Our second application is concerned with a setting where one has to make (high-stakes) decisions by testing $n$ hypotheses that are partitioned into $G$ groups according to some (protected) attributes. Let us consider the example of a bank’s loan approval process, where customers are grouped by gender or race. The null hypothesis in this case is that the loan should be approved for the customer. The bank faces the challenge of controlling the overall FDR, which means that the bank should not falsely decline too many customers’ applications, as it would reduce the loan volume and decrease income. Additionally, the bank also needs to control the group-wise FDR to ensure that certain racial or gender groups are not unfairly denied loans. If the group-wise FDR inflates, customers who believe they are unfairly denied loans may take their business elsewhere. Therefore, the bank aims to maintain control over the FDR for each group while also controlling the overall FDR across all hypotheses. Applying an FDR-controlling procedure to the whole set of hypotheses, ignoring the group structure will fail to control the FDR for individual groups. On the other hand, performing multiple testing for each group separately at level $\alpha$ does not ensure that the overall FDR is controlled at the same level $\alpha$. A naive procedure to achieve error control at both the group and overall levels is to conduct multiple testing for each group separately at the reduced level $\alpha/G$, which turns out to be too conservative, as seen in our (unreported) numerical studies. To address this challenge, we propose a fair multiple testing procedure that uses e-values as a bridge to assemble the results from different groups, controlling the FDR within each group and the overall FDR simultaneously. Specifically, we implement the BC procedure for each group, followed by constructing e-values by assembling e-values from different groups with proper weights, and using the constructed e-values as the input for the e-BH procedure. We show that the resulting procedure simultaneously controls the FDR within each group and the overall FDR in finite sample.
As the final application, we consider the problem of multiple testing with external structural information in the form of covariates, which has received much recent attention, as leveraging auxiliary information can enhance the power and interpretability of multiple testing results in many scientific applications. For instance, in differential expression analysis of RNA-seq data, which tests for differences in the mean expression of the genes between conditions, the sum of read counts per gene across all samples could be the auxiliary data since it is informative of the statistical power. In differential abundance analysis of microbiome sequencing data, which tests for differences in the mean abundance of the detected bacterial species between conditions, the genetic divergence among species is important auxiliary information since closely related species usually have similar physical characteristics and tend to covary with the condition of interest. A growing list of works has reflected the importance of this research direction in recent years, for instance, (Genovese et al., 2006; Hu et al., 2010; Sun et al., 2015; Ignatiadis et al., 2016; Boca and Leek, 2018; Lei and Fithian, 2018; Li and Barber, 2019; Ignatiadis and Huber, 2021; Yun et al., 2022; Zhang and Chen, 2022; Zhao and Zhou, 2023). However, these existing works suffer from different limitations. For example, the local FDR-based method (Sun et al., 2015; Cao et al., 2022) lacks finite sample FDR control and can only guarantee FDR control asymptotically. The weighted BH methods (Ignatiadis and Huber, 2021; Li and Barber, 2019) lead to suboptimal power, as observed in our numerical studies. To address these drawbacks, we propose a powerful multiple testing procedure that incorporates auxiliary information with guaranteed FDR control in finite samples. We randomly split samples into several groups and use the cross-fitting approach to estimate the rejection function in each group using all samples in other groups. Then, we implement the generalized BC procedure for each group. We assemble the BC procedure induced e-values from different groups with proper weights. The assembled e-values are used as the input for e-BH procedure. We show that the proposed procedure controls the FDR at the desired level in finite sample.
Our approach involves a data-dependent method for weighting e-values when aggregating them from the BH and BC procedures (the first application) or assembling them from different subsets of the data (the second and third applications). It is important to note that our weighting method differs from the “boosting factor” proposed by Wang and Ramdas (2022), which involves multiplying each e-value by a factor to boost them up before applying the e-BH procedure. In our approach, we construct weights that ensure the weighted e-values satisfying Condition (3) below. This condition is sufficient to maintain FDR control of the e-BH procedure at the desired level. Our numerical findings show that implementing the e-BH procedure with data-dependent weights improves its efficiency across all applications compared to the unweighted e-BH procedure.
The rest of the paper is organized as follows. In Section 2, we provide a brief review of the BH, BC, and e-BH procedures and establish the equivalence between the BH/BC and the e-BH procedures. In Section 3, we extend this equivalence to a generalized version of the BH and BC procedures. We present new multiple testing methodologies in three different applications, including a hybrid approach that integrates the BH and BC procedures, a multiple testing procedure that controls both the group-wise and overall FDRs, and a structure adaptive multiple testing procedure, in Sections 4.1, 4.2, and 4.3, respectively. Section 5 concludes. The appendix contains additional numerical results and all proofs of the main results.
2 Preliminaries
To begin with, we provide a brief overview of the BH, BC, and e-BH procedures, and establish the equivalence between the BH (BC) procedure and the e-BH procedure with a suitably defined set of e-values. This equivalence appears to be a new finding that has not been explicitly stated in the previous literature.
We are interested in testing $n$ hypotheses $\{H_{1},\dots,H_{n}\}$ simultaneously. Let $\bm{\theta}=\{\theta_{1},\dots,\theta_{n}\}\in\{0,1\}^{n}$ indicate the underlying truth of each hypothesis, where $\theta_{i}=0$ if $H_{i}$ is under the null and $\theta_{i}=1$ otherwise. Denote by $\bm{\delta}=(\delta_{1},\dots,\delta_{n})\in\{0,1\}^{n}$ a decision rule for the $n$ hypotheses, where we reject (accept) the $i$th hypothesis if $\delta_{i}=1$ ($\delta_{i}=0$). The FDR for the decision rule $\bm{\delta}$ is defined as the expectation of the false discovery proportion (FDP), i.e.,
$\mathrm{FDR}(\bm{\delta})=\mathbb{E}[\mathrm{FDP}(\bm{\delta})],\quad\mathrm{FDP}(\bm{\delta})=\frac{\sum_{i=1}^{n}(1-\theta_{i})\delta_{i}}{1\vee\sum_{i=1}^{n}\delta_{i}},$
where $a\vee b=\max\{a,b\}$. The goal of an FDR controlling procedure is to ensure that the FDR is bounded from above by a pre-specified number $\alpha\in(0,1)$.
2.1 BH procedure
The Benjamini-Hochberg (BH) procedure (Benjamini and Hochberg, 1995) is perhaps the most widely used FDR-controlling method. To describe the procedure, suppose we observe a p-value $p_{i}$ for each $H_{i}$. Sort the p-values in ascending order as $p_{(1)}\leq\dots\leq p_{(n)}$ and let $\hat{k}=\max\left\{i\colon p_{(i)}\leq(\alpha i)/n\right\}$. The BH procedure rejects all hypotheses $H_{(i)}$ with $i\leq\hat{k}$, where $H_{(i)}$ is the hypothesis associated with $p_{(i)}$. This procedure is equivalent to rejecting all $H_{i}$ with $p_{i}\leq T_{\mathrm{BH}}$, where $T_{\mathrm{BH}}$ is defined as
$T_{\mathrm{BH}}=\sup\left\{0<t\leq 1\colon\frac{nt}{1\vee R(t)}\leq\alpha\right\},$ (1)
with $R(t)=\sum_{i=1}^{n}\mathbbm{1}\{p_{i}\leq t\}$ being the number of rejections given the threshold $t$, and $\mathbbm{1}\{A\}$ denoting the indicator function associated with a set $A$. It is well known that if the null p-values are mutually independent and super-uniform, and are independent of the alternative p-values, the BH procedure at level $\alpha$ controls the FDR at the level $\alpha n_{0}/n\leq\alpha$, where $n_{0}$ is the number of hypotheses under the null.
2.2 BC procedure
In a seminal paper by (Barber and Candès, 2015), the authors proposed a model-free multiple testing procedure (BC procedure hereafter) that exploits the symmetry of the null p-values (or test statistics) to estimate the number of false rejections. More precisely, the BC procedure specifies a data-dependent threshold, denoted by $T_{\mathrm{BC}}$, which is determined as follows:
$T_{\mathrm{BC}}=\sup\left\{0<t<0.5\colon\frac{1+\sum^{n}_{i=1}\mathbbm{1}\{p_{i}\geq 1-t\}}{1\vee R(t)}\leq\alpha\right\},$ (2)
and it rejects all $H_{i}$ with $p_{i}\leq T_{\mathrm{BC}}$. The BC procedure has been shown to provide finite sample FDR control under suitable assumptions (Barber and Candès, 2015).
2.3 E-values and e-BH procedure
A non-negative random variable $E$ is called an e-value if $\mathbb{E}[E]\leq 1$ under the null hypothesis. Suppose we observe $n$ e-values $e_{1},\dots,e_{n}$ corresponding to the hypotheses $H_{1},\dots,H_{n}$. The $\alpha$-level e-BH procedure involves sorting the e-values in decreasing order as $e_{(1)}\geq\dots\geq e_{(n)}$ and rejecting the hypotheses associated with the $\hat{k}$ largest e-values, where $\hat{k}\coloneqq\max\left\{1\leq i\leq n\colon e_{(i)}\geq n/(i\alpha)\right\}$.
Proposition 1 (Theorem 2 of Wang and Ramdas (2022)).
Suppose the e-values satisfy
$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[e_{i}]\leq n,$ (3)
where $\mathcal{H}_{0}=\{1\leq i\leq n:\theta_{i}=0\}$. Then, the $\alpha$-level e-BH procedure controls the FDR at the level $\alpha$, regardless of the dependence structure among the e-values.
2.4 Connection between the BH and e-BH procedures
To see the connection between the BH and e-BH procedures, we define the e-value associated with $H_{i}$ to be
$e_{i}=T_{\mathrm{BH}}^{-1}\mathbbm{1}\{p_{i}\leq T_{\mathrm{BH}}\},$ (4)
where $T_{\mathrm{BH}}$ is given in (1). By Lemmas 3-4 in Storey et al. (2004), $\sum_{i\in\mathcal{H}_{0}}\mathbbm{1}\{p_{i}\leq t\}/t$ for $0<t\leq 1$ is martingale with time running backwards with respect to the filtration $\mathcal{F}_{t}=\sigma(\mathbbm{1}\{p_{i}\leq s\}:t\leq s\leq 1,i=1,2,\dots,n)$, and $T_{\mathrm{BH}}$ is a stopping time with respect to $\mathcal{F}_{t}$, where $\mathcal{F}_{t}$ is the sigma field generated by $\mathbbm{1}\{p_{i}\leq s\}$ for $t\leq s\leq 1.$ By the optional stopping time theorem, we have
$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[e_{i}]=\mathbb{E}\left[\frac{\sum_{i\in\mathcal{H}_{0}}\mathbbm{1}\{p_{i}\leq T_{\mathrm{BH}}\}}{T_{\mathrm{BH}}}\right]=\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}}\mathbbm{1}\{p_{i}\leq 1\}\right]=n_{0},$
which implies that the e-values defined by equation (4) satisfy (3). Thus by Proposition 1, the corresponding e-BH procedure controls the FDR at the desired level. Moreover, we claim that the e-BH procedure based on the e-values defined in (4) is equivalent to the BH procedure based on $\{p_{1},\dots,p_{n}\}$ in the sense that they produce the same set of rejections.
Theorem 1.
Let $\mathcal{S}_{\mathrm{BH}}$ be the set of rejections obtained through the BH procedure at the FDR level $\alpha$, and let $\mathcal{S}_{\mathrm{eBH}}$ represent the set of rejections obtained from the e-BH procedure at the same FDR level $\alpha$, with the e-values defined in (4). Then we have $\mathcal{S}_{\mathrm{BH}}=\mathcal{S}_{\mathrm{eBH}}$.
2.5 Connection between the BC and e-BH procedures
As noted in a recent work by Ren and Barber (2022), the BC procedure is equivalent to the e-BH procedure based on the following e-values:
$e_{i}=\frac{n\mathbbm{1}\{p_{i}\leq T_{\mathrm{BC}}\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC}}\}},$
where $T_{\mathrm{BC}}$ is the threshold defined in (2).
3 Generalized BH and BC procedures
In this section, we extend the BH and BC procedures to allow a general class of rejection rules. This generalization enables the testing procedure to utilize cross-sectional information among the p-values and external structural information for each hypothesis, which often results in a higher multiple testing power. We shall establish the connection between these generalized procedures and the corresponding e-BH procedures.
3.1 Generalized BH procedure
We generalize the BH procedure to allow the rejection rule to take the form of $\varphi_{i}(p_{i})\leq t$, where $\varphi_{i}$ is a strictly increasing function and can differ for each $i$. Let us define $F_{i}(t)=\varphi_{i}^{-1}(t)$ as the inverse function of $\varphi_{i}$, and consider the rejection threshold
$T_{\text{GBH}}=\sup\left\{0<t\leq 1:\frac{\sum^{n}_{i=1}F_{i}(t)}{1\vee R(t)}\leq\alpha\right\},$ (5)
where $R(t)=\sum^{n}_{i=1}\mathbbm{1}\{\varphi_{i}(p_{i})\leq t\}$. The generalized BH (GBH) procedure rejects $H_{i}$ whenever $\varphi_{i}(p_{i})\leq T_{\mathrm{GBH}}$. Similar to the BH procedure, the GBH procedure can be equivalently implemented in the following way. Define $\bar{F}(t)=n^{-1}\sum^{n}_{i=1}F_{i}(t)$ which is strictly increasing in $t$. We sort $q_{i}=\varphi_{i}(p_{i})$ in an ascending order, i.e., $q_{(1)}\leq\dots\leq q_{(n)}$ and find the largest $k$, represented as $\hat{k}$, for which $q_{(k)}\leq\bar{F}^{-1}\left(\frac{\alpha k}{n}\right).$ We then reject $H_{(i)}$ for all $i\leq\hat{k}$. The following proposition states that, under appropriate conditions in the form of $F_{i}$, the GBH procedure ensures FDR control.
Proposition 2.
Suppose that the null p-values $\{p_{i}\}_{i\in\mathcal{H}_{0}}$ are mutually independent and super-uniform, and are independent of the alternative p-values $\{p_{i}\}_{i\notin\mathcal{H}_{0}}$. If $F_{i}(t)=c_{i}h(t)$ where $c_{i}$ is some constant and $h$ is a strictly increasing function of $t$, the GBH procedure controls the FDR at level $\alpha$.
As illustrated in the following examples, the GBH procedure aligns with several methods in the literature of structure adaptive multiple testing.
Example 1.
Let $\varphi_{i}(p)=p/\omega_{i}$, where $\omega_{i}$ denotes the weight for the $i$th hypothesis with $\omega_{i}>0$ and $\sum_{i=1}^{n}\omega_{i}=n$. The GBH procedure associated with this choice of $\varphi_{i}$ corresponds to the weighted BH procedure first introduced by Genovese et al. (2006). In this case, $F_{i}(t)=w_{i}t$ and the rejection threshold can be expressed as
$T_{\text{GBH}}=\sup\left\{0<t\leq 1\colon\frac{\sum_{i=1}^{n}w_{i}t}{1\vee R(t)}\leq\alpha\right\}=\sup\left\{0<t\leq 1\colon\frac{nt}{1\vee R(t)}\leq\alpha\right\},$
where $R(t)=\sum_{i=1}^{n}\mathbbm{1}\{p_{i}/w_{i}\leq t\}$.
Example 2.
Another choice of $\varphi_{i}$ is the local FDR (Efron, 2005; Sun and Cai, 2007) under a two-group mixture model. Specifically, we assume that the p-values are independently generated from the mixture density given by $\pi_{i}+(1-\pi_{i})(1-\kappa)p^{-\kappa}$, where $\pi_{i}\in(0,1)$ is the mixing proportion and $\kappa\in(0,1)$ controls the shape of the p-value density under the alternative. The rejection rule based on the local FDR under this model is given by
$\varphi_{i}(p)=\frac{\pi_{i}}{\pi_{i}+(1-\pi_{i})(1-\kappa)p^{-\kappa}}\leq t,$ (6)
which is the probability that the $i$th hypothesis is under the null given the observed p-value being $p$. It follows that
$F_{i}(t)=\varphi_{i}^{-1}(t)=\left(\frac{\pi_{i}(1-t)}{(1-\pi_{i})t(1-\kappa)}\right)^{-1/\kappa}=c_{i}h(t),$
where $c_{i}=\left\{\pi_{i}/\bigl{(}(1-\pi_{i})(1-\kappa)\bigr{)}\right\}^{-1/\kappa}$ and $h(t)=\left\{(1-t)/t\right\}^{-1/\kappa}$.
3.2 Connection between the GBH and e-BH procedures
Analogous to the BH procedure, we show that the GBH procedure is equivalent to the e-BH procedure applied to the following e-values:
$e_{i}=\frac{n\mathbbm{1}\{\varphi_{i}(p_{i})\leq T_{\text{GBH}}\}}{\sum^{n}_{i=1}F_{i}(T_{\mathrm{GBH}})},$ (7)
where $T_{\text{GBH}}$ is defined in (5). By the leave-one-out argument, we prove the following result.
Proposition 3.
Under the assumptions in Proposition 2, the e-values $e_{1},\dots,e_{n}$ defined in (7) satisfy Condition (3). As a consequence, the corresponding e-BH procedure controls the FDR at the desired level.
Additionally, we can prove that the e-BH procedure and the GBH procedure deliver the same set of rejections.
Theorem 2.
Let $\mathcal{S}_{\mathrm{GBH}}$ be the set of rejections obtained through the GBH procedure at the FDR level $\alpha$, and let $\mathcal{S}_{\mathrm{eBH}}$ represent the set of rejections obtained from the e-BH procedure at the same FDR level $\alpha$, with the e-values defined in (7). Then we have $\mathcal{S}_{\mathrm{GBH}}=\mathcal{S}_{\mathrm{eBH}}$.
3.3 Generalized BC procedure
In this section, we generalize the BC procedure with the rejection rule given by $\varphi_{i}(p_{i})\leq t$. We assume that the null p-value satisfies the following condition:
$P(p_{i}\leq a)\leq P(p_{i}\geq 1-a)=P(1-p_{i}\leq a),\quad\text{ for all }0\leq a\leq 0.5.$ (8)
Condition (8) is weaker than the mirror conservativeness in Lei and Fithian (2018), and it can be shown that super-uniformity implies (8). Indeed,
$P(1-p_{i}\leq a)\geq 1-P(p_{i}\leq 1-a)\geq 1-(1-a)=a\geq P(p_{i}\leq a).$
Assume that $\varphi_{i}$ is an increasing and continuous function, and define $F_{i}(x)=\sup\{0\leq p\leq 1\colon\varphi_{i}(p)\leq x\}$. We claim that
$P(\varphi_{i}(p_{i})\leq b)=P(p_{i}\leq F_{i}(b)).$
To see this, consider two cases. If $\varphi_{i}(p_{i})\leq b$, by the definition of $F_{i}(b)$, we have $p_{i}\leq F_{i}(b)$. On the other hand, if $p_{i}\leq F_{i}(b)$, then $\varphi_{i}(p_{i})\leq\varphi_{i}(F_{i}(b))=\lim_{p\uparrow F_{i}(b)}\varphi_{i}(p)\leq b$, where we use the fact that $\varphi_{i}$ is increasing to get the two inequalities, and the equality is due to the continuity of $\varphi_{i}$. Therefore, the above claim together with equation (8) implies that
$P(\varphi_{i}(p_{i})\leq b)=P(p_{i}\leq F_{i}(b))\leq P(1-p_{i}\leq F_{i}(b))=P(\varphi_{i}(1-p_{i})\leq b),$ (9)
for all $\varphi_{i}(0)\leq b\leq\varphi_{i}(0.5)$. Hence, we have
$\frac{\sum_{i\in\mathcal{H}_{0}}\mathbbm{1}\{\varphi_{i}(p_{i})\leq t\}}{1\vee\sum_{i=1}^{n}\mathbbm{1}\{\varphi_{i}(p_{i})\leq t\}}$ $\approx\frac{\sum_{i\in\mathcal{H}_{0}}P(\varphi_{i}(p_{i})\leq t\})}{1\vee\sum_{i=1}^{n}\mathbbm{1}\{\varphi_{i}(p_{i})\leq t\}}$
$\leq\frac{\sum_{i\in\mathcal{H}_{0}}P(\varphi_{i}(1-p_{i})\leq t\})}{1\vee\sum_{i=1}^{n}\mathbbm{1}\{\varphi_{i}(p_{i})\leq t\}}$
$\approx\frac{\sum_{i\in\mathcal{H}_{0}}\mathbbm{1}\{\varphi_{i}(1-p_{i})\leq t\}}{1\vee\sum_{i=1}^{n}\mathbbm{1}\{\varphi_{i}(p_{i})\leq t\}}$
$\leq\frac{1+\sum_{i=1}^{n}\mathbbm{1}\{\varphi_{i}(1-p_{i})\leq t\}}{1\vee\sum_{i=1}^{n}\mathbbm{1}\{\varphi_{i}(p_{i})\leq t\}},$
where the last term can be viewed as a conservative estimate of the FDP. Motivated by this observation, we define the threshold for the GBC procedure as
$T_{\text{GBC}}=\sup\left\{0<t\leq T_{\mathrm{up}}\colon\frac{1+\sum_{i=1}^{n}\mathbbm{1}\{\varphi_{i}(1-p_{i})\leq t\}}{1\vee\sum_{i=1}^{n}\mathbbm{1}\{\varphi_{i}(p_{i})\leq t\}}\leq\alpha\right\},$ (10)
which is the largest cutoff such that the FDP estimate is bounded above by $\alpha$, where $T_{\mathrm{up}}$ satisfies $T_{\mathrm{up}}<\min_{i}\varphi_{i}(0.5)$. The GBC procedure rejects $H_{i}$ whenever $\varphi_{i}(p_{i})\leq T_{\text{GBC}}$.
Proposition 4.
Suppose that the null p-values $\{p_{i}\}_{i\in\mathcal{H}_{0}}$ are mutually independent and satisfy Condition (8), and are independent of the alternative p-values $\{p_{i}\}_{i\notin\mathcal{H}_{0}}$. Assuming $\varphi_{i}$ is a monotonic increasing and continuous function for all $i$, then the GBC procedure ensures the FDR control at level $\alpha$.
Compared to the GBH procedure, the GBC approach affords us greater flexibility in selecting $\varphi_{i}$, as it no longer requires $\varphi_{i}$ to be strictly increasing and its (generalized) inverse function does not have to fulfill the condition in Proposition 2.
Example 3.
Suppose the p-value $p_{i}$ is generated independently from the two-group mixture model: $\pi_{i}f_{0}+(1-\pi_{i})f_{1,i}$, where $\pi_{i}\in(0,1)$ is the mixing proportion and $f_{0}$ and $f_{1,i}$ denote the p-value distributions under the null and alternative respectively. The local FDR is defined as
$\mathrm{Lfdr}_{i}(p)=\frac{\pi_{i}f_{0}(p)}{\pi_{i}f_{0}(p)+(1-\pi_{i})f_{1,i}(p)},$
which is the posterior probability that the $i$th hypothesis is under the null given the observed p-value being $p$. The monotone likelihood ratio assumption (Sun and Cai, 2007) states that $f_{1,i}(p)/f_{0}(p)$ is decreasing in $p$. Under this assumption, $\varphi_{i}(p)=\mathrm{Lfdr}_{i}(p)$ is monotonically increasing in $p$ and thus fulfills the requirement in Proposition 4. Additionally, it has been shown in the literature that the rejection rule $\varphi_{i}(p_{i})=\mathrm{Lfdr}_{i}(p_{i})\leq t$ is optimal in the sense of maximizing the expected number of true positives among the decision rules that control the marginal FDR at level $\alpha$, see e.g., Sun and Cai (2007); Lei and Fithian (2018); Cao et al. (2022).
3.4 Connection between the GBC and e-BH procedures
We show that the GBC procedure is equivalent to the e-BH procedure with the following choice of e-values:
$e_{i}=\frac{n\mathbbm{1}\{\varphi_{i}(p_{i})\leq T_{\text{GBC}}\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{\text{GBC}}\}},$ (11)
where $T_{\text{GBC}}$ is defined in (10). By equation (B.1) in the proof of Proposition 4, we have
$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[e_{i}]\leq n,$
which implies that the corresponding e-BH procedure controls the FDR at the desired level. Furthermore, the following theorem shows that the e-BH procedure with the e-values defined above is equivalent to the GBC procedure.
Theorem 3.
Let $\mathcal{S}_{\mathrm{GBC}}$ be the set of rejections obtained through the GBC procedure at the FDR level $\alpha$, and let $\mathcal{S}_{\mathrm{eBH}}$ represent the set of rejections obtained from the e-BH procedure at the same FDR level $\alpha$, with the e-values defined in (11). Then we have $\mathcal{S}_{\mathrm{GBC}}=\mathcal{S}_{\mathrm{eBH}}$.
4 Applications
Building upon the insights from the previous section, we shall develop new multiple testing procedures by aggregating the e-values resulting from different procedures (e.g., the BH and BC procedures) or the same procedure but applied to different subsets of the data. By showing that the aggregated e-values fulfill Condition (3), we have the e-BH procedure based on the aggregated e-values controls the FDR in finite sample. We illustrate this idea through three concrete applications below. The first application develops a hybrid procedure to leverage the strengths from the BH and BC procedures, aiming to improve their robustness and efficiency. The second application investigates fairness in the multiple testing context. In particular, given a sensitive/protected variable that partitions the data into several groups, we develop a fair multiple testing procedure that ensures the FDR within each group and the overall FDR across all groups are controlled at the same desired level simultaneously. Our last application proposes a new structure adaptive multiple testing procedure through cross-fitting and e-value aggregation. We demonstrate that the proposed procedure has guaranteed FDR control and can deliver comparable or even higher power than existing competitors.
4.1 A hybrid procedure
Empirical results in the literature suggest that neither the BH procedure nor the BC procedure dominates the other. Specifically, in the case of sparse signals (i.e., $n_{0}/n\approx 1$), the BH procedure typically has higher power than the BC procedure (Arias-Castro and Chen, 2017). In contrast, when the signal is dense, the BH procedure becomes conservative, making the BC procedure more preferable. However, in real-world applications, it is often impossible to know the signal strength and density in advance, and thus it is unclear which method would perform better. Applying both methods and reporting the results associated with the one that delivers more rejections does not ensure FDR control. Given the insights from Sections 2.4-2.5, we ask the question of whether it is possible to develop a new multiple testing procedure that has finite sample FDR control and is powerful across a wider range of signals by borrowing the strengths from both the BH and BC procedures. To address this question, we introduce a hybrid approach that utilizes e-values as a bridge to combine the BH and BC procedures, aiming for better robustness and efficiency. Let $e_{\mathrm{BH},i}$ and $e_{\mathrm{BC},i}$ be the e-values from the BH and BC procedures (at levels $\alpha_{\mathrm{BH}}$ and $\alpha_{\mathrm{BC}}$) respectively for testing the $i$th hypothesis. We define the weighted e-values
$e_{i}=w_{\mathrm{BH},i}e_{\mathrm{BH},i}+w_{\mathrm{BC},i}e_{\mathrm{BC},i},$
that aggregates the information from both the BH and BC procedures, and apply the e-BH procedure to the resulting e-values, where $w_{\mathrm{BH},i}$ and $w_{\mathrm{BC},i}$ are non-negative weights. The detailed implementation of the hybrid procedure is described in Algorithm 1 below.
Algorithm 1 Hybrid procedure
1:p-values $p_{1},\dots,p_{n}$ and significance levels $\alpha_{\mathrm{BH}}$, $\alpha_{\mathrm{BC}}$ and $\alpha_{\mathrm{eBH}}$
2:Implement the BH procedure at the significance level $\alpha_{\mathrm{BH}}$. Compute the threshold $T_{\mathrm{BH}}$ using (1). Determine the e-value for the BH procedure with the formula:
$e_{\mathrm{BH},i}=\frac{1}{T_{\mathrm{BH}}}\mathbbm{1}\{p_{i}\leq T_{\mathrm{BH}}\}.$
3:Implement the BC procedure at the significance level $\alpha_{\mathrm{BC}}$. Compute the threshold $T_{\mathrm{BC}}$ using (2). Determine the e-value for the BC procedure with the formula:
$e_{\mathrm{BC},i}=\frac{n\mathbbm{1}\{p_{i}\leq T_{\mathrm{BC}}\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC}}\}}.$
4:Calculate the weighted averaged e-value by
$e_{i}=w_{\mathrm{BH},i}e_{\mathrm{BH},i}+w_{\mathrm{BC},i}e_{\mathrm{BC},i}.$ (12)
5:Run the e-BH procedure utilizing the weighted averaged e-values at the significance level $\alpha_{\mathrm{eBH}}$.
6:The indices of rejected hypotheses.
4.1.1 Choice of significance levels and weights
We now discuss the choices of the significance levels $(\alpha_{\mathrm{BH}},\alpha_{\mathrm{BC}})$ and the weights $(w_{\mathrm{BH},i},w_{\mathrm{BC},i})$ in Algorithm 1, which play important roles in the hybrid procedure. Different from the target FDR level $\alpha_{\mathrm{eBH}}$, the choice of $(\alpha_{\mathrm{BH}},\alpha_{\mathrm{BC}})$ does not affect the FDR control level but instead affect the power of the hybrid procedure. Following the discussions in Section 3.2 of Ren and Barber (2022), when there are $n_{a}$ non-nulls with extremely strong signals, we expect that
$1+\sum_{j=1}^{n}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC}}\}\approx\sum_{j\in\mathcal{H}_{0}}\mathbbm{1}\{p_{j}\leq T_{\mathrm{BC}}\}\approx\frac{\alpha_{\mathrm{BC}}n_{a}}{1-\alpha_{\mathrm{BC}}}.$
In a similar spirit, we expect the FDR of the BH procedure to be $\tau_{0}\alpha$, where $\tau_{0}=n_{0}/n.$ Let $R_{\mathrm{BH}}$ be the number of rejections in the BH procedure. Then we have $R_{\mathrm{BH}}\approx n_{a}/(1-\tau_{0}\alpha_{\mathrm{BH}})$ and
$T_{\mathrm{BH}}=\frac{\alpha_{\mathrm{BH}}R_{\mathrm{BH}}}{n}\approx\frac{\alpha_{\mathrm{BH}}n_{a}}{n(1-\tau_{0}\alpha_{\mathrm{BH}})}.$
Thus the hybrid procedure will reject $H_{i}$ with $i\notin\mathcal{H}_{0}$ when
$e_{i}\approx\frac{\omega_{\mathrm{BC},i}n(1-\alpha_{\mathrm{BC}})}{\alpha_{\mathrm{BC}}n_{a}}+\frac{\omega_{\mathrm{BH},i}n(1-\tau_{0}\alpha_{\mathrm{BH}})}{\alpha_{\mathrm{BH}}n_{a}}\geq\frac{n}{\alpha_{\mathrm{eBH}}n_{a}},$
which is equivalent to
$\frac{\omega_{\mathrm{BC},i}(1-\alpha_{\mathrm{BC}})}{\alpha_{\mathrm{BC}}}+\frac{\omega_{\mathrm{BH},i}(1-\tau_{0}\alpha_{\mathrm{BH}})}{\alpha_{\mathrm{BH}}}\geq\frac{1}{\alpha_{\mathrm{eBH}}}.$ (13)
We found that setting $\alpha_{\mathrm{BC}}=\alpha_{\mathrm{BH}}=\alpha_{\mathrm{eBH}}/(1+\alpha_{\mathrm{eBH}})$ fulfills the above constraint, and leads to good performance in our numerical studies.
Next, we discuss the selection of $(w_{\mathrm{BH},i},w_{\mathrm{BC},i})$ which balances the contribution from the two methods. A natural choice would be to set $w_{\mathrm{BH},i}=w_{\mathrm{BC},i}=0.5$ for all $i$. The theorem below shows that if the weights are independent of the p-values and satisfy $w_{\mathrm{BH},i}+\max_{j}w_{\mathrm{BC},j}\leq 1$ for all $i$, then the e-values defined in (12) satisfy Condition (3). Therefore, according to Proposition 1, the corresponding e-BH procedure controls the FDR at the desired level.
Theorem 4.
Suppose that the null p-values $\{p_{i}\}_{i\in\mathcal{H}_{0}}$ are mutually independent and super-uniform, and are independent of the alternative p-values $\{p_{i}\}_{i\notin\mathcal{H}_{0}}$. Assume that $w_{\mathrm{BH},i}$ and $w_{\mathrm{BC},i}$ are independent of the p-values, and $w_{\mathrm{BH},i}+\max_{j}w_{\mathrm{BC},j}\leq 1$ for all $i$. Then, the weighted e-values given in (12) satisfy Condition (3).
[…]
$\begin{split}w_{\mathrm{BH},i}&=\frac{T_{\mathrm{BH},i}}{T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j,i}\}\right)},\\ w_{\mathrm{BC},i}&=\frac{\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC}}\}\right)}{\max_{j}T_{\mathrm{BH},j}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC}}\}\right)}.\end{split}$ (14)
Theorem 5.
Suppose that the null p-values $\{p_{i}\}_{i\in\mathcal{H}_{0}}$ are mutually independent and super-uniform, and are independent of the alternative p-values $\{p_{i}\}_{i\notin\mathcal{H}_{0}}$. Let $w_{\mathrm{BH},i}$ and $w_{\mathrm{BC},i}$ be defined in (14). Then, the weighted average e-values specified in (12) with the data-dependent weights given in (14) satisfy Condition (3).
As a consequence of Theorem 5, the hybrid procedure with the data-dependent weights in (14) provides finite sample FDR control. Furthermore, in view of the proof of Theorem 5, the above conclusion remains true if we replace $T_{\mathrm{BH},i}$ in (14) by any function of $\tilde{p}_{-i}$.
4.1.2 Numerical studies
We investigate the finite sample performance of the hybrid procedure via several simulation examples. We fix the sample size $n=1,000$ and set the significance level at $\alpha_{\mathrm{eBH}}=0.05$. For each experimental setting, the average FDP (which estimates the FDR) and the average power based on 500 Monte Carlo replications are reported. We consider two different ways to implement the hybrid procedure: (1) eBH_Ada, for which the weights are calculated via (14) and $\alpha_{\mathrm{BH}}=\alpha_{\mathrm{BC}}=\alpha_{\mathrm{eBH}}/(1+\alpha_{\mathrm{eBH}})$; (2) eBH_Ave, for which the weights are set as $w_{\mathrm{BH},i}=w_{\mathrm{BC},i}=0.5$ for all $i=1,\cdots,n$ and $\alpha_{\mathrm{BH}}=\alpha_{\mathrm{BC}}=\alpha_{\mathrm{eBH}}/2$. Notice that the weights defined in (14) involve the term $T_{\mathrm{BC},j,i}$, which can be computationally expensive. To reduce the computational burden, we also consider a fast implementation of eBH_Ada, referred to as fast_eBH_Ada, which uses the weights
$w_{\mathrm{BH},i}=\frac{T_{\mathrm{BH},i}}{T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j}\}\right)},$
but otherwise is the same as eBH_Ada.
We consider two settings. In Setting S1, we set $n_{a}$ to be $100$ and $200$, corresponding to $10\%$ and $20\%$ signals. For the null hypotheses, the p-values are generated independently from the uniform distribution on $[0,1]$; for the alternative hypotheses, the p-values are generated independently from the uniform distribution on $[0,p_{1}]$, where $p_{1}$ takes values between $1.25\times 10^{-2}$ and $1.5\times 10^{-2}$. Here, a larger $p_{1}$ indicates a weaker signal strength on average. Figure 1 summarizes the results for Setting S1. All methods under consideration control the FDR at the 5% level. For the dense signal case (i.e., 20% of the hypotheses are non-nulls), the BH procedure is quite conservative, while the BC procedure delivers significantly higher power. The performance of eBH_Ada is between the BH and BC procedures. In contrast, eBH_Ave does not provide much improvement over the BH procedure. It is also worth noting that fast_eBH_Ada delivers almost identical results compared to eBH_Ada but with a much lower computational cost. When $n_{a}=100$, the eBH_Ada procedure has nearly the same power as the BC procedure, and it outperforms eBH_Ave whose power is almost identical to the BH procedure.
Figure 1: FDR and power for Setting S1. The left panel and the right panel correspond to the cases of $n_{a}=100$ and $n_{a}=200$, respectively. The results are based on $500$ Monte Carlo replications.
In Setting S2. we set $n_{a}$ to be $50$ and 100, corresponding to $5\%$ and $10\%$ of all hypotheses. Under the null hypothesis, we generate the test statistics $X_{i}$ from the standard normal distribution $\mathcal{N}(0,1)$. Under the alternatives, $X_{i}$ follows $\mathcal{N}(\mu\log(n),1)$, where the parameter $\mu$ controls the signal strength, which varies between $0.3$ and $0.5$. Subsequently, the p-values are computed as $p_{i}=1-\Phi(X_{i})$, where $\Phi$ denotes the distribution function of the standard normal distribution. The simulation results are presented in Figure 2. When the signal density is 5% (i.e., $n_{a}=50$), the BH procedure performs better than the BC procedure. eBH_Ada has nearly the same power as the BH procedure, and it again outperforms eBH_Ave, which slightly improves over the BC procedure. When the signal density increases to 10% (i.e., $n_{a}=100$), the performance gap between the BH and BC procedures diminishes. eBH_Ada and its fast version produce almost the same power as the BH procedure, outperforming both the eBH_Ave and BC procedures. In this case, the performance of eBH_Ave is worse than that of the BH and BC procedures.
Figure 2: FDR and power for Setting S2. The left panel and the right panel correspond to the cases of $n_{a}=50$ and $n_{a}=200$, respectively. Results are based on $500$ Monte Carlo replications.
In summary, eBH_Ada is capable of improving the worse-performing method within the BH and BC procedures under different scenarios. Interestingly, its power can be nearly identical to the better-performing method, which demonstrates the adaptivity of eBH_Ada. fast_eBH_Ada delivers almost the same results as eBH_Ada in all settings with a much lower computational burden. We therefore recommend the use of fast_eBH_Ada in practice.
4.2 Fair multiple testing
In recent years, there has been a growing interest in developing machine learning algorithms that ensure certain notion of fairness for groups categorized by protected attributes, such as gender, race, or ethnicity. However, the development of statistical inference procedures that guarantee specific notions of fairness has not received the same level of attention. To address this gap, we introduce a new notion of fairness in the context of multiple hypothesis testing, and propose a procedure to ensure this notion of fairness.
Suppose we have $n$ hypotheses which can be partitioned into $G$ distinct groups, denoted as $\{\mathcal{G}_{g}\colon g=1,\dots,G\}$, according to some protected attributes, where $\cup_{g=1}^{G}\mathcal{G}_{g}=[n]:=\{1,2,\dots,n\}$ and $\mathcal{G}_{g}\cap\mathcal{G}_{g^{\prime}}=\emptyset$ for $g\neq g^{\prime}$. Recall that $\theta_{i}\in\{0,1\}$ represents the underlying truth of $H_{i}$ and $\delta_{i}\in\{0,1\}$ denotes the decision rule for $H_{i}$. We define the group-wise FDP and FDR based on $\bm{\delta}$ respectively as:
$\mathrm{FDP}_{g}(\bm{\delta})=\frac{\sum_{i\in\mathcal{G}_{g}}(1-\theta_{i})\delta_{i}}{1\vee\sum_{i\in\mathcal{G}_{g}}\delta_{i}},\quad\text{FDR}_{g}(\bm{\delta})=\mathbb{E}[\text{FDP}_{g}(\bm{\delta})],\quad g=1,2,\dots,G.$
Definition 1.
A decision rule $\bm{\delta}$ with the target FDR level $\alpha$ is called a fair decision rule if it controls both the group-wise FDRs uniformly over $1\leq g\leq G$ and the overall FDR at the level $\alpha$, i.e., $\max_{1\leq g\leq G}\text{FDR}_{g}(\bm{\delta})\leq\alpha$ and $\text{FDR}(\bm{\delta})\leq\alpha$.
Remark 1.
Among various notions of fairness that have been discussed in the literature (Verma and Rubin, 2018), predictive parity in the classification context (Chouldechova, 2017) is most relevant to our definition of fairness. To elaborate, let us consider a binary classification problem where $Y\in\{0,1\}$ represents the true labels and $\widehat{Y}\in\{0,1\}$ denotes the predicted labels. In this scenario, the FDR is defined as $P(Y=0|\widehat{Y}=1)$. A classifier satisfies predictive parity if it ensures an equal FDR across all groups. However, the multiple testing problem differs from the classification problem as the underlying truth of each hypothesis is unobserved and cannot be used to construct the decision rule. In our context, it is more appropriate to control the FDRs across different groups at the same desired level rather than forcing them to be the same.
[…]
Algorithm A 0.039 0.006 0.067
Algorithm B 0.057 0.048 0.025
Table 1: FDR results for Setting E2, where the nominal FDR level is $5\%$.
To overcome the issues above, we consider a multiple testing procedure by assembling the e-values from the BC procedure applied to each group separately. Specifically, we implement the BC procedure at the level $\alpha$ for each individual group and let
$T_{g}=\sup\left\{0<t<0.5\colon\frac{1+\sum_{i\in\mathcal{G}_{g}}\mathbbm{1}\{p_{i}\geq 1-t\}}{1\vee\sum_{i\in\mathcal{G}_{g}}\mathbbm{1}\{p_{i}\leq t\}}\leq\alpha\right\}$ (15)
be the rejection threshold for the $g$th group with $1\leq g\leq G$. Define
$e_{i}=\frac{n_{g}w_{i}\mathbbm{1}\{p_{i}\leq T_{g}\}}{1+\sum_{j\in\mathcal{G}_{g}}\mathbbm{1}\{p_{j}\geq 1-T_{g}\}},$ (16)
for $i\in\mathcal{G}_{g}$, where $w_{i}>0$ represents the weight for the $i$th hypothesis, which will be specified in Section 4.2.1. After collecting the e-values from each group, we implement the e-BH procedure at level $\alpha$. The testing procedure is summarized in Algorithm 2.
Algorithm 2 Fair multiple testing
p-values $p_{1},\dots,p_{n}$, group indices $\mathcal{G}_{1},\dots,\mathcal{G}_{G}$, significance level $\alpha$
for $g=1,\dots,G$ do
     Implement the BC procedure utilizing the p-values $\{p_{i}\colon i\in\mathcal{G}_{g}\}$ at the level $\alpha$.
     Calculate the threshold $T_{g}$ using (15).
     for $i\in\mathcal{G}_{g}$ do
         Calculate the e-value $e_{i}$ using (16).
     end for
end for
Assemble the e-values from all groups.
Run the e-BH procedure utilizing the assembled e-values at the level $\alpha$.
The indices of rejected hypotheses.
It is important to note that only the nonzero e-values can be rejected in the e-BH procedure. As a result, the group-wise FDR is effectively controlled at level $\alpha$ for each group. In the following section, we will demonstrate that the e-BH procedure can effectively control the overall FDR even when the weights are selected in a data-dependent manner.
4.2.1 Choice of weights and FDR control
Controlling the overall FDR requires ensuring that the e-values defined in (16) satisfy Condition (3). One strategy is to set $w_{i}=1$ for all $i$. Another strategy takes into account the group size by setting $w_{i}=n/(Gn_{g})$ for all $i\in\mathcal{G}_{g}$ and $g=1,\dots,G$, where $n_{g}=|\mathcal{G}_{g}|$ is the cardinality of the $g$th group. It can be verified that Condition (3) is satisfied by both strategies. The e-BH procedures based on these two weight choices are referred to as eBH_1 and eBH_2, respectively.
Theorem 6.
Suppose that the null p-values $\{p_{i}\}_{i\in\mathcal{H}_{0}}$ are mutually independent and satisfy Condition (8), and are independent of the alternative p-values $\{p_{i}\}_{i\notin\mathcal{H}_{0}}$. Further suppose the weights $w_{i}$ are independent of the p-values and satisfy
$\sum_{g=1}^{G}n_{g}\max_{i\in\mathcal{G}_{g}}w_{i}\leq n.$ (17)
Then, the e-values specified in (16) satisfy Condition (3). Hence, the corresponding e-BH procedure controls the FDR in finite sample.
Our simulations show that eBH_1 and eBH_2 frequently suffer from low statistical power. To enhance efficiency, we suggest utilizing a data-dependent weight selection approach inspired by the leave-one-out technique (Barber et al., 2020).
Denote the p-values in the $g$th group by $\mathbf{p}_{g}=\{p_{i}\}_{i\in\mathcal{G}_{g}}$. Write $\tilde{p}_{i}=\min\{p_{i},1-p_{i}\}$, and let $\mathbf{p}_{g,i}$ for $i\in\mathcal{G}_{g}$ be the collection of p-values obtained by replacing $p_{i}$ with $\tilde{p}_{i}$ in $\mathbf{p}_{g}$. By viewing $T_{g}$ as a functions of $\mathbf{p}_{g}$, we define $T_{g,i}=T_{g}(\mathbf{p}_{g,i})$, i.e., the threshold of the BC procedure applied to the set of p-values $\mathbf{p}_{g,i}$. We define the data-dependent weights as
$w_{i}=\frac{\frac{n}{n_{g}}\left(1+\sum_{j\neq i,j\in\mathcal{G}_{g}}\mathbbm{1}\{p_{j}\geq 1-T_{g}\}\right)}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{g}}\mathbbm{1}\{p_{j}\geq 1-T_{g}\}\right)+\sum_{g^{\prime}\neq g}\sum_{j\in\mathcal{G}_{g^{\prime}}}\mathbbm{1}\{p_{j}\geq 1-T_{g^{\prime},j}\}},$ (18)
for $i\in\mathcal{G}_{g}$. When $G=2$, we have
$w_{i}=\frac{\frac{n}{n_{1}}\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1}\}\right)}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1}\}\right)+\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{p_{j}\geq 1-T_{2,j}\}}$
for $i\in\mathcal{G}_{1}$ and
$w_{i}=\frac{\frac{n}{n_{2}}\left(1+\sum_{j\neq i,j\in\mathcal{G}_{2}}\mathbbm{1}\{p_{j}\geq 1-T_{2}\}\right)}{\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1,j}\}+\left(1+\sum_{j\neq i,j\in\mathcal{G}_{2}}\mathbbm{1}\{p_{j}\geq 1-T_{2}\}\right)}$
for $i\in\mathcal{G}_{2}$. The e-BH procedure based on the weights specified in (18) is referred to as the eBH_Ada method in the following discussions. The theorem below states that eBH_Ada has finite sample FDR control.
Theorem 7.
Suppose that the null p-values $\{p_{i}\}_{i\in\mathcal{H}_{0}}$ are mutually independent and satisfy Condition (8), and are independent of the alternative p-values $\{p_{i}\}_{i\notin\mathcal{H}_{0}}$. Then, the e-values specified in (16) with the weights defined via (18) satisfy Condition (3). Hence, the corresponding e-BH procedure controls the FDR in finite sample.
4.2.2 Numerical studies
We shall compare the finite sample performance of the proposed method with two naive approaches through simulations. The first method disregards the group information and directly applies the BC procedure to all p-values. We refer to this method as BC_Com for future reference. BC_Com has two shortcomings. Firstly, it may fail to control the group-wise FDRs, as illustrated in Setting E2. Secondly, it fails to ensure comparable power across different groups, resulting in the possibility of one group having high power while the other group has nearly zero power, as illustrated in Setting E1. The same issue is also encountered by eBH_1. An alternative approach involves implementing the BC procedure for each group separately and combining all rejections. We call this method BC_Sep. Although BC_Sep effectively controls the FDR for individual groups, it does not guarantee the overall FDR control.
Remark 2.
We also tried BC_Sep at the level $\alpha/G$, a naive method that controls both the group-wise and overall FDRs. Indeed, let $\hat{n}_{ga}$ be the number of rejections for the $g$th group, and denote the number of false rejections for the $g$th group by $\hat{n}_{g0}$. Then we have $\mathbb{E}[\hat{n}_{g0}/(1\vee\hat{n}_{ga})]\leq\alpha/G$, which implies that
$\mathbb{E}\left[\frac{\sum_{g=1}^{G}\hat{n}_{g0}}{1\vee\sum_{g=1}^{G}\hat{n}_{ga}}\right]\leq\mathbb{E}\left[\sum_{g=1}^{G}\frac{\hat{n}_{g0}}{1\vee\hat{n}_{ga}}\right]\leq\alpha.$
However, this method has nearly zero power in all our simulation settings. Therefore, we have decided not to include its results in the tables below.
[…]
In all settings, we assume that the p-values follow the uniform distribution on $[0,1]$ under the null. For the first group, the p-value is assumed to follow $\mathrm{Unif}[0,u_{1}]$ under the alternatives, while for the second group, it follows $\mathrm{Unif}[0,u_{2}]$ under the alternatives. The parameter values for different settings are detailed in Table 2.
Setting $n_{1}$ $n_{1a}$ $u_{1}$ $n_{2}$ $n_{2a}$ $u_{2}$
E1 $100$ $20$ $1.5\times 10^{-2}$ $1000$ $20$ $1.5\times 10^{-3}$
E2 $100$ $20$ $1.5\times 10^{-3}$ $1000$ $20$ $1.5\times 10^{-3}$
Table 2: Parameter settings for the case of $G=2$. Here, $n_{g}$ represents the number of hypotheses for the $g$th group; $n_{ga}$ denotes the number of non-null hypotheses in the $g$th group. The parameter $u_{g}$ is the upper bound of the uniform distribution for the p-values under the alternatives for the $g$th group.
Setting E1 corresponds to a scenario in which, for instance, the first group consists of ethnic minorities while the second group comprises of majorities. The number of non-nulls are the same across the two groups. The alternative p-values in the first group are larger than those in the second group on average as $u_{1}>u_{2}$. Table 3 summarizes the results for this setting. BC_Com exhibits high power for the second group, yet its power in the first group are rather low. This is because the non-null p-values from the first group are not sufficiently small, and a combined analysis of the two groups demands a lower threshold and thus fails to reject them. Additionally, BC_Sep has a slightly inflated overall FDR in this case.
[…]
BC_Com 0.144 0.041 0.248 0.023 0.012 0.025
BC_Sep 0.296 0.352 0.24 0.053 0.036 0.025
eBH_1 0.07 0 0.141 0.02 0 0.02
eBH_2 0.078 0.077 0.078 0.008 0.008 0.008
eBH_Ada 0.181 0.212 0.149 0.028 0.021 0.015
Table 3: FDR and power for Setting E1, where the nominal FDR level is 5%.
Setting E2 is the same as Setting E1 except that $u_{1}=u_{2}=1.5\times 10^{-2}$. The results for Setting E2 are presented in Table 4. We observe that BC_Com fails to control the FDR for the second group, which can be explained as follows. Due to the fact that the non-null p-values have a similar scale for both groups and the sample size of the first group is significantly smaller than that of the second group, BC_Com has a higher threshold compared to the BC procedure applied only to the second group. This can result in an FDR inflation in the second group for BC_Com. We also observe that BC_Sep suffers from an overall FDR inflation. In contrast, all three variants of the e-BH procedure control the group-wise and overall FDRs at the desired level. eBH_Ada has a much higher power than the other two e-value based methods.
[…]
BC_Com 0.605 0.605 0.606 0.039 0.006 0.067
BC_Sep 0.573 0.906 0.24 0.057 0.048 0.025
eBH_1 0.07 0 0.141 0.02 0 0.02
eBH_2 0.216 0.217 0.215 0.017 0.011 0.022
eBH_Ada 0.384 0.552 0.216 0.036 0.029 0.023
Table 4: FDR and power for Setting E2, where the nominal FDR level is 5%.
We present the results for $G=4$ in Appendix C.1, where we consider three different scenarios (Settings F1-F3). In particular, we note that BC_Com suffers from a severe FDR inflation with the empirical FDR being 0.312 at the 5% target level in Setting F2. In Setting F3, BC_Sep has an empirical overall FDR being 0.346, which is much higher than the 20% level.
To summarize, as seen in Settings E2 and F2, BC_Com has no guarantee in controlling the group-wise FDR. On the other hand, BC_Sep fails to control the overall FDR, as observed in all the settings, particularly Settings E2 and F3 in Appendix C.1. The e-BH based approaches provide both group-wise and overall FDR control, making them fair according to Definition 1. However, eBH_1 and eBH_2 may suffer from power loss under certain scenarios. In contrast, eBH_Ada demonstrates consistent effectiveness across all settings by achieving FDR fairness and reasonable power.
4.3 Structure adaptive multiple testing
Having access to various types of auxiliary information that reflect the structural relationships among hypotheses is becoming increasingly common. Taking advantage of such auxiliary information can improve the statistical power in multiple testing. In this section, we consider the scenarios where, in addition to the p-value $p_{i}$, there is associated structural information in the form of a covariate $x_{i}$ for each hypothesis. This side information represents heterogeneity among the p-values and may affect the prior probabilities of the null hypotheses being true or the signal strength under alternatives. Our goal is to develop a multiple testing procedure that can incorporate such external structural information to improve statistical power and has guaranteed FDR control in finite sample. The high-level idea behind our approach is to relax the p-value thresholds for hypotheses that are more likely to be non-null and tighten the thresholds for the other hypotheses through the use of hypothesis-specific rejection rule, i.e., $\varphi_{i}(p_{i})\leq t$, so that the FDR can be controlled.
Our proposed method combines the cross-fitting technique (Ignatiadis and Huber, 2021) (a sample-splitting and fitting approach that enables learning the hypothesis-specific rejection function $\varphi_{i}$ without overfitting as long as the hypotheses can be partitioned into independent folds) with the GBC procedure introduced in Section 3. First, we randomly split the data into $G$ distinct groups, denoted as $\{\mathcal{G}_{g}\colon g=1,\dots,G\}$, where $\bigcup_{g=1}^{G}\mathcal{G}_{g}=[n]$ and $\mathcal{G}_{g}\cap\mathcal{G}_{g^{\prime}}=\emptyset$ for $g\neq g^{\prime}$. We estimate the rejection function $\varphi_{i}$ for the hypothesis $i$ in the $g$th group using the data from all other groups, which ensures that the estimated rejection function is independent of the p-values in group $g$. We then apply the GBC procedure based on the estimated rejection functions separately to each group and obtain the corresponding e-values. Finally, we collect all the e-values and apply the e-BH procedure at the target level to control the FDR.
To describe the cross-fitting procedure, let us assume that $\varphi_{i}(p)=\varphi(p,x_{i};\beta)$ for some unknown parameter $\beta$ that needs to be estimated from the data. We define the cross-fitted estimate as
$\hat{\beta}_{-g}=\operatorname*{arg\,min}_{\beta}\sum_{i\notin\mathcal{G}_{g}}\mathcal{L}(p_{i},x_{i},\beta).$
Here $\mathcal{L}$ is some loss function such as negative log-likelihood under the two-group mixture model in Example 3, in which case we assume $\pi_{i}(p)=\pi(p,x_{i};\beta_{1})$, $f_{1,i}(p)=f(p,x_{i};\beta_{2})$ and $\beta=(\beta_{1},\beta_{2}).$ It should be noted that the choice of the functional form of $\varphi_{i}$ is highly flexible and does not impact the FDR control due to cross-fitting that avoids over-fitting. Given $\hat{\beta}_{-g}$, we define
$\hat{\varphi}_{i}(p)=\varphi(p,x_{i};\hat{\beta}_{-g})\quad\text{for}\quad i\in\mathcal{G}_{g}.$
Next, we apply the GBC procedure using the cross-fitted functions $\{\hat{\varphi}_{i}(\cdot)\}_{i\in\mathcal{G}_{g}}$ at the level $\alpha_{\mathrm{GBC}}$. The corresponding threshold for the $g$th group is given by
$T_{g}=\sup\left\{0<t\leq T_{g,\mathrm{up}}\colon\frac{1+\sum_{i\in\mathcal{G}_{g}}\mathbbm{1}\{\hat{\varphi}_{i}(1-p_{i})\leq t\}}{1\vee\sum_{i\in\mathcal{G}_{g}}\mathbbm{1}\{\hat{\varphi}_{i}(p_{i})\leq t\}}\leq\alpha_{\mathrm{GBC}}\right\},$ (19)
where $T_{g,\mathrm{up}}$ satisfies $T_{g,\mathrm{up}}<\min_{i\in\mathcal{G}_{g}}\hat{\varphi}_{i}(0.5)$. We define the e-value for all $i\in\mathcal{G}_{g}$ as
$e_{i}=\frac{n_{g}w_{i}\mathbbm{1}\{\hat{\varphi}_{i}(p_{i})\leq T_{g}\}}{1+\sum_{j\in\mathcal{G}_{g}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{g}\}},$ (20)
where $w_{i}>0$ represents the e-value weight for hypothesis $i$ in group $G_{g}$. Finally, we aggregate all e-values from each group and implement the e-BH procedure. A detailed description of our procedure is given in Algorithm 3.
Algorithm 3 Cross-fitting based structure adaptive multiple testing
p-values $p_{1},\dots,p_{n}$, covariates $x_{1},\dots,x_{n}$, group indices $\mathcal{G}_{1},\dots,\mathcal{G}_{G}$, significance levels $\alpha_{\mathrm{GBC}}$ and $\alpha_{\mathrm{eBH}}$
for $g=1,\dots,G$ do
     Compute the estimate of the rejection rule $\hat{\varphi}_{i}(\cdot)$ for $i\in\mathcal{G}_{g}$ using $\mathbf{p}_{-g}$ and $\mathbf{x}_{-g}$.
     Calculate the threshold $T_{g}$ using (19).
     for $i\in\mathcal{G}_{g}$ do
         Calculate the e-value $e_{i}$ using (20).
     end for
end for
Assemble the e-values from all groups.
Run the e-BH procedure utilizing the assembled e-values at the level $\alpha_{\mathrm{eBH}}$.
The indices of rejected hypotheses.
4.3.1 Choice of weights and FDR control
To ensure that the FDR is controlled at the desired level, it is crucial to verify that the e-values defined in equation (20) satisfy Condition (3). We shall show that under certain conditions on the weights, the e-values defined by (20) satisfy (3), and as a result, the corresponding e-BH procedure controls the FDR at the desired level. Before stating the main theorem, let us first introduce some notation. Define $\tilde{p}_{i}=\min\{p_{i},1-p_{i}\}$, $\mathbf{p}_{g}=\{p_{i}\}_{i\in\mathcal{G}_{g}}$, and $\mathbf{p}_{g,i}$ as the collection of p-values obtained by replacing $p_{i}$ with $\tilde{p}_{i}$ in $\mathbf{p}_{g}$ for $i\in\mathcal{G}_{g}$. Also, let $\mathbf{p}_{-g}=\{p_{i}\}_{i=1}^{n}\setminus\mathbf{p}_{g}$. Due to cross-fitting, the estimated function $\hat{\varphi}_{i}(\cdot)$ for $i\in\mathcal{G}_{g}$ only depends on $\mathbf{p}_{-g}$. Moreover, given the fitted functions $\hat{\varphi}_{i}$ for $i\in\mathcal{G}_{g}$, the threshold $T_{g}$ defined in (19) can be treated as a function of $\mathbf{p}_{g}$. Let $T_{g,j}=T_{g}(\mathbf{p}_{g,j};\{\hat{\varphi}_{l}\}_{l\in\mathcal{G}_{g}})$. We impose the following assumptions.
Assumption 1.
Let $\{(p_{i},x_{i})\}$ for $1\leq i\leq n$ be the p-value and covariate pairs.
  1. (A)
    The null pairs $\{(p_{i},x_{i})\}_{i\in\mathcal{H}_{0}}$ are mutually independent.
  2. (B)
    The null pairs $\{(p_{i},x_{i})\}_{i\in\mathcal{H}_{0}}$ are independent of the alternative pairs $\{(p_{i},x_{i})\}_{i\notin\mathcal{H}_{0}}$.
  3. (C)
    For $i\in\mathcal{H}_{0}$, $p_{i}$ is independent of $x_{i}$ and satisfies Condition (8).
Assumption 2.
For all $1\leq i\leq n$, $\varphi_{i}(\cdot,x_{i};\beta)$ is a monotonic increasing and continuous function given any $\beta$ and $x_{i}$.
Assumption 1 concerns the dependence of the null pairs, which is standard in the literature; see, e.g., Assumption 1 of Ignatiadis and Huber (2021) and Zhao and Zhou (2023). Assumption 2 implies that $P(\varphi_{i}(p_{i},x_{i};\beta)\leq b)\leq P(\varphi_{i}(1-p_{i},x_{i};\beta)\leq b)$ for all $\varphi_{i}(0,x_{i};\beta)\leq b\leq\varphi_{i}(0.5,x_{i};\beta)$, which will be used in our proof. We describe a concrete choice of $\varphi_{i}$ in Section 4.3.2 below.
Theorem 8.
Suppose Assumptions 1 and 2 hold. If the weights $\{w_{i}\}$ are independent of the p-values and covariate information and satisfy (17), then the e-values defined in (20) fulfills Condition (3).
A naive choice is to set $w_{i}=1$ for all $i=1,2,\cdots,n$, which satisfy (17). However, this choice of weights often leads to low statistical power in simulations. To improve efficiency, we propose a data-dependent approach for constructing the weights. Given the group index $g$, for $g^{\prime}\neq g$, $i\in\mathcal{G}_{g}$ and $j\in\mathcal{G}_{g^{\prime}}$, let $\hat{\varphi}_{j}^{g,i,p}$ be the cross-fitted function obtained by replacing $\mathbf{p}_{g}$ with $\mathbf{p}_{g,i}(p)$, where $\mathbf{p}_{g,i}(p)$ is the collection of p-values with $p_{i}$ replaced by $p$. Define $T_{g^{\prime},j}^{g,i,p}=T_{g^{\prime}}(\mathbf{p}_{g^{\prime},j};\{\hat{\varphi}_{l}^{g,i,p}\}_{l\in\mathcal{G}_{g^{\prime}}})$. For $i\in\mathcal{G}_{g}$ with $1\leq g\leq G$, we propose the following e-value weights:
$w_{i}=\frac{\frac{n}{n_{g}}\left(1+\sum_{j\neq i,j\in\mathcal{G}_{g}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{g}\}\right)}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{g}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{g}\}\right)+\sup_{p\in[0,1]}\sum_{g^{\prime}\neq g}\sum_{j\in\mathcal{G}_{g^{\prime}}}\mathbbm{1}\{\hat{\varphi}_{j}^{g,i,p}(1-p_{j})\leq T_{g^{\prime},j}^{g,i,p}\}}.$ (21)
In the case of $G=2$, we have
$w_{i}=\frac{\frac{n}{n_{1}}\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1}\}\right)}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1}\}\right)+\sup_{p\in[0,1]}\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{\hat{\varphi}_{j}^{1,i,p}(1-p_{j})\leq T_{2,j}^{1,i,p}\}}$
for $i\in\mathcal{G}_{1}$ and
$w_{i}=\frac{\frac{n}{n_{2}}\left(1+\sum_{j\neq i,j\in\mathcal{G}_{2}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{2}\}\right)}{\sup_{p\in[0,1]}\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}^{2,i,p}(1-p_{j})\leq T_{1,j}^{2,i,p}\}+\left(1+\sum_{j\neq i,j\in\mathcal{G}_{2}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{2}\}\right)}$
for $i\in\mathcal{G}_{2}$. The construction of $w_{i}$ involves taking supremum over $p$, which is crucial for the proof to go through. On the one hand, it renders $w_{i}$ independent of $p_{i}$, which is a useful fact in the proof. On the other hand, it makes the weight sufficiently small in the sense that we can upper bound $w_{i}$ with the term $\sup_{p\in[0,1]}\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}^{2,i,p}(1-p_{j})\leq T_{1,j}^{2,i,p}\}$ in its denominator replaced by $\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1,j}\}$, which is another fact used in our argument.
Theorem 9.
Suppose Assumptions 1 and 2 hold. Then, the e-values defined in (20) with weights specified by (21) satisfy Condition (3).
4.3.2 Simulation studies
We shall compare the finite sample performance of the proposed method with several existing approaches through simulation studies. Throughout, we fix the sample size $n=3,000$ and set the target FDR level at $\alpha_{\mathrm{eBH}}=0.1$. For each experimental setting, we conduct 100 simulations, and report the average FDP (as an estimate of the FDR) and power over the independent simulation runs.
We begin by providing the implementation details of the proposed method. In the GBC procedure, we utilize the rejection rule based on the local FDR under the two group mixture model in Example 3, where we set $f_{0}(p)=\mathbbm{1}\{p\in[0,1]\}$ and $f_{1,i}(p)=(1-\pi_{i})(1-\kappa_{i})p^{-\kappa_{i}}$ with $\kappa_{i}\in(0,1).$ We consider the working models that link $(\pi_{i},\kappa_{i})$ with the external covariates:
$\pi_{i}$ $=\pi_{\beta_{\pi}}(x_{i})=\frac{1}{1+\exp(-\beta_{\pi,0}-\beta_{\pi,1}^{\top}x_{i})},$
$\kappa_{i}$ $=\kappa_{\beta_{\kappa}}(x_{i})=\frac{1}{1+\exp(-\beta_{\kappa,0}-\beta_{\kappa,1}^{\top}x_{i})},$
where the parameters $\beta_{\pi}=(\beta_{\pi,0},\beta_{\pi,1}^{\top})\in\mathbb{R}^{d+1}$ and $\beta_{\kappa}=(\beta_{\kappa,0},\beta_{\kappa,1}^{\top})\in\mathbb{R}^{d+1}$ can be estimated by maximizing the pseudo-log-likelihood using the EM algorithm. Please refer to Zhang and Chen (2022) for more optimization details. After obtaining the estimates $\hat{\beta}_{\pi}$ and $\hat{\beta}_{\kappa}$ from the EM algorithm, we define
$\hat{\pi}_{i}=\begin{cases}\epsilon_{1}&\text{if }1/\bigl{(}1+\exp(-\hat{\beta}_{\pi,0}-\hat{\beta}_{\pi,1}^{\top}x_{i})\bigr{)}\leq\epsilon_{1},\\ 1/\bigl{(}1+\exp(-\hat{\beta}_{\pi,0}-\hat{\beta}_{\pi,1}^{\top}x_{i})\bigr{)}&\text{if }\epsilon_{1}<1/\bigl{(}1+\exp(-\hat{\beta}_{\pi,0}-\hat{\beta}_{\pi,1}^{\top}x_{i})\bigr{)}<1-\epsilon_{2},\\ 1-\epsilon_{2}&\text{otherwise,}\end{cases}$
where winsorization is used to prevent $\hat{\pi}_{i}$ from being too close to zero or one to stabilize the algorithm. We define the rejection rule
$\hat{\varphi}_{i}(p)=\frac{1}{\hat{\pi}_{i}+(1-\hat{\pi}_{i})(1-\hat{\kappa}_{i})p^{-\hat{\kappa}_{i}}}\leq t,$
where $\hat{\kappa}_{i}=1/\bigl{(}1+\exp(-\beta_{\kappa,0}-\hat{\beta}_{\kappa,1}^{\top}x_{i})\bigr{)}$. Additionally, to speed up the procedure, we propose the following weight that is computationally less expensive:
$w_{i}=\frac{\frac{n}{n_{g}}\left(1+\sum_{j\neq i,j\in\mathcal{G}_{g}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{g}\}\right)}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{g}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{g}\}\right)+\sum_{g^{\prime}\neq g}\sum_{j\in\mathcal{G}_{g^{\prime}}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{g,j}\}}.$
We fix $\epsilon_{1}=0.1$, $\epsilon_{2}=1\times 10^{-5}$, $G=2$, and $\alpha_{\mathrm{GBC}}=\alpha_{\mathrm{eBH}}/(1+\alpha_{\mathrm{eBH}})$. We refer to this method as eBH_GBC for future reference. We compare the proposed method with the following competing methods:
  • BH: The BH procedure (Benjamini and Hochberg, 1995). We implement this method using the p.adjust function in R.
  • IHW_storey: The covariate-powered cross-weighted method with the Storey’s procedure to estimate the null-proportion (Ignatiadis and Huber, 2021). We implement this method using the ihw_bh function in the R package IHWStatsPaper.
  • IHW_betamix: The covariate-powered cross-weighted method with the beta mixture model (Ignatiadis and Huber, 2021). We implement this method using the ihw_betamix_censored function in the R package IHWStatsPaper.
  • AdaPT: The adaptive p-value thresholding procedure (Lei and Fithian, 2018). We implement this method using the adapt_glm function in the R package adaptMT.
  • SABHA: The structure adaptive BH procedure (Li and Barber, 2019). The code was downloaded from the link provided by the original paper.
To illustrate the effect of the covariate, we generate a single covariate $x_{i}$ from the standard normal distribution. Given the value of $x_{i}$, we define $\pi_{i}$ as
$\pi_{i}=\frac{\exp(a_{0}+a_{1}x_{i})}{1+\exp(a_{0}+a_{1}x_{i})},$
where $a_{0}$ and $a_{1}$ determine the baseline signal density and the informativeness of the covariate, respectively. The values of $a_{0}$ and $a_{1}$ are fixed for each simulated dataset. Specifically, we set $a_{0}$ to take on values from the set $\{3.5,1.5,1.5\}$, achieving signal densities of approximately $3\%$, $8\%$, and $18\%$, respectively, representing sparse, medium, and dense signals. Furthermore, we set $a_{1}$ to take on values from the set $\{1.5,2,2.5\}$, representing a less informative, moderately informative, and strongly informative covariate, respectively. The underlying truth $\theta_{i}$ is then simulated based on $\pi_{i}:\theta_{i}\sim\text{Bernoulli}(1-\pi_{i}).$ We next generate the covariate that affects the alternative function $f_{1,i}$. Specifically, we sample another covariate $x_{i}^{\prime}\sim\mathcal{N}(0,1)$ and define
$\eta_{i}=\frac{2\exp(a_{f}x_{i}^{\prime})}{1+\exp(a_{f}x_{i}^{\prime})},$
where we set $a_{f}\in\{0,0.5,1\}$ for no informativeness, less informativeness, and strong informativeness. Then, the z-scores are sampled from
$z_{i}\sim\mathcal{N}(\eta_{i}\mu\theta_{i},1),$
where $\mu$ denotes the signal strength with the values evenly distributed in the interval $[2.5,3.4]$. These $z$-scores are transformed into p-values using the one-sided formula $1-\Phi(z_{i})$. The p-values, along with the corresponding covariates $x_{i}$ and $x_{i}^{\prime}$, serve as the input for the structure adaptive multiple testing methods.
Figure 3: Empirical FDR and power for $a_{f}=1$.
Figure 3 shows the results for $a_{f}=1$. All methods successfully controlled the FDR at the desired level. When the signal is sparse ($a_{0}=3.5$), eBH_GBC is the most powerful method. The SABHA method has the second-best performance, while the two versions of IHW show only slight improvement over the BH procedure. When the covariate is less informative ($a_{1}=1.5$), AdaPT is less powerful than the BH procedure. As the covariate becomes strongly informative ($a_{1}=2.5$), all structure adaptive methods outperform the BH procedure in terms of power. Our proposed method exhibits the highest power in most cases, with only AdaPT surpassing eBH_GBC in power when the signal is dense ($a_{0}=1.5$) and the covariate is strongly informative ($a_{1}=2.5$). The results for the settings with $a_{f}=0$ and $a_{f}=0.5$ are deferred to Appendix C.2.
4.3.3 Real data examples
We analyzed three omics datasets: Airway (Himes et al., 2014), Bottomly (Bottomly et al., 2011), and MWAS (McDonald et al., 2018). The Airway and Bottomly datasets are transcriptomics data obtained from RNA-seq experiments, which have been analyzed previously by Ignatiadis et al. (2016); Lei and Fithian (2018); Zhang and Chen (2022). These two datasets are available in the airway and IHWStatsPaper packages, respectively. For both datasets, we used the logarithm of the “basemean” as the covariate and removed the samples with missing values, leaving us with 18,028 and 11,709 tests, respectively. We obtained the MWAS dataset from the publicly available data of the AmericanGut project (McDonald et al., 2018). We focused on a subset of subjects with ages greater than thirteen and with complete sex and country information. We excluded OTUs (clustered sequencing units representing bacterial species) observed in fewer than ten subjects, resulting in 3,394 OTUs tested using the Wilcoxon rank sum test on normalized abundances. We used the library size of samples as the external covariate.
Figure 4: Number of discoveries of various methods with the target FDR level ranging from $0.01$ to $0.1$ in three real datasets.
The results of different methods for target FDR levels ranging from $0.01$ to $0.1$ are presented in Figure 4. AdaPT and eBH_GBC are the two methods that make the most discoveries (except for the MWAS data set with a target FDR level below 0.025). For the airway dataset, AdaPT consistently produces the most discoveries when the target FDR level is high. This is due to the high signal density of this data set. For instance, when the target FDR level is 10%, AdaPT is able to identify 6,053 discoveries out of 18,028 tests. It is worth noting that the proposed method performs similarly to AdaPT when the target FDR level is below 4%. We observe a similar phenomenon for the Bottomly dataset. For the MWAS dataset, AdaPT and eBH_GBC fail to make any discoveries when the target FDR level is 1%. This is a limitation of the BC-based method, which may have reduced power when the signal is very sparse. However, as the target FDR level increases, AdaPT and eBH_GBC quickly surpass the other methods in terms of the number of discoveries. eBH_GBC outperforms AdaPT with more discoveries when the FDR level is above 5%. Overall, eBH_GBC performs comparably to AdaPT.
5 Discussions
In this paper, we have shown that the BH and BC procedures, as well as their generalized versions, are all equivalent to the e-BH procedure based on some forms of e-values. This simple insight opens up new avenues for constructing multiple testing procedures in different contexts. Specifically, leveraging this idea, one can translate the testing results from different procedures or the same procedure for different subsets of the data into e-values. By aggregating or assembling these e-values, one can obtain a combined set of e-values that contains valuable information about various multiple testing procedures and subsets of the data. We have demonstrated the potential of this strategy in three applications, backed by rigorous theoretical analysis and numerical studies. An important aspect of our proposed methods is the use of data-dependent weights, which are constructed using the leave-one-out technique to aggregate and assemble e-values. Using these weights, we can obtain valid e-values that ensure the corresponding e-BH procedure controls the FDR in finite sample. The e-BH procedure based on the data-dependent weight is often more powerful than the unweighted version, but it comes at the cost of more expensive computation. Through simulations, we have demonstrated that a more cost-friendly version of these weights exists, which achieves nearly identical performance with a much lower computational burden.
We envision the idea of aggregating different multiple testing results through e-values to be useful in other contexts, such as meta-analysis or federated learning. Other interesting future research problems include finding the optimal way of combining the e-values with respect to certain criteria and investigating the robustness of the proposed methods when the data exhibit dependence.
Appendix
Appendix A Proof of the main results
We first state the following propositions whose proofs are deferred to Appendix  B. These results will be used frequently in the subsequent proofs of the main theorems.
Proposition 5 (Lemma 6 of Barber et al. [2020]).
Let $T_{\mathrm{BC},i}$ be the threshold for the BC methods when $p_{i}$ is replaced with $\min\{p_{i},1-p_{i}\}$. For any $i$, $j$, if $\min(p_{i},p_{j})\geq 1-\max\{T_{\mathrm{BC},i},T_{\mathrm{BC},j}\}$, then we have $T_{\mathrm{BC},i}=T_{\mathrm{BC},j}$.
Proposition 6.
Suppose the assumptions in Proposition 4 hold. Let $T_{i}$ be the threshold for the generalized BC methods when $p_{i}$ is replaced with $\min\{p_{i},1-p_{i}\}$. For any $i$, $j$, if $\max\{\varphi_{i}(1-p_{i}),\varphi_{j}(1-p_{j})\}\leq\max\{T_{i},T_{j}\}$, then we have $T_{i}=T_{j}$.
A.1 Proof of Theorem 1
Proof.
Theorem 1 is a special case of Theorem 2 by choosing $\varphi_{i}$ as the identity function for all $1\leq i\leq n$. ∎
A.2 Proof of Theorem 2
Proof.
Let $\hat{k}$ be the cardinality of $\mathcal{S}_{\mathrm{GBH}}$. If $i\notin\mathcal{S}_{\mathrm{GBH}}$, then $\varphi_{i}(p_{i})>T_{\mathrm{GBH}}$ and thus $e_{i}=0$. This implies that the $i$th hypothesis is not rejected by the e-BH procedure, and hence $\mathcal{S}_{\mathrm{eBH}}\subset\mathcal{S}_{\mathrm{GBH}}$. Conversely, if $i\in\mathcal{S}_{\mathrm{GBH}}$, we have $\varphi_{i}(p_{i})\leq T_{\mathrm{GBH}}\leq\bar{F}^{-1}\left(\frac{\alpha\hat{k}}{n}\right)$, leading to
$e_{i}=\frac{1}{\bar{F}(T_{\mathrm{GBH}})}\geq\frac{n}{\alpha\hat{k}}$
as $\bar{F}$ is strictly increasing. Define $e_{(i)}=n\mathbbm{1}\{q_{(i)}\leq T_{\mathrm{GBH}}\}/\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}})$ for $1\leq i\leq n,$ where $q_{(1)}\leq\dots\leq q_{(n)}$ are the order statistics of $\{q_{i}=\varphi_{i}(p_{i})\}^{n}_{i=1}$. Let $\hat{k}$ represent the maximum $i$ for which $q_{(i)}\leq T_{\mathrm{GBH}}$. We get
$e_{(\hat{k})}\geq\frac{n}{\alpha\hat{k}},$ (A.1)
which indicates that $|\mathcal{S}_{\mathrm{eBH}}|\geq\hat{k}$. Because $\mathcal{S}_{\mathrm{eBH}}\subset\mathcal{S}_{\mathrm{GBH}}$, it is clear that $\mathcal{S}_{\mathrm{eBH}}=\mathcal{S}_{\mathrm{GBH}}$. ∎
A.3 Proof of Theorem 3
Let $\hat{k}$ be the cardinality of $\mathcal{S}_{\mathrm{GBC}}$. If $i\notin\mathcal{S}_{\mathrm{GBC}}$, then $\varphi_{i}(p_{i})>T_{\mathrm{GBC}}$ and thus $e_{i}=0$. Hence the $i$th hypothesis is not rejected by the e-BH procedure, which implies that $\mathcal{S}_{\mathrm{eBH}}\subset\mathcal{S}_{\mathrm{GBC}}$. For the other direction, note that if $i\in\mathcal{S}_{\mathrm{GBC}}$, then $\varphi_{i}(p_{i})\leq T_{\mathrm{GBC}}$ and $1+\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{\mathrm{GBC}}\}\leq\hat{k}\alpha$. Hence, we have
$e_{i}\geq\frac{n}{\hat{k}\alpha}.$
We sort the e-values in descending order as $e_{(1)}\geq\dots\geq e_{(n)}$. It is clear that $e_{(\hat{k})}\geq n/(\hat{k}\alpha)$. Thus, $|\mathcal{S}_{\mathrm{eBH}}|\geq\hat{k}$, which implies that $\mathcal{S}_{\mathrm{eBH}}=\mathcal{S}_{\mathrm{GBC}}$.
A.4 Proof of Theorem 4
Proof.
Consider the BH procedure and observe that for a given number of rejections $R_{\mathrm{BH}}$, $T_{\mathrm{BH}}=T_{\mathrm{BH}}(R_{\mathrm{BH}})$ is a deterministic function of $R_{\mathrm{BH}}$. Let $R_{\mathrm{BH}}(p_{i}\rightarrow 0)$ be the number of rejections obtained by replacing the p-value $p_{i}$ with 0. Using the above fact and that the weights are independent of the p-values, we have
$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[w_{\mathrm{BH},i}e_{\mathrm{BH},i}]=$ $\sum_{i\in\mathcal{H}_{0}}w_{\mathrm{BH},i}\mathbb{E}\left[\frac{1}{T_{\mathrm{BH}}}\mathbbm{1}\{p_{i}\leq T_{\mathrm{BH}}\}\right]$
$=$ $\sum_{i\in\mathcal{H}_{0}}w_{\mathrm{BH},i}\sum_{k=1}^{n}\mathbb{E}\left[\frac{1}{T_{\mathrm{BH}}(k)}\mathbbm{1}\{p_{i}\leq T_{\mathrm{BH}}(k),R_{\mathrm{BH}}(p_{i}\rightarrow 0)=k\}\right]$
$\leq$ $\sum_{i\in\mathcal{H}_{0}}w_{\mathrm{BH},i}\sum_{k=1}^{n}\mathbb{E}[\mathbbm{1}\{R_{\mathrm{BH}}(p_{i}\rightarrow 0)=k\}]$
$=$ $\sum_{i\in\mathcal{H}_{0}}w_{\mathrm{BH},i}.$
For the BC procedure, denote $\tilde{p}_{i}\coloneqq\min\{p_{i},1-p_{i}\}$ and $\mathbf{p}_{-i}\coloneqq\{p_{1},\dots,p_{i-1},\,\tilde{p}_{i},\,p_{i+1},\dots,p_{n}\}$. By viewing $T_{\mathrm{BC}}$ as a function of the p-values, we define $T_{\mathrm{BC},i}=T_{\mathrm{BC}}(\mathbf{p}_{-i})$. Let $\mathcal{F}_{i}$ be the sigma algebra generated by $\mathbf{p}_{-i}$. Then, we have
$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[w_{\mathrm{BC},i}e_{\mathrm{BC},i}]=$ $\sum_{i\in\mathcal{H}_{0}}w_{\mathrm{BC},i}\mathbb{E}\left[\frac{n\mathbbm{1}\{p_{i}\leq T_{\mathrm{BC}}\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC}}\}}\right]$
$=$ $\sum_{i\in\mathcal{H}_{0}}w_{\mathrm{BC},i}\mathbb{E}\left[\frac{n\mathbbm{1}\{p_{i}\leq T_{\mathrm{BC},i}\}}{1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}}\right]$
$=$ $\sum_{i\in\mathcal{H}_{0}}w_{\mathrm{BC},i}\mathbb{E}\left[\frac{n}{1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}}\mathbb{E}[\mathbbm{1}\{p_{i}\leq T_{\mathrm{BC},i}\}\lvert\mathcal{F}_{i}]\right]$
$\leq$ $\sum_{i\in\mathcal{H}_{0}}w_{\mathrm{BC},i}\mathbb{E}\left[\frac{n}{1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}}\mathbb{E}[\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}\lvert\mathcal{F}_{i}]\right]$
$=$ $\sum_{i\in\mathcal{H}_{0}}w_{\mathrm{BC},i}\mathbb{E}\left[\frac{n\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}}{1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}}\right],$
where (i) we have used the fact that $T_{\mathrm{BC}}=T_{\mathrm{BC},i}$ when $p_{i}\leq T_{BC}<0.5$ to get the second equation, (ii) the third equation follows because $T_{\mathrm{BC},i}$ are measurable with respect to $\mathcal{F}_{i}$, and (iii) the inequality is due to the assumption that $p_{i}$ follows the super-uniform distribution on $[0,1]$, and thus it satisfies Condition (8) under the null.
By Proposition (5), we have
$\frac{\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}}{1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}}=\frac{\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}}{\sum_{j=1}^{n}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j}\}}.$ (A.2)
If $p_{i}<1-T_{\mathrm{BC},i}$, both sides are equal to 00. If $p_{i}\geq 1-T_{\mathrm{BC},i}$, we claim that $\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}=\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j}\}$. Indeed, if $p_{j}\geq 1-T_{\mathrm{BC},i}$ but $p_{j}<1-T_{\mathrm{BC},j}$, we have $T_{\mathrm{BC},i}>T_{\mathrm{BC},j}$. This implies that $\min(p_{i},p_{j})\geq 1-\max\{T_{\mathrm{BC},i},T_{\mathrm{BC},j}\}$. By Proposition 5, we have $T_{\mathrm{BC},i}=T_{\mathrm{BC},j}$, which contradicts with the fact that $T_{\mathrm{BC},i}>T_{\mathrm{BC},j}$. The other direction can be proved similarly. Hence we have
$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[w_{\mathrm{BC},i}e_{\mathrm{BC},i}]\leq n\max_{i\in\mathcal{H}_{0}}w_{\mathrm{BC},i}\mathbb{E}\left[\frac{\sum_{i\in\mathcal{H}_{0}}\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}}{\sum_{j=1}^{n}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j}\}}\right]\leq n\max_{i}w_{\mathrm{BC},i}.$
By the assumption $w_{\mathrm{BH},i}+\max_{j}w_{\mathrm{BC},j}\leq 1$, we have
$\mathbb{E}[w_{\mathrm{BH},i}e_{\mathrm{BH},i}+w_{\mathrm{BC},i}e_{\mathrm{BC},i}]\leq\sum_{i=1}^{n}w_{\mathrm{BH},i}+n\max_{j}w_{\mathrm{BC},j}\leq n.$
A.5 Proof of Theorem 5
Proof.
Consider the BH procedure and observe that for a given number of rejections $R_{\mathrm{BH}}$, $T_{\mathrm{BH}}=T_{\mathrm{BH}}(R_{\mathrm{BH}})$ is a deterministic function of $R_{\mathrm{BH}}$. Let $R_{\mathrm{BH}}(p_{i}\rightarrow 0)$ be the number of rejections obtained by replacing the p-value $p_{i}$ with 0. Using the above fact and the leave-one-out argument, we have
$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[w_{\mathrm{BH},i}e_{\mathrm{BH},i}]$
$=$ $\sum_{i\in\mathcal{H}_{0}}\mathbb{E}\left[\frac{T_{\mathrm{BH},i}}{T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j,i}\}\right)}\frac{1}{T_{\mathrm{BH}}}\mathbbm{1}\{p_{i}\leq T_{\mathrm{BH}}\}\right]$
$=$ $\sum_{i\in\mathcal{H}_{0}}\sum_{k=1}^{n}\mathbb{E}\left[\frac{T_{\mathrm{BH},i}}{T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j,i}\}\right)}\frac{1}{T_{\mathrm{BH}}(k)}\mathbbm{1}\{p_{i}\leq T_{\mathrm{BH}}(k),R_{\mathrm{BH}}(p_{i}\rightarrow 0)=k\}\right],$
where to get the second equality we have used the fact that when the $i$th hypothesis is rejected (i.e., $p_{i}\leq T_{\mathrm{BH}}$), $R_{\mathrm{BH}}=R_{\mathrm{BH}}(p_{i}\rightarrow 0)$. Let $\mathcal{F}_{i}$ be the sigma algebra generated by $\{p_{1},\cdots,p_{i-1},0,p_{i+1},\cdots,p_{n}\}$. We have
$\sum_{k=1}^{n}\mathbb{E}\left[\frac{T_{\mathrm{BH},i}}{T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j,i}\}\right)}\frac{1}{T_{\mathrm{BH}}(k)}\mathbbm{1}\{p_{i}\leq T_{\mathrm{BH}}(k),R_{\mathrm{BH}}(p_{i}\rightarrow 0)=k\}\Bigg{|}\mathcal{F}_{i}\right]$
$=$ $\sum_{k=1}^{n}\frac{T_{\mathrm{BH},i}}{T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j,i}\}\right)}\frac{1}{T_{\mathrm{BH}}(k)}P(p_{i}\leq T_{\mathrm{BH}}(k))\mathbb{E}[\mathbbm{1}\{R_{\mathrm{BH}}(p_{i}\rightarrow 0)=k\}\lvert\mathcal{F}_{i}]$
$\leq$ $\frac{T_{\mathrm{BH},i}}{T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j,i}\}\right)},$
where we used the fact that $T_{\mathrm{BH},i}$ and $T_{\mathrm{BC},j,i}$ are both measurable with respect to $\mathcal{F}_{i}$. Thus,
$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[w_{\mathrm{BH},i}e_{\mathrm{BH},i}]$
$\leq$ $\sum_{i\in\mathcal{H}_{0}}\mathbb{E}\left[\frac{T_{\mathrm{BH},i}}{T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j,i}\}\right)}\right]$
$\leq$ $\sum_{i\in\mathcal{H}_{0}}\mathbb{E}\left[\frac{\max_{i}T_{\mathrm{BH},i}}{\max_{i}T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j,i}\}\right)}\right].$
Note that $T_{\mathrm{BC},j,i}\geq T_{\mathrm{BC},j}$ and hence $\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j,i}\}\geq\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j}\}$. It follows that
$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[w_{\mathrm{BH},i}e_{\mathrm{BH},i}]\leq$ $\sum_{i\in\mathcal{H}_{0}}\mathbb{E}\left[\frac{\max_{i}T_{\mathrm{BH},i}}{\max_{i}T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j}\}\right)}\right]$
$\leq$ $\mathbb{E}\left[\frac{n\max_{i}T_{\mathrm{BH},i}}{\max_{i}T_{\mathrm{BH},i}+\frac{1}{n}\sum_{j=1}^{n}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j}\}}\right].$
For the BC procedure, let ~$\widetilde{\mathcal{F}}_{i}$ be the sigma algebra generated by $\mathbf{p}_{-i}$. Then, we have
$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[w_{\mathrm{BC},i}e_{\mathrm{BC},i}]$
$=$ $\sum_{i\in\mathcal{H}_{0}}\mathbb{E}\left[\frac{\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC}}\}\right)}{\max_{i}T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC}}\}\right)}\frac{n\mathbbm{1}\{p_{i}\leq T_{\mathrm{BC}}\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC}}\}}\right]$
$=$ $\sum_{i\in\mathcal{H}_{0}}\mathbb{E}\left[\frac{\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}\right)}{\max_{i}T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}\right)}\frac{n\mathbbm{1}\{p_{i}\leq T_{\mathrm{BC},i}\}}{1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}}\right]$
$\leq$ $\sum_{i\in\mathcal{H}_{0}}\mathbb{E}\left[\frac{1}{\max_{i}T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}\right)}\mathbb{E}[\mathbbm{1}\{p_{i}\leq T_{\mathrm{BC},i}\}\lvert\widetilde{\mathcal{F}}_{i}]\right]$
$=$ $\sum_{i\in\mathcal{H}_{0}}\mathbb{E}\left[\frac{\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}}{\max_{i}T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}\right)}\right],$
where (i) we have used the fact that $T_{\mathrm{BC}}=T_{\mathrm{BC},i}$ when $p_{i}\leq T_{BC}<0.5$ to get the second equation, (ii) the third equation follows because both $\max_{i}T_{\mathrm{BH},i}$ and $T_{\mathrm{BC},i}$ are measurable with respect to ~$\widetilde{\mathcal{F}}_{i}$, and (iii) the inequality is due to the assumption that $p_{i}$ follows the super-uniform distribution on $[0,1]$ and thus satisfies Condition (8).
By Proposition 5 and following the same discussion used to prove (A.2), we obtain
$\frac{\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}}{\max_{i}T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}\right)}=\frac{\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}}{\max_{i}T_{\mathrm{BH},i}+\frac{1}{n}\left(\sum_{j=1}^{n}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j}\}\right)}.$
Hence we have
$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[w_{\mathrm{BC},i}e_{\mathrm{BC},i}]=$ $\mathbb{E}\left[\frac{\sum_{i\in\mathcal{H}_{0}}\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}}{\max_{i}T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j}\}\right)}\right]$
$\leq$ $\mathbb{E}\left[\frac{\sum_{i=1}^{n}\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}}{\max_{i}T_{\mathrm{BH},i}+\frac{1}{n}\sum_{j=1}^{n}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j}\}}\right].$
Combining the arguments, we obtain
$\sum_{i\in\mathcal{H}_{0}}\{\mathbb{E}[w_{\mathrm{BH},i}e_{\mathrm{BH},i}]+\mathbb{E}[w_{\mathrm{BC},i}e_{\mathrm{BC},i}]\}\leq\mathbb{E}\left[\frac{n\max_{i}T_{\mathrm{BH},i}+\sum_{i=1}^{n}\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}}{\max_{i}T_{\mathrm{BH},i}+\frac{1}{n}\sum_{j=1}^{n}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j}\}}\right]=n.$
A.6 Proof of Theorem 6
Proof.
We only present the proof for the case of $G=2$. The arguments can be generalized to the general case without essential difficulty. Let us consider the first group. Following the same discussion in the proof of Theorem 4, we have
$\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}\mathbb{E}[w_{i}e_{i}]=$ $\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}w_{i}\mathbb{E}\left[\frac{n_{1}\mathbbm{1}\{p_{i}\leq T_{1}\}}{1+\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1}\}}\right]$
$=$ $\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}w_{i}\mathbb{E}\left[\frac{n_{1}\mathbbm{1}\{p_{i}\leq T_{1,i}\}}{1+\sum_{j\in\mathcal{G}_{1},j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{1,i}\}}\right]$
$\leq$ $\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}w_{i}\mathbb{E}\left[\frac{n_{1}\mathbbm{1}\{p_{i}\geq 1-T_{1,i}\}}{1+\sum_{j\in\mathcal{G}_{1},j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{1,i}\}}\right].$
By Proposition 5 and the discussion in the proof of Theorem 4, we have
$\frac{n_{1}\mathbbm{1}\{p_{i}\geq 1-T_{1,i}\}}{1+\sum_{j\in\mathcal{G}_{1},j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{1,i}\}}=\frac{n_{1}\mathbbm{1}\{p_{i}\geq 1-T_{1,i}\}}{\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1,j}\}}.$
Thus,
$\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}\mathbb{E}[w_{i}e_{i}]\leq\max_{i\in\mathcal{G}_{1}}w_{i}\mathbb{E}\left[\frac{n_{1}\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}\mathbbm{1}\{p_{i}\geq 1-T_{1,i}\}}{\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1,j}\}}\right]\leq n_{1}\max_{i\in\mathcal{G}_{1}}w_{i}.$
Using the same argument for the second group, we obtain
$\sum_{i\in\mathcal{G}_{2}\cap\mathcal{H}_{0}}\mathbb{E}[w_{i}e_{i}]\leq n_{2}\max_{i\in\mathcal{G}_{2}}w_{i}.$
Hence, by (17), we deduce that
$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[w_{i}e_{i}]\leq n,$
which completes the proof. ∎
A.7 Proof of Theorem 7
Proof.
We only present the proof for the case of $G=2$. The arguments can be generalized to the general case without essential difficulty. Let us consider the first group. Following the same discussion in the proof of Theorem 4, we have
$\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}\mathbb{E}[w_{i}e_{i}]$
$=$ $\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}\mathbb{E}\left[\frac{\frac{n}{n_{1}}\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1}\}\right)}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1}\}\right)+\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{p_{j}\geq 1-T_{2,j}\}}\frac{n_{1}\mathbbm{1}\{p_{i}\leq T_{1}\}}{1+\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1}\}}\right]$
$=$ $\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}\mathbb{E}\left[\frac{n\mathbbm{1}\{p_{i}\leq T_{1,i}\}}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1,i}\}\right)+\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{p_{j}\geq 1-T_{2,j}\}}\right]$
$\leq$ $\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}\mathbb{E}\left[\frac{n\mathbbm{1}\{p_{i}\geq 1-T_{1,i}\}}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1,i}\}\right)+\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{p_{j}\geq 1-T_{2,j}\}}\right].$
By Proposition 5 and the discussion in the proof of Theorem 4, we have
$\frac{n\mathbbm{1}\{p_{i}\geq 1-T_{1,i}\}}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1,i}\}\right)+\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{p_{j}\geq 1-T_{2,j}\}}$
$=$ $\frac{n\mathbbm{1}\{p_{i}\geq 1-T_{1,i}\}}{\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1,j}\}+\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{p_{j}\geq 1-T_{2,j}\}}.$
Thus,
$\sum_{i\in\mathcal{H}_{0}\cap\mathcal{G}_{1}}\mathbb{E}[w_{i}e_{i}]\leq\frac{n\sum_{i\in\mathcal{G}_{1}}\mathbbm{1}\{p_{i}\geq 1-T_{1,i}\}}{\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1,j}\}+\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{p_{2,j}\geq 1-T_{2,j}\}}.$
Using the same argument for the second group, we obtain
$\sum_{i\in\mathcal{H}_{0}\cap\mathcal{G}_{2}}\mathbb{E}[w_{i}e_{i}]\leq\frac{n\sum_{i\in\mathcal{G}_{2}}\mathbbm{1}\{p_{i}\geq 1-T_{2,i}\}}{\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1,j}\}+\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{p_{2,j}\geq 1-T_{2,j}\}}.$
Hence,
$\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}\mathbb{E}[w_{i}e_{i}]+\sum_{i\in\mathcal{G}_{2}\cap\mathcal{H}_{0}}\mathbb{E}[w_{i}e_{i}]\leq n.$
A.8 Proof of Theorem 8
Proof.
We only prove the result for $G=2$ (the same argument applies to the case of a general $G$). Note that when $\hat{\varphi}_{i}(p_{i})\leq T_{g}\leq T_{g,\mathrm{up}}<\hat{\varphi}_{i}(0.5)$, Assumption 2 implies that $p_{i}<0.5$ and hence $p_{i}=\tilde{p}_{i}$. By the definition of $T_{g,i}$, $T_{g}=T_{g,i}$. Therefore, for the first group, we have
$\sum_{i\in\mathcal{H}_{0}\cap\mathcal{G}_{1}}\mathbb{E}[w_{i}e_{i}]=$ $\sum_{i\in\mathcal{H}_{0}\cap\mathcal{G}_{1}}\mathbb{E}\left[\frac{n_{1}w_{i}\mathbbm{1}\{\hat{\varphi}_{i}(p_{i})\leq T_{1}\}}{1+\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{i})\leq T_{1}\}}\right]$
$=$ $\sum_{i\in\mathcal{H}_{0}\cap\mathcal{G}_{1}}\mathbb{E}\left[\frac{n_{1}w_{i}\mathbbm{1}\{\hat{\varphi}_{i}(p_{i})\leq T_{1,i}\}}{1+\sum_{j\in\mathcal{G}_{1},j\neq i}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{i})\leq T_{1,i}\}}\right].$
Let $\mathcal{F}_{i}$ denote the sigma algebra generated by $\mathbf{p}_{-i}$. Since $T_{1,i}$, $\hat{\varphi}_{j}$ and $w_{i}$ are all measurable with respect to $\mathcal{F}_{i}$, we deduce that
$\sum_{i\in\mathcal{H}_{0}\cap\mathcal{G}_{1}}\mathbb{E}[w_{i}e_{i}]=$ $\sum_{i\in\mathcal{H}_{0}\cap\mathcal{G}_{1}}\mathbb{E}\left[\frac{n_{1}w_{i}}{1+\sum_{j\in\mathcal{G}_{1},j\neq i}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{i})\leq T_{1,i}\}}\mathbb{E}[\mathbbm{1}\{\hat{\varphi}_{i}(p_{i})\leq T_{1,i}\}\lvert\mathcal{F}_{i}]\right]$
$\leq$ $\sum_{i\in\mathcal{H}_{0}\cap\mathcal{G}_{1}}\mathbb{E}\left[\frac{n_{1}w_{i}}{1+\sum_{j\in\mathcal{G}_{1},j\neq i}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{i})\leq T_{1,i}\}}\mathbb{E}[\mathbbm{1}\{\hat{\varphi}_{i}(1-p_{i})\leq T_{1,i}\}\lvert\mathcal{F}_{i}]\right]$
$=$ $\sum_{i\in\mathcal{H}_{0}\cap\mathcal{G}_{1}}\mathbb{E}\left[\frac{n_{1}w_{i}\mathbbm{1}\{\hat{\varphi}_{i}(1-p_{i})\leq T_{1,i}\}}{1+\sum_{j\in\mathcal{G}_{1},j\neq i}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{i})\leq T_{1,i}\}}\right],$
where we use Assumption 1(C) to get the inequality.
By Proposition  6 and Assumption 2, we have
$\frac{\mathbbm{1}\{\hat{\varphi}_{i}(1-p_{i})\leq T_{1,i}\}}{1+\sum_{j\in\mathcal{G}_{1},j\neq i}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1,i}\}}$ $=\frac{\mathbbm{1}\{\hat{\varphi}_{i}(1-p_{i})\leq T_{1,i}\}}{1+\sum_{j\in\mathcal{G}_{1},j\neq i}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1,j}\}}$
$=\frac{\mathbbm{1}\{\hat{\varphi}_{i}(1-p_{i})\leq T_{1,i}\}}{\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1,j}\}}.$
If $\hat{\varphi}_{i}(1-p_{i})>T_{1,i}$, both sides are equal to 00. If $\hat{\varphi}_{i}(1-p_{i})\leq T_{1,i}$, we claim that $\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1,i}\}=\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1,j}\}$. Indeed, if $\hat{\varphi}_{j}(1-p_{j})>T_{1,i}$ but $\hat{\varphi}_{j}(1-p_{j})\leq T_{1,j}$, then we have $T_{1,i}<T_{1,j}$. Hence, $\hat{\varphi}_{i}(1-p_{i})\leq T_{1,i}<T_{1,j}$. By proposition 6, we have $T_{1,i}=T_{1,j}$, which contradicts with $T_{1,i}<T_{1,j}$. The other direction can be proved similarly.
If the weights $\{w_{i}\}$ are independent of the p-values and covariate information, then we have
$\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}\mathbb{E}[w_{i}e_{i}]\leq n_{1}\max_{i\in\mathcal{G}_{1}}w_{i}\mathbb{E}\left[\frac{\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}\mathbbm{1}\{\hat{\varphi}_{i}(1-p_{i})\leq T_{1,i}\}}{\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1,j}\}}\right]\leq n_{1}\max_{i\in\mathcal{G}_{1}}w_{i}.$
Using the same argument for the second group, we obtain
$\sum_{i\in\mathcal{G}_{2}\cap\mathcal{H}_{0}}\mathbb{E}[w_{i}e_{i}]\leq n_{2}\max_{i\in\mathcal{G}_{2}}w_{i}.$
Hence, by (17), we deduce that
$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[w_{i}]\leq n.$
A.9 Proof of Theorem 9
Proof.
We only prove the result for $G=2$ (the same argument applies to the case of a general $G$). Note that when $\hat{\varphi}_{i}(p_{i})\leq T_{g}\leq T_{g,\mathrm{up}}<\hat{\varphi}_{i}(0.5)$, Assumption 2 implies that $p_{i}<0.5$ and thus $p_{i}=\tilde{p}_{i}$. Thus, $T_{g}=T_{g,i}$ by the definition of $T_{g,i}$. Therefore, for the first group, we have
$\sum_{i\in\mathcal{H}_{0}\cap\mathcal{G}_{1}}\mathbb{E}[w_{i}e_{i}]$
$=$ $\sum_{i\in\mathcal{H}_{0}\cap\mathcal{G}_{1}}\mathbb{E}\left[\frac{\frac{n}{n_{1}}\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1}\}\right)}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1}\}\right)+\sup_{p\in[0,1]}\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{\hat{\varphi}_{j}^{1,i,p}(1-p_{j})\leq T_{2,j}^{1,i,p}\}}\right.$
×$\qquad\times\left.\frac{n_{1}\mathbbm{1}\{\hat{\varphi}_{i}(p_{i})\leq T_{1}\}}{1+\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1}\}}\right]$
$\leq$ $\sum_{i\in\mathcal{H}_{0}\cap\mathcal{G}_{1}}\mathbb{E}\left[\frac{n\mathbbm{1}\{\hat{\varphi}_{i}(p_{i})\leq T_{1,i}\}}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1,i}\}\right)+\sup_{p\in[0,1]}\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{\hat{\varphi}_{j}^{1,i,p}(1-p_{j})\leq T_{2,j}^{1,i,p}\}}\right].$
Let $\mathcal{F}_{i}$ denote the sigma algebra generated by $\mathbf{p}_{-i}$. Since $T_{1,i}$, $\hat{\varphi}_{j}$ for $j\in\mathcal{G}_{1}$, $\hat{\varphi}_{j}^{1,i,p}$, and $T_{2,j}^{1,i,p}$ for $j\in\mathcal{G}_{2}$ are all measurable with respect to $\mathcal{F}_{i}$, we deduce that
$\sum_{i\in\mathcal{H}_{0}\cap\mathcal{G}_{1}}\mathbb{E}[w_{i}e_{i}]$
$=$ $\sum_{i\in\mathcal{H}_{0}\cap\mathcal{G}_{1}}\mathbb{E}\left[\mathbb{E}\left[\frac{n\mathbbm{1}\{\hat{\varphi}_{i}(p_{i})\leq T_{1,i}\}}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1,i}\}\right)+\sup_{p\in[0,1]}\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{\hat{\varphi}_{j}^{1,i,p}(1-p_{j})\leq T_{2,j}^{1,i,p}\}}\Bigg{\lvert}\mathcal{F}_{i}\right]\right]$
$\leq$ $\sum_{i\in\mathcal{H}_{0}\cap\mathcal{G}_{1}}\mathbb{E}\left[\mathbb{E}\left[\frac{n\mathbbm{1}\{\hat{\varphi}_{i}(1-p_{i})\leq T_{1,i}\}}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1,i}\}\right)+\sup_{p\in[0,1]}\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{\hat{\varphi}_{j}^{1,i,p}(1-p_{j})\leq T_{2,j}^{1,i,p}\}}\Bigg{\lvert}\mathcal{F}_{i}\right]\right]$
$\leq$ $\sum_{i\in\mathcal{H}_{0}\cap\mathcal{G}_{1}}\mathbb{E}\left[\frac{n\mathbbm{1}\{\hat{\varphi}_{i}(1-p_{i})\leq T_{1,i}\}}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1,i}\}\right)+\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{2,j}\}}\right],$
where we have used Assumption 1(C) to obtain the first inequality and the second inequality is due to the fact that
$\sum_{j\in G_{2}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{2,j}\}$ $=\sum_{j\in G_{2}}\mathbbm{1}\{\hat{\varphi}_{j}^{1,i,p}(1-p_{j})\leq T_{2,j}^{1,i,p}\}\lvert_{p=p_{i}}$
$\leq\sup_{p\in[0,1]}\sum_{j\in G_{2}}\mathbbm{1}\{\hat{\varphi}_{j}^{1,i,p}(1-p_{j})\leq T_{2,j}^{1,i,p}\}.$
By Proposition 6 and the argument in the proof of Theorem 8, we have
$\frac{n\mathbbm{1}\{\hat{\varphi}_{i}(1-p_{i})\leq T_{1,i}\}}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1,i}\}\right)+\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{2,j}\}}$
$=$ $\frac{n\mathbbm{1}\{\hat{\varphi}_{i}(1-p_{i})\leq T_{1,i}\}}{\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1,j}\}+\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{2,j}\}}.$
Hence, for the first group, we get
$\sum_{i\in\mathcal{H}_{0}\cap\mathcal{G}_{1}}\mathbb{E}[w_{i}e_{i}]\leq\mathbb{E}\left[\frac{n\sum_{i\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{i}(1-p_{i})\leq T_{1,i}\}}{\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1,j}\}+\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{2,j}\}}\right].$
Following the same discussion, we have
$\sum_{i\in\mathcal{H}_{0}\cap\mathcal{G}_{2}}\mathbb{E}[w_{i}e_{i}]\leq\mathbb{E}\left[\frac{n\sum_{i\in\mathcal{G}_{2}}\mathbbm{1}\{\hat{\varphi}_{i}(1-p_{i})\leq T_{1,i}\}}{\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1,j}\}+\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{2,j}\}}\right].$
Combing the above results leads to
$\sum_{g=1}\sum_{i\in\mathcal{H}_{0}\cap\mathcal{G}_{g}}\mathbb{E}[w_{i}e_{i}]\leq n.$
Appendix B Additional proofs
B.1 Proof of Proposition 1
Proof.
Note that
[…]
$\leq$ $\sum_{i=1}^{n}\frac{\mathbbm{1}\{ie_{(i)}\geq n/\alpha,H_{(i)}\text{ is under the null}\}}{1\vee i}$
$\leq$ $\sum_{i=1}^{n}\mathbbm{1}\{H_{(i)}\text{ is under the null}\}\frac{\alpha e_{(i)}}{n}=\frac{\alpha}{n}\sum_{i\in\mathcal{H}_{0}}e_{i}.$
Under Condition (3), we have
$\text{FDR}=\mathbb{E}[\text{FDP}]\leq\alpha.$
B.2 Proof of Proposition 2
Proof.
Let $V_{i}=\mathbbm{1}\{H_{i}\text{ is rejected}\}$. We have
$\mathrm{FDP}(T_{\mathrm{GBH}})=\sum_{i\in\mathcal{H}_{0}}\frac{V_{i}}{R(T_{\mathrm{GBH}})\vee 1}=\sum_{i\in\mathcal{H}_{0}}\frac{V_{i}}{\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}})}\frac{\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}})}{R(T_{\mathrm{GBH}})\vee 1}\leq\alpha\sum_{i\in\mathcal{H}_{0}}\frac{V_{i}}{\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}})}.$
Therefore, it suffices to prove
$\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}}\frac{V_{i}}{\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}})}\right]\leq 1.$
Observing that, for a given $R$, $T_{\mathrm{GBH}}=T_{\mathrm{GBH}}(R)$ is a deterministic function of $R$, we have:
$\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}}\frac{V_{i}}{\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}})}\right]=\sum_{i\in H_{0}}\sum_{k=1}^{n}\mathbb{E}\left[\frac{V_{i}\mathbbm{1}\{R=k\}}{\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}}(k))}\right]=\sum_{i\in H_{0}}\sum_{k=1}^{n}\mathbb{E}\left[\frac{V_{i}\mathbbm{1}\{R(p_{i}\rightarrow 0)=k\}}{\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}}(k))}\right],$
where $R(p_{i}\rightarrow 0)$ is the number of rejections obtained by replacing the p-value $p_{i}$ with 0. To clarify the second equality, note that if $V_{i}=0$, the equation is trivially true. When $V_{i}=1$, setting $p_{i}$ to 00 does not change the number of rejections. Using the assumption that $F_{j}(x)=c_{j}h(x)$, we obtain:
$\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}}\frac{V_{i}}{\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}})}\right]$ $=\sum_{i\in\mathcal{H}_{0}}\sum_{k=1}^{n}\frac{1}{\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}}(k))}\mathbb{E}[\mathbbm{1}\{\varphi_{i}(p_{i})\leq T_{\mathrm{GBH}}(k)\}]\mathbb{E}[\mathbbm{1}\{R(p_{i}\rightarrow 0)=k\}]$
$\leq\sum_{i\in\mathcal{H}_{0}}\sum_{k=1}^{n}\frac{F_{i}(T_{\mathrm{GBH}}(k))}{\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}}(k))}\mathbb{E}[\mathbbm{1}\{R(p_{i}\rightarrow 0)=k\}]$
$=\sum_{i\in\mathcal{H}_{0}}\sum_{k=1}^{n}\frac{c_{i}}{\sum_{j=1}^{n}c_{j}}\mathbb{E}[\mathbbm{1}\{R(p_{i}\rightarrow 0)=k\}]$
$=\frac{\sum_{i\in\mathcal{H}_{0}}c_{i}}{\sum_{j=1}^{n}c_{j}}\leq 1.$
B.3 Proof of Proposition 3
Proof.
Recall from the proof of Proposition 2 that $T_{\mathrm{GBH}}=T_{\mathrm{GBH}}(R_{\mathrm{GBH}})$ is a deterministic function of $R_{\mathrm{GBH}}$. Additionally, for any $i\in\mathcal{H}_{0}$ and $\varphi_{i}(p_{i})\leq T_{\mathrm{GBH}}$, replacing the p-value $p_{i}$ with 0 does not change the number of rejections, i.e., $R_{\mathrm{GBH}}(p_{i}\leftarrow 0)=R_{\mathrm{GBH}}$. Therefore, we can infer that
$\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}}e_{i}\right]=$ $\sum_{i\in\mathcal{H}_{0}}\sum^{n}_{k=1}\mathbb{E}\left[\frac{n\mathbbm{1}\{\varphi_{i}(p_{i})\leq T_{\mathrm{GBH}},R_{\mathrm{GBH}}=k\}}{\sum^{n}_{j=1}F_{j}(T_{\mathrm{GBH}})}\right]$
$=$ $\sum_{i\in\mathcal{H}_{0}}\sum^{n}_{k=1}\frac{n\mathbb{E}[\mathbbm{1}\{\varphi_{i}(p_{i})\leq T_{\mathrm{GBH}}(k),R_{\mathrm{GBH}}=k\}]}{\sum^{n}_{j=1}F_{j}(T_{\mathrm{GBH}}(k))}$
$=$ $\sum_{i\in\mathcal{H}_{0}}\sum^{n}_{k=1}\frac{n\mathbb{E}[\mathbbm{1}\{\varphi_{i}(p_{i})\leq T_{\mathrm{GBH}}(k),R_{\mathrm{GBH}}(p_{i}\leftarrow 0)=k\}]}{\sum^{n}_{j=1}F_{j}(T_{\mathrm{GBH}}(k))}$
$\leq$ $\sum_{i\in\mathcal{H}_{0}}\sum^{n}_{k=1}\frac{nF_{i}(T_{\mathrm{GBH}}(k))\mathbb{E}[\mathbbm{1}\{R_{\mathrm{GBH}}(p_{i}\leftarrow 0)=k\}]}{\sum^{n}_{j=1}F_{j}(T_{\mathrm{GBH}}(k))}$
$=$ $\sum_{i\in\mathcal{H}_{0}}\sum_{k=1}^{n}\frac{nc_{i}h(T_{\mathrm{GBH}}(k))}{\sum_{j=1}^{n}c_{j}h(T_{\mathrm{GBH}}(k))}\mathbb{E}[\mathbbm{1}\{R_{\mathrm{GBH}}(p_{i}\leftarrow 0)=k\}]$
$=$ $n\frac{\sum_{i\in\mathcal{H}_{0}}c_{i}}{\sum_{j=1}^{n}c_{j}}\leq n,$
which completes the proof. ∎
B.4 Proof of Proposition 4
Write $T=T_{\mathrm{GBC}}$ for the ease of notation. First note that
$\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}}\frac{\mathbbm{1}\{\varphi_{i}(p_{i})\leq T\}}{1\vee\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(p_{j})\leq T\}}\right]$
$=$ $\sum_{i\in\mathcal{H}_{0}}\mathbb{E}\left[\frac{\mathbbm{1}\{\varphi_{i}(p_{i})\leq T\}}{1\vee\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(p_{j})\leq T\}}\frac{1+\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T\}}\right]$
$\leq$ $\alpha\sum_{i\in\mathcal{H}_{0}}\mathbb{E}\left[\frac{\mathbbm{1}\{\varphi_{i}(p_{i})\leq T\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T\}}\right].$
Hence, we only need to show that
$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}\left[\frac{\mathbbm{1}\{\varphi_{i}(p_{i})\leq T\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T\}}\right]\leq 1.$
One approach to prove FDR control is through the construction of a super-martingale and the use of the optional stopping time theorem. Here we employ an alternative argument based on the leave-one-out technique. Let $\tilde{p}_{i}=\min\{p_{i},1-p_{i}\}$ and $\mathbf{p}_{-i}=(p_{1},\dots,p_{i-1},\tilde{p}_{i},p_{i+1},\dots,p_{n})$. Define $T_{i}=T(\mathbf{p}_{-i})$, where we view $T$ as a function of the p-values. Notice that if $\varphi_{i}(p_{i})\leq T$, then we have
$\varphi_{i}(p_{i})\leq T\leq T_{\mathrm{up}}<\varphi_{i}(0.5),$
which implies that $p_{i}<0.5$ since $\varphi_{i}$ is increasing. Hence, if the $i$th hypothesis is rejected, then $p_{i}=\tilde{p}_{i}$. Thus, $\mathbbm{1}\{\varphi_{i}(p_{i})\leq T\}=\mathbbm{1}\{\varphi_{i}(p_{i})\leq T_{i}\}$, which further implies that
$\mathbb{E}\left[\frac{\mathbbm{1}\{\varphi_{i}(p_{i})\leq T\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T\}}\right]$ $=\mathbb{E}\left[\frac{\mathbbm{1}\{\varphi_{i}(p_{i})\leq T_{i}\}}{1+\sum_{j\neq i}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{i}\}}\right],$
where we use the fact that if $p_{i}<0.5$, then $\varphi_{i}(1-p_{i})\geq\varphi_{i}(0.5)>T_{\mathrm{up}}\geq T_{i}$. Let $\mathcal{F}_{i}$ be the sigma algebra generated by $\mathbf{p}_{-i}$. For $i\in\mathcal{H}_{0}$, we have
$\mathbb{E}\left[\frac{\mathbbm{1}\{\varphi_{i}(p_{i})\leq T\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T\}}\right]=$ $\mathbb{E}\left[\mathbb{E}\left[\frac{\mathbbm{1}\{\varphi_{i}(p_{i})\leq T_{i}\}}{1+\sum_{j\neq i}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{i}\}}\Bigg{|}\mathcal{F}_{i}\right]\right]$
$=$ $\mathbb{E}\left[\frac{1}{1+\sum_{j\neq i}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{i}\}}\mathbb{E}\left[\mathbbm{1}\{\varphi_{i}(p_{i})\leq T_{i}\}\lvert\mathcal{F}_{i}\right]\lvert\right]$
$\leq$ $\mathbb{E}\left[\frac{1}{1+\sum_{j\neq i}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{i}\}}\mathbb{E}\left[\mathbbm{1}\{\varphi_{i}(1-p_{i})\leq T_{i}\}\lvert\mathcal{F}_{i}\right]\lvert\right]$
$=$ $\mathbb{E}\left[\frac{\mathbbm{1}\{\varphi_{i}(1-p_{i})\leq T_{i}\}}{1+\sum_{j\neq i}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{i}\}}\right],$
where we use the assumption that $p_{i}$ satisfies Condition (8) to get the inequality. By Proposition  6, we have
$\frac{\mathbbm{1}\{\varphi_{i}(1-p_{i})\leq T_{i}\}}{1+\sum_{j\neq i}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{i}\}}=\frac{\mathbbm{1}\{\varphi_{i}(1-p_{i})\leq T_{i}\}}{1+\sum_{j\neq i}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{j}\}}=\frac{\mathbbm{1}\{\varphi_{i}(1-p_{i})\leq T_{i}\}}{\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{j}\}}.$
If $\varphi_{i}(1-p_{i})>T_{i}$, both sides are equal to 00. If $\varphi_{i}(1-p_{i})\leq T_{i}$, we claim that $\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{i}\}=\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{j}\}$. Indeed, if $\varphi_{j}(1-p_{j})>T_{i}$ but $\varphi_{j}(1-p_{j})\leq T_{j}$, then we have $T_{i}<T_{j}$. Hence, $\varphi_{i}(1-p_{i})\leq T_{i}<T_{j}$. By proposition 6, we have $T_{i}=T_{j}$, which contradicts with the assumption $T_{i}<T_{j}$. The other direction can be proved similarly.
Hence,
$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}\left[\frac{\mathbbm{1}\{\varphi_{i}(p_{i})\leq T\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T\}}\right]\leq\mathbb{E}\left[\frac{\sum_{i\in\mathcal{H}_{0}}\mathbbm{1}\{\varphi_{i}(1-p_{i})\leq T_{i}\}}{\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{j}\}}\right]\leq 1,$ (B.1)
which finishes the proof.
B.5 Proof of Proposition 5
Proof.
Proposition 5 is a special case of Proposition 6 by choosing $\varphi_{i}$ as the identity function for all $1\leq i\leq n$. ∎
B.6 Proof of Proposition 6
Proof.
Write $T=T_{\mathrm{GBC}}$ for the ease of notation. First, given a p-value vector $\mathbf{p}=(p_{1},\cdots,p_{n})$, recall that the threshold $T$ is defined as
$T=\max\left\{0<t\leq T_{\mathrm{up}}\colon\underbrace{\frac{1+\sum_{l=1}^{n}\mathbbm{1}\{\varphi_{l}(1-p_{l})\leq t\}}{\sum_{l=1}^{n}\mathbbm{1}\{\varphi_{l}(p_{l})\leq t\}}}_{\coloneqq g(\mathbf{p},t)}\leq\alpha\right\},$
where $T_{\mathrm{up}}$ satisfies $T_{\mathrm{up}}<\varphi_{l}(0.5)$ for all $l$.
Without loss of generality, let us assume $T_{i}\geq T_{j}$. By the assumption that $\max\{\varphi_{i}(1-p_{i}),\varphi_{j}(1-p_{j})\}\leq\max\{T_{i},T_{j}\}$, we have $\varphi_{i}(1-p_{i})\leq T_{i}$ and $\varphi_{j}(1-p_{j})\leq T_{i}$. Since $\varphi_{i}$ is an increasing function, we have $\varphi_{i}(1-p_{i})\leq T_{\mathrm{up}}<\varphi_{i}(0.5)$, which implies $0.51-p_{i}<$ 0.5. Thus $\varphi_{i}(p_{i})\geq\varphi_{i}(0.5)>T_{\mathrm{up}}\geq T_{i}$. The same discussion for $p_{j}$ leads to $\varphi_{j}(p_{j})>T_{i}$.
Denote $\tilde{p}_{i}=\min\{p_{i},1-p_{i}\}$ and $\mathbf{p}_{-i}=(p_{1},\cdots,p_{i-1},\tilde{p}_{i},p_{i+1},\cdots,p_{n})$ for all $i$. Consider the function
$g(\mathbf{p}_{-j},T_{i})=\frac{1+\sum_{l=1}^{n}\mathbbm{1}\{\varphi_{l}(1-p_{-j,l})\leq T_{i}\}}{\sum_{l=1}^{n}\mathbbm{1}\{\varphi_{l}(p_{-j,l})\leq T_{i}\}},$
where $p_{-j,l}$ is the $l$th entry of $p_{-j}$. For the denominator, we have
$\sum_{l=1}^{n}\mathbbm{1}\{\varphi_{l}(p_{-j,l})\leq T_{i}\}$
$=$ $\sum_{l=1}^{n}\mathbbm{1}\{\varphi_{l}(p_{-i,l})\leq T_{i}\}+\underbrace{\mathbbm{1}\{\varphi_{j}(p_{-j,j})\leq T_{i}\}}_{=1}+\underbrace{\mathbbm{1}\{\varphi_{i}(p_{-j,i})\leq T_{i}\}}_{=0}$
$\qquad-\underbrace{\mathbbm{1}\{\varphi_{j}(p_{-i,j})\leq T_{i}\}}_{=0}-\underbrace{\mathbbm{1}\{\varphi_{i}(p_{-i,i})\leq T_{i}\}}_{=1}$
$=$ $\sum_{l=1}^{n}\mathbbm{1}\{\varphi_{l}(p_{-i,l})\leq T_{i}\}.$
Similarly, for the numerator, we have
$\sum_{l=1}^{n}\mathbbm{1}\{\varphi_{l}(1-p_{-j,l})\leq T_{i}\}$
$=$ $\sum_{l=1}^{n}\mathbbm{1}\{\varphi_{l}(1-p_{-i,l})\leq T_{i}\}+\underbrace{\mathbbm{1}\{\varphi_{j}(1-p_{-j,j})\leq T_{i}\}}_{=0}$
$\qquad+\underbrace{\mathbbm{1}\{\varphi_{i}(1-p_{-j,i})\leq T_{i}\}}_{=1}-\underbrace{\mathbbm{1}\{\varphi_{j}(1-p_{-i,j})\leq T_{i}\}}_{=1}-\underbrace{\mathbbm{1}\{\varphi_{i}(1-p_{-i,i})\leq T_{i}\}}_{0}$
$=$ $\sum_{l=1}^{n}\mathbbm{1}\{\varphi_{l}(1-p_{-i,l})\leq T_{i}\}.$
Hence, $g(\mathbf{p}_{-j},T_{i})=g(\mathbf{p}_{-i},T_{i})\leq\alpha$. By the definition of $T_{j}$, we must have $T_{i}\leq T_{j}$. Similarly, we get $T_{j}\leq T_{i}$ and hence $T_{i}=T_{j}$. ∎
Appendix C Additional numerical results
C.1 Additional numerical results for fair multiple testing
[…]
In all settings, we assume that the p-values follow the uniform distribution on $[0,1]$ under the null. For the $g$th group, the p-value is assumed to follow $\mathrm{Unif}[0,u_{g}]$ under the alternatives. The parameter values for different settings are detailed in Table C.2.
Table C.2 displays the results for Setting F1. It can be seen that BC_Com fails to control the FDR for the third and fourth groups with the empirical FDR reaching 0.073 compared to the 5% target level. BC_Sep has an empirical FDR being 0.064, higher than the nominal level. The results for Setting F2 are presented in Table C.4. We note that BC_Com suffers from a severe FDR inflation with the empirical FDR being 0.312 at the 5% target level. In Setting F3, we raise the target FDR level to $20\%$. As seen from Table C.4, BC_Sep significantly inflates the overall FDR with the empirical FDR being 0.346. Throughout all settings, the e-BH based approach controls both the group-wise FDR and the overall FDR. Furthermore, eBH_Ada outperforms both eBH_1 and eBH_2 in terms of power.
C.2 Additional numerical results for structure adaptive multiple testing
The results for $a_{f}=0.5$ are presented in Figure C.1. When the signal is sparse, and the covariate is less informative, slight FDR inflation is observed in SABHA and the two versions of IHW. eBH_GBC has the highest power, followed by SABHA and the two versions of IHW. AdaPT, on the other hand, shows a power loss when compared to the BH procedure. However, as the covariate becomes more informative, all structure adaptive methods outperform the BH procedure, and eBH_GBC has the most true discoveries when the signal is sparse. Furthermore, when the signal becomes dense, eBH_GBC, AdaPT, and SABHA have similar performance in power.
Figure C.2 shows the results for $a_{f}=0$, i.e., the alternative p-value distribution is independent of the covariates. In this case, AdaPT performs the best, followed by eBG_GBC and SABHA, which dominate IHW and the BH procedure.
Table C.1: Parameter settings for the case of $G=4$. Here, $n_{g}$ represents the number of hypotheses for the $g$th group; $n_{ga}$ denotes the number of non-null hypotheses in the $g$th group. The parameter $u_{g}$ is the upper bound of the uniform distribution for the p-values under the alternatives for the $g$th group.
Setting Target FDR level $n_{1}$ $n_{1a}$ $p_{1}$ $n_{2}$ $n_{2a}$ $p_{2}$ $n_{3}$ $n_{3a}$ $p_{3}$ $n_{4}$ $n_{4a}$ $p_{4}$
Setting F1 0.05 100 20 $1.5\times 10^{-3}$ 100 20 $1.5\times 10^{-3}$ 1000 20 $1.5\times 10^{-3}$ 1000 20 $1.5\times 10^{-3}$
Setting F2 0.05 100 1 $5\times 10^{-4}$ 100 20 $2\times 10^{-3}$ 100 20 $3\times 10^{-3}$ 100 20 $4\times 10^{-3}$
Setting F3 0.2 50 2 $1\times 10^{-3}$ 100 2 $1\times 10^{-3}$ 50 4 $2\times 10^{-3}$ 100 4 $2\times 10^{-3}$
[…]
BC_Com 0.698 0.697 0.698 0.698 0.698 0.042 0.006 0.006 0.072 0.073
BC_Sep 0.579 0.906 0.893 0.245 0.273 0.064 0.048 0.045 0.025 0.03
eBH_1 0.022 0 0 0.045 0.045 0.006 0 0 0.005 0.006
eBH_2 0.051 0.051 0.051 0.051 0.05 0.004 0.003 0.002 0.005 0.006
eBH_Ada 0.22 0.313 0.313 0.118 0.137 0.021 0.016 0.015 0.012 0.015
Table C.1: Parameter settings for the case of $G=4$. Here, $n_{g}$ represents the number of hypotheses for the $g$th group; $n_{ga}$ denotes the number of non-null hypotheses in the $g$th group. The parameter $u_{g}$ is the upper bound of the uniform distribution for the p-values under the alternatives for the $g$th group.
Table C.2: FDR and power for Setting F1, where the nominal FDR level is 5%.
Table C.3: FDR and power for Setting F2, where the nominal FDR level is 5%.
[…]
BC_Com 0.93 0.947 0.94 0.932 0.919 0.044 0.312 0.032 0.03 0.031
BC_Sep 0.788 0 0.863 0.794 0.747 0.055 0 0.044 0.043 0.041
eBH_1 0.001 0 0.001 0.001 0.001 0 0 0 0 0
eBH_2 0.001 0 0.001 0.001 0.001 0 0 0 0 0
eBH_Ada 0.541 0 0.56 0.551 0.538 0.032 0 0.029 0.03 0.03
[…]
BC_Com 0.955 0.967 0.968 0.948 0.95 0.174 0.147 0.236 0.095 0.151
BC_Sep 0.393 0.149 0.139 0.528 0.507 0.346 0.1 0.094 0.173 0.171
eBH_1 0.002 0.001 0.003 0.001 0.003 0.002 0.001 0.002 0 0.002
eBH_2 0.007 0.005 0.004 0.008 0.008 0.004 0.003 0.003 0.003 0.003
eBH_Ada 0.038 0.012 0.006 0.053 0.051 0.021 0.008 0.004 0.016 0.018
Table C.3: FDR and power for Setting F2, where the nominal FDR level is 5%.
Table C.4: FDR and power for Setting F3, where the nominal FDR level is 20%.
Figure C.1: Empirical FDR and power for $a_{f}=0.5$.
Figure C.2: Empirical FDR and power for $a_{f}=0$.
wzAYj8cwTZPwPH9/sVg8PXweDAauqqr2cDjEer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v2

@@
-E-values, Multiple Testing and Beyond
+A General Framework for Multiple Testing via E-value Aggregation and Data-Dependent Weighting
 Abstract
-We discover a connection between the Benjamini-Hochberg (BH) procedure and the recently proposed e-BH procedure Wang and Ramdas (2022) with a suitably defined set of e-values.
-This insight extends to a generalized version of the BH procedure and the model-free multiple testing procedure in Barber and Candès (2015) (BC) with a general form of rejection rules.
-The connection provides an effective way of developing new multiple testing procedures by aggregating or assembling e-values resulting from the BH and BC procedures and their use in different subsets of the data.
-In particular, we propose new multiple testing methodologies in three applications, including a hybrid approach that integrates the BH and BC procedures, a multiple testing procedure aimed at ensuring a new notion of fairness by controlling both the group-wise and overall false discovery rates (FDR), and a structure adaptive multiple testing procedure that can incorporate external covariate information to boost detection power.
-One notable feature of the proposed methods is that we use a data-dependent approach for assigning weights to e-values, significantly enhancing the efficiency of the resulting e-BH procedure.
-The construction of the weights is non-trivial and is motivated by the leave-one-out analysis for the BH and BC procedures.
-In theory, we prove that the proposed e-BH procedures with data-dependent weights in the three applications ensure finite sample FDR control.
-Furthermore, we demonstrate the efficiency of the proposed methods through numerical studies in the three applications.
-Keywords: Benjamini-Hochberg procedure, Cross-fitting, E-values, False discovery rate, Leave-one-out analysis, Multiple testing
+Motivated by recent findings in Li and Zhang (2025), which established an equivalence between certain p-value-based multiple testing procedures and the e-Benjamini-Hochberg procedure (Wang and Ramdas, 2022), we introduce a general framework for constructing novel multiple testing methods through the aggregation and combination of e-values.
+Specifically, we propose methodologies for three distinct scenarios: (i) assembly of e-values obtained from different subsets of data, simultaneously controlling group-wise and overall false discovery rates; (ii) aggregation of e-values derived from different procedures or the same procedure employing different test statistics; and (iii) adaptive multiple testing methods that incorporate external structural information to enhance statistical power.
+A notable feature of our approach is the use of data-dependent weighting of e-values, significantly improving the efficiency of the resulting e-Benjamini-Hochberg procedures.
+The construction of these weights is non-trivial and inspired by leave-one-out analysis, a widely utilized technique for proving false discovery rate control in p-value-based methodologies.
+We theoretically establish that the proposed e-Benjamini-Hochberg procedures, when equipped with data-dependent weights, guarantee finite-sample false discovery rate control across all three considered applications.
+Additionally, numerical studies illustrate the efficacy and advantages of the proposed methods within each application scenario.
+Keywords: Cross-fitting, E-values, False discovery rate, Leave-one-out analysis, Multiple testing
 1 Introduction
-When working with high-dimensional data in modern scientific fields, the problem of multiple testing often arises when we explore a vast number of hypotheses with the goal of detecting signals while also controlling some error measures, such as the false discovery rate (FDR).
-The Benjamini-Hochberg (BH) procedure (Benjamini and Hochberg, 1995) is perhaps the most widely used FDR-controlling procedure that rejects a hypothesis whenever its p-value is less than or equal to an adaptive rejection threshold determined by the whole set of p-values.
-Barber and Candès (2015) proposed a model-free FDR-controlling (BC) procedure that estimates the number of false rejections by leveraging the symmetry of p-values under the null and compares each p-value with an adaptive threshold.
-More recently, there is a growing literature on utilizing e-values for statistical inference under different contexts, see, e.g., (Grünwald et al., 2020; Shafer, 2021; Vovk and Wang, 2021; Xu et al., 2021; Ignatiadis et al., 2022; Dunn et al., 2023; Xu and Ramdas, 2023).
-In particular, Wang and Ramdas (2022) proposed a multiple testing procedure (named e-BH procedure) by applying the BH procedure to e-values, which was shown to control the FDR even when the e-values exhibit arbitrary dependence.
-In this work, we establish a connection between the BH and e-BH procedures with a suitably defined set of e-values, proving that they yield identical rejection sets.
-This connection extends to a generalized version of the BH and BC procedures, which can have a more general form for the rejection rule.
-Based on these findings, we propose new multiple testing procedures by aggregating e-values from the BH and BC procedures or assembling e-values from different subsets of the data.
-Specifically, we consider three concrete applications, which we will illustrate in detail below.
-Our first application concerns the development of a robust and efficient multiple testing procedure by leveraging the strength of the BH and BC procedures.
-The BH and BC procedures employ different strategies to estimate the number of false rejections, leading to distinct performances depending on the underlying signal density and strength.
-Empirical results in the literature suggest that neither one dominates the other.
-For instance, the BH procedure is usually more powerful when the underlying signal is sparse, while the BC procedure excels in the case of dense signal (Arias-Castro and Chen, 2017).
-It is often impossible to know in advance which method will perform better in real-world applications.
-Reporting results from the better-performing method can lead to an inflated FDR and is considered as a type of data snooping.
-To develop a more efficient procedure with guaranteed FDR control, we propose a hybrid method that utilizes e-values as a bridge to integrate the testing results from the BH and BC procedures.
-Specifically, we construct a set of e-values by properly weighting the e-values from both the BH and BC procedures.
-We then use the constructed e-values as the input for the e-BH procedure.
-We show that the resulting hybrid procedure controls the FDR and can significantly improve the worse-performing method in finite sample.
-Our second application is concerned with a setting where one has to make (high-stakes) decisions by testing $n$ hypotheses that are partitioned into $G$ groups according to some (protected) attributes.
-Let us consider the example of a bank’s loan approval process, where customers are grouped by gender or race.
-The null hypothesis in this case is that the loan should be approved for the customer.
-The bank faces the challenge of controlling the overall FDR, which means that the bank should not falsely decline too many customers’ applications, as it would reduce the loan volume and decrease income.
-Additionally, the bank also needs to control the group-wise FDR to ensure that certain racial or gender groups are not unfairly denied loans.
-If the group-wise FDR inflates, customers who believe they are unfairly denied loans may take their business elsewhere.
-Therefore, the bank aims to maintain control over the FDR for each group while also controlling the overall FDR across all hypotheses.
-Applying an FDR-controlling procedure to the whole set of hypotheses, ignoring the group structure will fail to control the FDR for individual groups.
-On the other hand, performing multiple testing for each group separately at level $\alpha$ does not ensure that the overall FDR is controlled at the same level $\alpha$.
-A naive procedure to achieve error control at both the group and overall levels is to conduct multiple testing for each group separately at the reduced level $\alpha/G$, which turns out to be too conservative, as seen in our (unreported) numerical studies.
-To address this challenge, we propose a fair multiple testing procedure that uses e-values as a bridge to assemble the results from different groups, controlling the FDR within each group and the overall FDR simultaneously.
-Specifically, we implement the BC procedure for each group, followed by constructing e-values by assembling e-values from different groups with proper weights, and using the constructed e-values as the input for the e-BH procedure.
-We show that the resulting procedure simultaneously controls the FDR within each group and the overall FDR in finite sample.
-As the final application, we consider the problem of multiple testing with external structural information in the form of covariates, which has received much recent attention, as leveraging auxiliary information can enhance the power and interpretability of multiple testing results in many scientific applications.
-For instance, in differential expression analysis of RNA-seq data, which tests for differences in the mean expression of the genes between conditions, the sum of read counts per gene across all samples could be the auxiliary data since it is informative of the statistical power.
-In differential abundance analysis of microbiome sequencing data, which tests for differences in the mean abundance of the detected bacterial species between conditions, the genetic divergence among species is important auxiliary information since closely related species usually have similar physical characteristics and tend to covary with the condition of interest.
-A growing list of works has reflected the importance of this research direction in recent years, for instance, (Genovese et al., 2006; Hu et al., 2010; Sun et al., 2015; Ignatiadis et al., 2016; Boca and Leek, 2018; Lei and Fithian, 2018; Li and Barber, 2019; Ignatiadis and Huber, 2021; Yun et al., 2022; Zhang and Chen, 2022; Zhao and Zhou, 2023).
-However, these existing works suffer from different limitations.
-For example, the local FDR-based method (Sun et al., 2015; Cao et al., 2022) lacks finite sample FDR control and can only guarantee FDR control asymptotically.
+In modern scientific research involving high-dimensional data, multiple testing frequently arises as a fundamental challenge.
+This occurs when simultaneously evaluating a large number of hypotheses to identify significant signals, necessitating careful control of error rates such as the false discovery rate (FDR) to ensure statistical validity.
+The Benjamini-Hochberg (BH) procedure (Benjamini and Hochberg, 1995) and the Barber-Candès (BC) procedure (Barber and Candès, 2015) are the most commonly employed methods for controlling the FDR using p-values.
+Recently, there has been increasing interest in employing e-values for FDR control; see, for example, Ignatiadis et al.
+(2024a, b); Xu and Ramdas (2024).
+In particular, Wang and Ramdas (2022) proposed a multiple testing approach called the e-Benjamini-Hochberg (e-BH) procedure, which applies the BH procedure directly to e-values.
+They demonstrated that the e-BH procedure controls the FDR even when e-values exhibit arbitrary dependence structures.
+Compared to p-values, which are defined through tail probabilities, e-values, defined via expectation, offer greater flexibility for combining multiple e-values to obtain new e-values (Vovk and Wang, 2021).
+For comprehensive reviews of methods based on e-values, we refer readers to Ramdas and Wang (2024).
+In recent work, Li and Zhang (2025) introduced a unified framework for multiple testing procedures based on p-values, which includes the BH and BC procedures as special cases.
+The authors established the equivalence between these p-value-based methods and the e-BH procedure when appropriate sets of e-values are utilized.
+Here, equivalence means that these methods yield identical rejection sets.
+Motivated by these findings, we propose new multiple testing procedures that aggregate e-values derived from different methods or the same method with different test statistics, or combine e-values obtained from different subsets of data.
+Specifically, we explore three concrete scenarios, which are detailed in the subsequent sections.
+In our first scenario, we consider the setting with $L$ sets of e-values derived from $L$ distinct datasets.
+Our goal is to aggregate these e-values into a single vector to incorporate all group-level information.
+We propose a procedure designed to control both the overall FDR and the group-wise FDR simultaneously.
+A key application involves high-stakes decision-making, where we test $n$ hypotheses partitioned into $G$ groups according to certain protected attributes.
+In a loan-approval system, for example, customers may be grouped by gender or race; the null hypothesis states that a given loan should be approved.
+Management must control both the overall FDR—so that too few loans are rejected—and the group-wise FDR—so that no group is unfairly treated.
+Conventional solutions fail: applying the FDR-controlling method to all $n$ tests fails to control FDR for certain groups; testing each group at level $\alpha$ fails to control overall FDR; and a Bonferroni adjustment to $\alpha/G$ is too conservative.
+To address this challenge, we propose a multiple-testing procedure that uses e-values as a bridge to combine the results from different groups, controlling the FDR within each group and the overall FDR simultaneously.
+Specifically, we apply the BC procedure to each group, then assemble the resulting e-values with appropriate weights to form a unified e-value vector, which we pass to the e-BH procedure.
+We show that the resulting method simultaneously controls the FDR within each group and the overall FDR in finite samples.
+In our second scenario, we consider a setting with $L$ distinct sets of e-values, which we aim to aggregate into a single e-value vector to incorporate all available information.
+By effectively combining e-values from diverse sources, the proposed approach enables the integration of multiple results while rigorously controlling the overall FDR.
+This general scenario has several specific applications.
+The first specific application is the development of a robust and efficient knockoff method capable of accommodating various underlying relationships between the response variable and predictors.
+In the knockoff methodology (Barber and Candès, 2015), several test statistics can be employed: Lasso-based methods excel for (near-)linear models, whereas random forest-based methods suit nonlinear settings (Candes et al., 2018).
+Because the true form of the dependence is rarely known, we develop a robust and efficient knockoff procedure that can leverage the strengths of different test statistics simultaneously.
+Specifically, we construct a unified set of e-values by aggregating the e-values derived from knockoff methods based on various test statistics.
+Subsequently, we input these aggregated e-values into the e-BH procedure.
+We demonstrate that the resulting hybrid knockoff procedure effectively controls the FDR and maintains robust performance regardless of the true underlying relationship between the response variable and predictors.
+Our second application involves developing a robust and efficient multiple-testing procedure by combining the strengths of the BH and BC procedures.
+The BH and BC procedures use different strategies for estimating the number of false rejections, leading to distinct performance characteristics depending on signal density and strength.
+In real-world applications, reporting results from the better-performing method can lead to inflated FDR and is considered a form of data snooping.
+To address this issue, we propose a hybrid approach that employs e-values as a bridge to integrate results from both the BH and BC procedures.
+To this end, we construct a unified set of e-values by suitably weighting the e-values derived from the BH and BC procedures.
+These combined e-values are then input into the e-BH procedure.
+We show that the resulting hybrid procedure maintains rigorous FDR control and can significantly improve performance relative to the weaker individual method in finite-sample settings.
+In the final scenario, we consider the problem of multiple testing with external structural information in the form of covariates, which has received significant recent attention, as leveraging auxiliary information can enhance the power and interpretability of multiple-testing results in many scientific applications.
+Typical covariates include (i) total read counts in RNA-seq, which modulate gene-level power, and (ii) phylogenetic distances in microbiome studies, where related species share abundance patterns.
+A growing list of works has reflected the importance of this research direction in recent years—for instance, Hu et al.
+(2010); Ignatiadis et al.
+(2016); Lei and Fithian (2018); Li and Barber (2019); Ignatiadis and Huber (2021); Zhang and Chen (2022); Zhao and Zhou (2024).
+However, these existing works suffer from various limitations.
+For example, the local FDR-based methods (Sun et al., 2015; Cao et al., 2022) lack finite-sample FDR control and only guarantee FDR control asymptotically.
 The weighted BH methods (Ignatiadis and Huber, 2021; Li and Barber, 2019) lead to suboptimal power, as observed in our numerical studies.
-To address these drawbacks, we propose a powerful multiple testing procedure that incorporates auxiliary information with guaranteed FDR control in finite samples.
-We randomly split samples into several groups and use the cross-fitting approach to estimate the rejection function in each group using all samples in other groups.
-Then, we implement the generalized BC procedure for each group.
-We assemble the BC procedure induced e-values from different groups with proper weights.
-The assembled e-values are used as the input for e-BH procedure.
-We show that the proposed procedure controls the FDR at the desired level in finite sample.
-Our approach involves a data-dependent method for weighting e-values when aggregating them from the BH and BC procedures (the first application) or assembling them from different subsets of the data (the second and third applications).
+To address these drawbacks, we propose a powerful multiple-testing procedure that incorporates auxiliary information while guaranteeing FDR control in finite samples.
+Specifically, we randomly split the data into several disjoint groups and use a cross-fitting approach (Ignatiadis and Huber, 2021) to estimate the rejection function for each group using all samples in the other groups.
+We then apply the flexible BC procedure (Li and Zhang, 2025) within each group, assemble the resulting group-wise BC e-values with appropriate weights, and feed the combined vector into the e-BH procedure.
+We show that the proposed method controls the FDR at the desired level in finite samples and achieves competitive power relative to state-of-the-art methods.
+Our approach involves a data-dependent method for weighting e-values when aggregating them from the BH and BC procedures (the second scenario) or assembling them from different subsets of the data (the first and third scenarios).
 It is important to note that our weighting method differs from the “boosting factor” proposed by Wang and Ramdas (2022), which involves multiplying each e-value by a factor to boost them up before applying the e-BH procedure.
-In our approach, we construct weights that ensure the weighted e-values satisfying Condition (3) below.
-This condition is sufficient to maintain FDR control of the e-BH procedure at the desired level.
-Our numerical findings show that implementing the e-BH procedure with data-dependent weights improves its efficiency across all applications compared to the unweighted e-BH procedure.
-The rest of the paper is organized as follows.
-In Section 2, we provide a brief review of the BH, BC, and e-BH procedures and establish the equivalence between the BH/BC and the e-BH procedures.
-In Section 3, we extend this equivalence to a generalized version of the BH and BC procedures.
-We present new multiple testing methodologies in three different applications, including a hybrid approach that integrates the BH and BC procedures, a multiple testing procedure that controls both the group-wise and overall FDRs, and a structure adaptive multiple testing procedure, in Sections 4.1, 4.2, and 4.3, respectively.
-Section 5 concludes.
-The appendix contains additional numerical results and all proofs of the main results.
+It is also worth mentioning the connection to the work of Ignatiadis et al.
+(2024b), where the authors used e-values as unnormalized weights for p-values in order to improve the testing power.
+The authors constructed these e-values using Basu’s theorem, which makes their weights independent of the p-values.
+In contrast, our weights are dependent on the e-values.
+The construction of our weights is motivated by the leave-one-out analysis for the BH and BC procedures.
+It ensures that the weighted e-values satisfy Condition (2) below, which is sufficient for the corresponding e-BH procedure to maintain FDR control at the desired level.
+Our numerical findings show that implementing the e-BH procedure with data-dependent weights improves its efficiency across all applications compared to the unweighted version.
+The remainder of the paper is organized as follows.
+Section 2 briefly reviews several multiple testing procedures and their relationship to the e-BH procedure.
+Sections 3-5 respectively present three distinct scenarios along with our newly proposed multiple testing methodologies tailored to different applications: (i) a multiple testing procedure controlling both group-wise and overall FDR, (ii) a hybrid knockoff method and a hybrid approach that integrates the BH and BC procedures, and (iii) a structure-adaptive multiple testing procedure.
+Section 6 provides concluding remarks.
+The Supplement includes additional numerical results as well as complete proofs of all main theoretical results.
 2 Preliminaries
-To begin with, we provide a brief overview of the BH, BC, and e-BH procedures, and establish the equivalence between the BH (BC) procedure and the e-BH procedure with a suitably defined set of e-values.
-This equivalence appears to be a new finding that has not been explicitly stated in the previous literature.
-We are interested in testing $n$ hypotheses $\{H_{1},\dots,H_{n}\}$ simultaneously.
-Let $\bm{\theta}=\{\theta_{1},\dots,\theta_{n}\}\in\{0,1\}^{n}$ indicate the underlying truth of each hypothesis, where $\theta_{i}=0$ if $H_{i}$ is under the null and $\theta_{i}=1$ otherwise.
-Denote by $\bm{\delta}=(\delta_{1},\dots,\delta_{n})\in\{0,1\}^{n}$ a decision rule for the $n$ hypotheses, where we reject (accept) the $i$th hypothesis if $\delta_{i}=1$ ($\delta_{i}=0$).
-The FDR for the decision rule $\bm{\delta}$ is defined as the expectation of the false discovery proportion (FDP), i.e.,
-$\mathrm{FDR}(\bm{\delta})=\mathbb{E}[\mathrm{FDP}(\bm{\delta})],\quad\mathrm{FDP}(\bm{\delta})=\frac{\sum_{i=1}^{n}(1-\theta_{i})\delta_{i}}{1\vee\sum_{i=1}^{n}\delta_{i}},$
-where $a\vee b=\max\{a,b\}$. The goal of an FDR controlling procedure is to ensure that the FDR is bounded from above by a pre-specified number $\alpha\in(0,1)$.
-2.1 BH procedure
-The Benjamini-Hochberg (BH) procedure (Benjamini and Hochberg, 1995) is perhaps the most widely used FDR-controlling method.
-To describe the procedure, suppose we observe a p-value $p_{i}$ for each $H_{i}$.
-Sort the p-values in ascending order as $p_{(1)}\leq\dots\leq p_{(n)}$ and let $\hat{k}=\max\left\{i\colon p_{(i)}\leq(\alpha i)/n\right\}$.
-The BH procedure rejects all hypotheses $H_{(i)}$ with $i\leq\hat{k}$, where $H_{(i)}$ is the hypothesis associated with $p_{(i)}$.
-This procedure is equivalent to rejecting all $H_{i}$ with $p_{i}\leq T_{\mathrm{BH}}$, where $T_{\mathrm{BH}}$ is defined as
-$T_{\mathrm{BH}}=\sup\left\{0<t\leq 1\colon\frac{nt}{1\vee R(t)}\leq\alpha\right\},$ (1)
-with $R(t)=\sum_{i=1}^{n}\mathbbm{1}\{p_{i}\leq t\}$ being the number of rejections given the threshold $t$, and $\mathbbm{1}\{A\}$ denoting the indicator function associated with a set $A$.
-It is well known that if the null p-values are mutually independent and super-uniform, and are independent of the alternative p-values, the BH procedure at level $\alpha$ controls the FDR at the level $\alpha n_{0}/n\leq\alpha$, where $n_{0}$ is the number of hypotheses under the null.
-2.2 BC procedure
-In a seminal paper by (Barber and Candès, 2015), the authors proposed a model-free multiple testing procedure (BC procedure hereafter) that exploits the symmetry of the null p-values (or test statistics) to estimate the number of false rejections.
-More precisely, the BC procedure specifies a data-dependent threshold, denoted by $T_{\mathrm{BC}}$, which is determined as follows:
-$T_{\mathrm{BC}}=\sup\left\{0<t<0.5\colon\frac{1+\sum^{n}_{i=1}\mathbbm{1}\{p_{i}\geq 1-t\}}{1\vee R(t)}\leq\alpha\right\},$ (2)
-and it rejects all $H_{i}$ with $p_{i}\leq T_{\mathrm{BC}}$. The BC procedure has been shown to provide finite sample FDR control under suitable assumptions (Barber and Candès, 2015).
-2.3 E-values and e-BH procedure
-A non-negative random variable $E$ is called an e-value if $\mathbb{E}[E]\leq 1$ under the null hypothesis.
-Suppose we observe $n$ e-values $e_{1},\dots,e_{n}$ corresponding to the hypotheses $H_{1},\dots,H_{n}$.
-The $\alpha$-level e-BH procedure involves sorting the e-values in decreasing order as $e_{(1)}\geq\dots\geq e_{(n)}$ and rejecting the hypotheses associated with the $\hat{k}$ largest e-values, where $\hat{k}\coloneqq\max\left\{1\leq i\leq n\colon e_{(i)}\geq n/(i\alpha)\right\}$.
-Proposition 1 (Theorem 2 of Wang and Ramdas (2022)).
-Suppose the e-values satisfy
-$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[e_{i}]\leq n,$ (3)
-where $\mathcal{H}_{0}=\{1\leq i\leq n:\theta_{i}=0\}$. Then, the $\alpha$-level e-BH procedure controls the FDR at the level $\alpha$, regardless of the dependence structure among the e-values.
-2.4 Connection between the BH and e-BH procedures
-To see the connection between the BH and e-BH procedures, we define the e-value associated with $H_{i}$ to be
-$e_{i}=T_{\mathrm{BH}}^{-1}\mathbbm{1}\{p_{i}\leq T_{\mathrm{BH}}\},$ (4)
-where $T_{\mathrm{BH}}$ is given in (1).
-By Lemmas 3-4 in Storey et al.
-(2004), $\sum_{i\in\mathcal{H}_{0}}\mathbbm{1}\{p_{i}\leq t\}/t$ for $0<t\leq 1$ is martingale with time running backwards with respect to the filtration $\mathcal{F}_{t}=\sigma(\mathbbm{1}\{p_{i}\leq s\}:t\leq s\leq 1,i=1,2,\dots,n)$, and $T_{\mathrm{BH}}$ is a stopping time with respect to $\mathcal{F}_{t}$, where $\mathcal{F}_{t}$ is the sigma field generated by $\mathbbm{1}\{p_{i}\leq s\}$ for $t\leq s\leq 1.$ By the optional stopping time theorem, we have
-$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[e_{i}]=\mathbb{E}\left[\frac{\sum_{i\in\mathcal{H}_{0}}\mathbbm{1}\{p_{i}\leq T_{\mathrm{BH}}\}}{T_{\mathrm{BH}}}\right]=\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}}\mathbbm{1}\{p_{i}\leq 1\}\right]=n_{0},$
-which implies that the e-values defined by equation (4) satisfy (3).
-Thus by Proposition 1, the corresponding e-BH procedure controls the FDR at the desired level.
-Moreover, we claim that the e-BH procedure based on the e-values defined in (4) is equivalent to the BH procedure based on $\{p_{1},\dots,p_{n}\}$ in the sense that they produce the same set of rejections.
-Theorem 1.
-Let $\mathcal{S}_{\mathrm{BH}}$ be the set of rejections obtained through the BH procedure at the FDR level $\alpha$, and let $\mathcal{S}_{\mathrm{eBH}}$ represent the set of rejections obtained from the e-BH procedure at the same FDR level $\alpha$, with the e-values defined in (4).
-Then we have $\mathcal{S}_{\mathrm{BH}}=\mathcal{S}_{\mathrm{eBH}}$.
-2.5 Connection between the BC and e-BH procedures
-As noted in a recent work by Ren and Barber (2022), the BC procedure is equivalent to the e-BH procedure based on the following e-values:
-$e_{i}=\frac{n\mathbbm{1}\{p_{i}\leq T_{\mathrm{BC}}\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC}}\}},$
-where $T_{\mathrm{BC}}$ is the threshold defined in (2).
-3 Generalized BH and BC procedures
-In this section, we extend the BH and BC procedures to allow a general class of rejection rules.
-This generalization enables the testing procedure to utilize cross-sectional information among the p-values and external structural information for each hypothesis, which often results in a higher multiple testing power.
-We shall establish the connection between these generalized procedures and the corresponding e-BH procedures.
-3.1 Generalized BH procedure
-We generalize the BH procedure to allow the rejection rule to take the form of $\varphi_{i}(p_{i})\leq t$, where $\varphi_{i}$ is a strictly increasing function and can differ for each $i$.
-Let us define $F_{i}(t)=\varphi_{i}^{-1}(t)$ as the inverse function of $\varphi_{i}$, and consider the rejection threshold
-$T_{\text{GBH}}=\sup\left\{0<t\leq 1:\frac{\sum^{n}_{i=1}F_{i}(t)}{1\vee R(t)}\leq\alpha\right\},$ (5)
-where $R(t)=\sum^{n}_{i=1}\mathbbm{1}\{\varphi_{i}(p_{i})\leq t\}$.
-The generalized BH (GBH) procedure rejects $H_{i}$ whenever $\varphi_{i}(p_{i})\leq T_{\mathrm{GBH}}$.
-Similar to the BH procedure, the GBH procedure can be equivalently implemented in the following way.
-Define $\bar{F}(t)=n^{-1}\sum^{n}_{i=1}F_{i}(t)$ which is strictly increasing in $t$.
-We sort $q_{i}=\varphi_{i}(p_{i})$ in an ascending order, i.e., $q_{(1)}\leq\dots\leq q_{(n)}$ and find the largest $k$, represented as $\hat{k}$, for which $q_{(k)}\leq\bar{F}^{-1}\left(\frac{\alpha k}{n}\right).$ We then reject $H_{(i)}$ for all $i\leq\hat{k}$.
-The following proposition states that, under appropriate conditions in the form of $F_{i}$, the GBH procedure ensures FDR control.
-Proposition 2.
-Suppose that the null p-values $\{p_{i}\}_{i\in\mathcal{H}_{0}}$ are mutually independent and super-uniform, and are independent of the alternative p-values $\{p_{i}\}_{i\notin\mathcal{H}_{0}}$.
-If $F_{i}(t)=c_{i}h(t)$ where $c_{i}$ is some constant and $h$ is a strictly increasing function of $t$, the GBH procedure controls the FDR at level $\alpha$.
-As illustrated in the following examples, the GBH procedure aligns with several methods in the literature of structure adaptive multiple testing.
-Example 1.
-Let $\varphi_{i}(p)=p/\omega_{i}$, where $\omega_{i}$ denotes the weight for the $i$th hypothesis with $\omega_{i}>0$ and $\sum_{i=1}^{n}\omega_{i}=n$.
-The GBH procedure associated with this choice of $\varphi_{i}$ corresponds to the weighted BH procedure first introduced by Genovese et al.
-(2006).
-In this case, $F_{i}(t)=w_{i}t$ and the rejection threshold can be expressed as
-$T_{\text{GBH}}=\sup\left\{0<t\leq 1\colon\frac{\sum_{i=1}^{n}w_{i}t}{1\vee R(t)}\leq\alpha\right\}=\sup\left\{0<t\leq 1\colon\frac{nt}{1\vee R(t)}\leq\alpha\right\},$
-where $R(t)=\sum_{i=1}^{n}\mathbbm{1}\{p_{i}/w_{i}\leq t\}$.
-Example 2.
-Another choice of $\varphi_{i}$ is the local FDR (Efron, 2005; Sun and Cai, 2007) under a two-group mixture model.
-Specifically, we assume that the p-values are independently generated from the mixture density given by $\pi_{i}+(1-\pi_{i})(1-\kappa)p^{-\kappa}$, where $\pi_{i}\in(0,1)$ is the mixing proportion and $\kappa\in(0,1)$ controls the shape of the p-value density under the alternative.
-The rejection rule based on the local FDR under this model is given by
-$\varphi_{i}(p)=\frac{\pi_{i}}{\pi_{i}+(1-\pi_{i})(1-\kappa)p^{-\kappa}}\leq t,$ (6)
-which is the probability that the $i$th hypothesis is under the null given the observed p-value being $p$. It follows that
-$F_{i}(t)=\varphi_{i}^{-1}(t)=\left(\frac{\pi_{i}(1-t)}{(1-\pi_{i})t(1-\kappa)}\right)^{-1/\kappa}=c_{i}h(t),$
-where $c_{i}=\left\{\pi_{i}/\bigl{(}(1-\pi_{i})(1-\kappa)\bigr{)}\right\}^{-1/\kappa}$ and $h(t)=\left\{(1-t)/t\right\}^{-1/\kappa}$.
-3.2 Connection between the GBH and e-BH procedures
-Analogous to the BH procedure, we show that the GBH procedure is equivalent to the e-BH procedure applied to the following e-values:
-$e_{i}=\frac{n\mathbbm{1}\{\varphi_{i}(p_{i})\leq T_{\text{GBH}}\}}{\sum^{n}_{i=1}F_{i}(T_{\mathrm{GBH}})},$ (7)
-where $T_{\text{GBH}}$ is defined in (5). By the leave-one-out argument, we prove the following result.
-Proposition 3.
-Under the assumptions in Proposition 2, the e-values $e_{1},\dots,e_{n}$ defined in (7) satisfy Condition (3). As a consequence, the corresponding e-BH procedure controls the FDR at the desired level.
-Additionally, we can prove that the e-BH procedure and the GBH procedure deliver the same set of rejections.
-Theorem 2.
-Let $\mathcal{S}_{\mathrm{GBH}}$ be the set of rejections obtained through the GBH procedure at the FDR level $\alpha$, and let $\mathcal{S}_{\mathrm{eBH}}$ represent the set of rejections obtained from the e-BH procedure at the same FDR level $\alpha$, with the e-values defined in (7).
-Then we have $\mathcal{S}_{\mathrm{GBH}}=\mathcal{S}_{\mathrm{eBH}}$.
-3.3 Generalized BC procedure
-In this section, we generalize the BC procedure with the rejection rule given by $\varphi_{i}(p_{i})\leq t$. We assume that the null p-value satisfies the following condition:
-$P(p_{i}\leq a)\leq P(p_{i}\geq 1-a)=P(1-p_{i}\leq a),\quad\text{ for all }0\leq a\leq 0.5.$ (8)
-Condition (8) is weaker than the mirror conservativeness in Lei and Fithian (2018), and it can be shown that super-uniformity implies (8). Indeed,
-$P(1-p_{i}\leq a)\geq 1-P(p_{i}\leq 1-a)\geq 1-(1-a)=a\geq P(p_{i}\leq a).$
-Assume that $\varphi_{i}$ is an increasing and continuous function, and define $F_{i}(x)=\sup\{0\leq p\leq 1\colon\varphi_{i}(p)\leq x\}$. We claim that
-$P(\varphi_{i}(p_{i})\leq b)=P(p_{i}\leq F_{i}(b)).$
-To see this, consider two cases.
-If $\varphi_{i}(p_{i})\leq b$, by the definition of $F_{i}(b)$, we have $p_{i}\leq F_{i}(b)$.
-On the other hand, if $p_{i}\leq F_{i}(b)$, then $\varphi_{i}(p_{i})\leq\varphi_{i}(F_{i}(b))=\lim_{p\uparrow F_{i}(b)}\varphi_{i}(p)\leq b$, where we use the fact that $\varphi_{i}$ is increasing to get the two inequalities, and the equality is due to the continuity of $\varphi_{i}$.
-Therefore, the above claim together with equation (8) implies that
-$P(\varphi_{i}(p_{i})\leq b)=P(p_{i}\leq F_{i}(b))\leq P(1-p_{i}\leq F_{i}(b))=P(\varphi_{i}(1-p_{i})\leq b),$ (9)
-for all $\varphi_{i}(0)\leq b\leq\varphi_{i}(0.5)$. Hence, we have
-$\frac{\sum_{i\in\mathcal{H}_{0}}\mathbbm{1}\{\varphi_{i}(p_{i})\leq t\}}{1\vee\sum_{i=1}^{n}\mathbbm{1}\{\varphi_{i}(p_{i})\leq t\}}$ $\approx\frac{\sum_{i\in\mathcal{H}_{0}}P(\varphi_{i}(p_{i})\leq t\})}{1\vee\sum_{i=1}^{n}\mathbbm{1}\{\varphi_{i}(p_{i})\leq t\}}$
-$\leq\frac{\sum_{i\in\mathcal{H}_{0}}P(\varphi_{i}(1-p_{i})\leq t\})}{1\vee\sum_{i=1}^{n}\mathbbm{1}\{\varphi_{i}(p_{i})\leq t\}}$
-$\approx\frac{\sum_{i\in\mathcal{H}_{0}}\mathbbm{1}\{\varphi_{i}(1-p_{i})\leq t\}}{1\vee\sum_{i=1}^{n}\mathbbm{1}\{\varphi_{i}(p_{i})\leq t\}}$
-$\leq\frac{1+\sum_{i=1}^{n}\mathbbm{1}\{\varphi_{i}(1-p_{i})\leq t\}}{1\vee\sum_{i=1}^{n}\mathbbm{1}\{\varphi_{i}(p_{i})\leq t\}},$
-where the last term can be viewed as a conservative estimate of the FDP. Motivated by this observation, we define the threshold for the GBC procedure as
-$T_{\text{GBC}}=\sup\left\{0<t\leq T_{\mathrm{up}}\colon\frac{1+\sum_{i=1}^{n}\mathbbm{1}\{\varphi_{i}(1-p_{i})\leq t\}}{1\vee\sum_{i=1}^{n}\mathbbm{1}\{\varphi_{i}(p_{i})\leq t\}}\leq\alpha\right\},$ (10)
-which is the largest cutoff such that the FDP estimate is bounded above by $\alpha$, where $T_{\mathrm{up}}$ satisfies $T_{\mathrm{up}}<\min_{i}\varphi_{i}(0.5)$.
-The GBC procedure rejects $H_{i}$ whenever $\varphi_{i}(p_{i})\leq T_{\text{GBC}}$.
-Proposition 4.
-Suppose that the null p-values $\{p_{i}\}_{i\in\mathcal{H}_{0}}$ are mutually independent and satisfy Condition (8), and are independent of the alternative p-values $\{p_{i}\}_{i\notin\mathcal{H}_{0}}$.
-Assuming $\varphi_{i}$ is a monotonic increasing and continuous function for all $i$, then the GBC procedure ensures the FDR control at level $\alpha$.
-Compared to the GBH procedure, the GBC approach affords us greater flexibility in selecting $\varphi_{i}$, as it no longer requires $\varphi_{i}$ to be strictly increasing and its (generalized) inverse function does not have to fulfill the condition in Proposition 2.
-Example 3.
-Suppose the p-value $p_{i}$ is generated independently from the two-group mixture model: $\pi_{i}f_{0}+(1-\pi_{i})f_{1,i}$, where $\pi_{i}\in(0,1)$ is the mixing proportion and $f_{0}$ and $f_{1,i}$ denote the p-value distributions under the null and alternative respectively.
-The local FDR is defined as
-$\mathrm{Lfdr}_{i}(p)=\frac{\pi_{i}f_{0}(p)}{\pi_{i}f_{0}(p)+(1-\pi_{i})f_{1,i}(p)},$
-which is the posterior probability that the $i$th hypothesis is under the null given the observed p-value being $p$.
-The monotone likelihood ratio assumption (Sun and Cai, 2007) states that $f_{1,i}(p)/f_{0}(p)$ is decreasing in $p$.
-Under this assumption, $\varphi_{i}(p)=\mathrm{Lfdr}_{i}(p)$ is monotonically increasing in $p$ and thus fulfills the requirement in Proposition 4.
-Additionally, it has been shown in the literature that the rejection rule $\varphi_{i}(p_{i})=\mathrm{Lfdr}_{i}(p_{i})\leq t$ is optimal in the sense of maximizing the expected number of true positives among the decision rules that control the marginal FDR at level $\alpha$, see e.g., Sun and Cai (2007); Lei and Fithian (2018); Cao et al.
-(2022).
-3.4 Connection between the GBC and e-BH procedures
-We show that the GBC procedure is equivalent to the e-BH procedure with the following choice of e-values:
-$e_{i}=\frac{n\mathbbm{1}\{\varphi_{i}(p_{i})\leq T_{\text{GBC}}\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{\text{GBC}}\}},$ (11)
-where $T_{\text{GBC}}$ is defined in (10). By equation (B.1) in the proof of Proposition 4, we have
-$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[e_{i}]\leq n,$
-which implies that the corresponding e-BH procedure controls the FDR at the desired level.
-Furthermore, the following theorem shows that the e-BH procedure with the e-values defined above is equivalent to the GBC procedure.
-Theorem 3.
-Let $\mathcal{S}_{\mathrm{GBC}}$ be the set of rejections obtained through the GBC procedure at the FDR level $\alpha$, and let $\mathcal{S}_{\mathrm{eBH}}$ represent the set of rejections obtained from the e-BH procedure at the same FDR level $\alpha$, with the e-values defined in (11).
-Then we have $\mathcal{S}_{\mathrm{GBC}}=\mathcal{S}_{\mathrm{eBH}}$.
-4 Applications
-Building upon the insights from the previous section, we shall develop new multiple testing procedures by aggregating the e-values resulting from different procedures (e.g., the BH and BC procedures) or the same procedure but applied to different subsets of the data.
-By showing that the aggregated e-values fulfill Condition (3), we have the e-BH procedure based on the aggregated e-values controls the FDR in finite sample.
-We illustrate this idea through three concrete applications below.
-The first application develops a hybrid procedure to leverage the strengths from the BH and BC procedures, aiming to improve their robustness and efficiency.
-The second application investigates fairness in the multiple testing context.
-In particular, given a sensitive/protected variable that partitions the data into several groups, we develop a fair multiple testing procedure that ensures the FDR within each group and the overall FDR across all groups are controlled at the same desired level simultaneously.
-Our last application proposes a new structure adaptive multiple testing procedure through cross-fitting and e-value aggregation.
-We demonstrate that the proposed procedure has guaranteed FDR control and can deliver comparable or even higher power than existing competitors.
-4.1 A hybrid procedure
-Empirical results in the literature suggest that neither the BH procedure nor the BC procedure dominates the other.
-Specifically, in the case of sparse signals (i.e., $n_{0}/n\approx 1$), the BH procedure typically has higher power than the BC procedure (Arias-Castro and Chen, 2017).
-In contrast, when the signal is dense, the BH procedure becomes conservative, making the BC procedure more preferable.
-However, in real-world applications, it is often impossible to know the signal strength and density in advance, and thus it is unclear which method would perform better.
-Applying both methods and reporting the results associated with the one that delivers more rejections does not ensure FDR control.
-Given the insights from Sections 2.4-2.5, we ask the question of whether it is possible to develop a new multiple testing procedure that has finite sample FDR control and is powerful across a wider range of signals by borrowing the strengths from both the BH and BC procedures.
-To address this question, we introduce a hybrid approach that utilizes e-values as a bridge to combine the BH and BC procedures, aiming for better robustness and efficiency.
-Let $e_{\mathrm{BH},i}$ and $e_{\mathrm{BC},i}$ be the e-values from the BH and BC procedures (at levels $\alpha_{\mathrm{BH}}$ and $\alpha_{\mathrm{BC}}$) respectively for testing the $i$th hypothesis.
-We define the weighted e-values
-$e_{i}=w_{\mathrm{BH},i}e_{\mathrm{BH},i}+w_{\mathrm{BC},i}e_{\mathrm{BC},i},$
-that aggregates the information from both the BH and BC procedures, and apply the e-BH procedure to the resulting e-values, where $w_{\mathrm{BH},i}$ and $w_{\mathrm{BC},i}$ are non-negative weights.
-The detailed implementation of the hybrid procedure is described in Algorithm 1 below.
-Algorithm 1 Hybrid procedure
-1:p-values $p_{1},\dots,p_{n}$ and significance levels $\alpha_{\mathrm{BH}}$, $\alpha_{\mathrm{BC}}$ and $\alpha_{\mathrm{eBH}}$
-2:Implement the BH procedure at the significance level $\alpha_{\mathrm{BH}}$. Compute the threshold $T_{\mathrm{BH}}$ using (1). Determine the e-value for the BH procedure with the formula:
-$e_{\mathrm{BH},i}=\frac{1}{T_{\mathrm{BH}}}\mathbbm{1}\{p_{i}\leq T_{\mathrm{BH}}\}.$
-3:Implement the BC procedure at the significance level $\alpha_{\mathrm{BC}}$. Compute the threshold $T_{\mathrm{BC}}$ using (2). Determine the e-value for the BC procedure with the formula:
-$e_{\mathrm{BC},i}=\frac{n\mathbbm{1}\{p_{i}\leq T_{\mathrm{BC}}\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC}}\}}.$
-4:Calculate the weighted averaged e-value by
-$e_{i}=w_{\mathrm{BH},i}e_{\mathrm{BH},i}+w_{\mathrm{BC},i}e_{\mathrm{BC},i}.$ (12)
-5:Run the e-BH procedure utilizing the weighted averaged e-values at the significance level $\alpha_{\mathrm{eBH}}$.
-6:The indices of rejected hypotheses.
-4.1.1 Choice of significance levels and weights
-We now discuss the choices of the significance levels $(\alpha_{\mathrm{BH}},\alpha_{\mathrm{BC}})$ and the weights $(w_{\mathrm{BH},i},w_{\mathrm{BC},i})$ in Algorithm 1, which play important roles in the hybrid procedure.
-Different from the target FDR level $\alpha_{\mathrm{eBH}}$, the choice of $(\alpha_{\mathrm{BH}},\alpha_{\mathrm{BC}})$ does not affect the FDR control level but instead affect the power of the hybrid procedure.
-Following the discussions in Section 3.2 of Ren and Barber (2022), when there are $n_{a}$ non-nulls with extremely strong signals, we expect that
-$1+\sum_{j=1}^{n}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC}}\}\approx\sum_{j\in\mathcal{H}_{0}}\mathbbm{1}\{p_{j}\leq T_{\mathrm{BC}}\}\approx\frac{\alpha_{\mathrm{BC}}n_{a}}{1-\alpha_{\mathrm{BC}}}.$
-In a similar spirit, we expect the FDR of the BH procedure to be $\tau_{0}\alpha$, where $\tau_{0}=n_{0}/n.$ Let $R_{\mathrm{BH}}$ be the number of rejections in the BH procedure.
-Then we have $R_{\mathrm{BH}}\approx n_{a}/(1-\tau_{0}\alpha_{\mathrm{BH}})$ and
-$T_{\mathrm{BH}}=\frac{\alpha_{\mathrm{BH}}R_{\mathrm{BH}}}{n}\approx\frac{\alpha_{\mathrm{BH}}n_{a}}{n(1-\tau_{0}\alpha_{\mathrm{BH}})}.$
-Thus the hybrid procedure will reject $H_{i}$ with $i\notin\mathcal{H}_{0}$ when
-$e_{i}\approx\frac{\omega_{\mathrm{BC},i}n(1-\alpha_{\mathrm{BC}})}{\alpha_{\mathrm{BC}}n_{a}}+\frac{\omega_{\mathrm{BH},i}n(1-\tau_{0}\alpha_{\mathrm{BH}})}{\alpha_{\mathrm{BH}}n_{a}}\geq\frac{n}{\alpha_{\mathrm{eBH}}n_{a}},$
-which is equivalent to
-$\frac{\omega_{\mathrm{BC},i}(1-\alpha_{\mathrm{BC}})}{\alpha_{\mathrm{BC}}}+\frac{\omega_{\mathrm{BH},i}(1-\tau_{0}\alpha_{\mathrm{BH}})}{\alpha_{\mathrm{BH}}}\geq\frac{1}{\alpha_{\mathrm{eBH}}}.$ (13)
-We found that setting $\alpha_{\mathrm{BC}}=\alpha_{\mathrm{BH}}=\alpha_{\mathrm{eBH}}/(1+\alpha_{\mathrm{eBH}})$ fulfills the above constraint, and leads to good performance in our numerical studies.
-Next, we discuss the selection of $(w_{\mathrm{BH},i},w_{\mathrm{BC},i})$ which balances the contribution from the two methods.
-A natural choice would be to set $w_{\mathrm{BH},i}=w_{\mathrm{BC},i}=0.5$ for all $i$.
-The theorem below shows that if the weights are independent of the p-values and satisfy $w_{\mathrm{BH},i}+\max_{j}w_{\mathrm{BC},j}\leq 1$ for all $i$, then the e-values defined in (12) satisfy Condition (3).
-Therefore, according to Proposition 1, the corresponding e-BH procedure controls the FDR at the desired level.
-Theorem 4.
-Suppose that the null p-values $\{p_{i}\}_{i\in\mathcal{H}_{0}}$ are mutually independent and super-uniform, and are independent of the alternative p-values $\{p_{i}\}_{i\notin\mathcal{H}_{0}}$.
-Assume that $w_{\mathrm{BH},i}$ and $w_{\mathrm{BC},i}$ are independent of the p-values, and $w_{\mathrm{BH},i}+\max_{j}w_{\mathrm{BC},j}\leq 1$ for all $i$.
-Then, the weighted e-values given in (12) satisfy Condition (3).
-[…]
-$\begin{split}w_{\mathrm{BH},i}&=\frac{T_{\mathrm{BH},i}}{T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j,i}\}\right)},\\ w_{\mathrm{BC},i}&=\frac{\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC}}\}\right)}{\max_{j}T_{\mathrm{BH},j}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC}}\}\right)}.\end{split}$ (14)
-Theorem 5.
-Suppose that the null p-values $\{p_{i}\}_{i\in\mathcal{H}_{0}}$ are mutually independent and super-uniform, and are independent of the alternative p-values $\{p_{i}\}_{i\notin\mathcal{H}_{0}}$.
-Let $w_{\mathrm{BH},i}$ and $w_{\mathrm{BC},i}$ be defined in (14).
-Then, the weighted average e-values specified in (12) with the data-dependent weights given in (14) satisfy Condition (3).
-As a consequence of Theorem 5, the hybrid procedure with the data-dependent weights in (14) provides finite sample FDR control.
-Furthermore, in view of the proof of Theorem 5, the above conclusion remains true if we replace $T_{\mathrm{BH},i}$ in (14) by any function of $\tilde{p}_{-i}$.
-4.1.2 Numerical studies
-We investigate the finite sample performance of the hybrid procedure via several simulation examples.
-We fix the sample size $n=1,000$ and set the significance level at $\alpha_{\mathrm{eBH}}=0.05$.
-For each experimental setting, the average FDP (which estimates the FDR) and the average power based on 500 Monte Carlo replications are reported.
-We consider two different ways to implement the hybrid procedure: (1) eBH_Ada, for which the weights are calculated via (14) and $\alpha_{\mathrm{BH}}=\alpha_{\mathrm{BC}}=\alpha_{\mathrm{eBH}}/(1+\alpha_{\mathrm{eBH}})$; (2) eBH_Ave, for which the weights are set as $w_{\mathrm{BH},i}=w_{\mathrm{BC},i}=0.5$ for all $i=1,\cdots,n$ and $\alpha_{\mathrm{BH}}=\alpha_{\mathrm{BC}}=\alpha_{\mathrm{eBH}}/2$.
-Notice that the weights defined in (14) involve the term $T_{\mathrm{BC},j,i}$, which can be computationally expensive.
-To reduce the computational burden, we also consider a fast implementation of eBH_Ada, referred to as fast_eBH_Ada, which uses the weights
-$w_{\mathrm{BH},i}=\frac{T_{\mathrm{BH},i}}{T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j}\}\right)},$
-but otherwise is the same as eBH_Ada.
-We consider two settings.
-In Setting S1, we set $n_{a}$ to be $100$ and $200$, corresponding to $10\%$ and $20\%$ signals.
-For the null hypotheses, the p-values are generated independently from the uniform distribution on $[0,1]$; for the alternative hypotheses, the p-values are generated independently from the uniform distribution on $[0,p_{1}]$, where $p_{1}$ takes values between $1.25\times 10^{-2}$ and $1.5\times 10^{-2}$.
-Here, a larger $p_{1}$ indicates a weaker signal strength on average.
-Figure 1 summarizes the results for Setting S1.
-All methods under consideration control the FDR at the 5% level.
-For the dense signal case (i.e., 20% of the hypotheses are non-nulls), the BH procedure is quite conservative, while the BC procedure delivers significantly higher power.
-The performance of eBH_Ada is between the BH and BC procedures.
-In contrast, eBH_Ave does not provide much improvement over the BH procedure.
-It is also worth noting that fast_eBH_Ada delivers almost identical results compared to eBH_Ada but with a much lower computational cost.
-When $n_{a}=100$, the eBH_Ada procedure has nearly the same power as the BC procedure, and it outperforms eBH_Ave whose power is almost identical to the BH procedure.
-Figure 1: FDR and power for Setting S1.
-The left panel and the right panel correspond to the cases of $n_{a}=100$ and $n_{a}=200$, respectively.
-The results are based on $500$ Monte Carlo replications.
-In Setting S2.
-we set $n_{a}$ to be $50$ and 100, corresponding to $5\%$ and $10\%$ of all hypotheses.
-Under the null hypothesis, we generate the test statistics $X_{i}$ from the standard normal distribution $\mathcal{N}(0,1)$.
-Under the alternatives, $X_{i}$ follows $\mathcal{N}(\mu\log(n),1)$, where the parameter $\mu$ controls the signal strength, which varies between $0.3$ and $0.5$.
-Subsequently, the p-values are computed as $p_{i}=1-\Phi(X_{i})$, where $\Phi$ denotes the distribution function of the standard normal distribution.
-The simulation results are presented in Figure 2.
-When the signal density is 5% (i.e., $n_{a}=50$), the BH procedure performs better than the BC procedure.
-eBH_Ada has nearly the same power as the BH procedure, and it again outperforms eBH_Ave, which slightly improves over the BC procedure.
-When the signal density increases to 10% (i.e., $n_{a}=100$), the performance gap between the BH and BC procedures diminishes.
-eBH_Ada and its fast version produce almost the same power as the BH procedure, outperforming both the eBH_Ave and BC procedures.
-In this case, the performance of eBH_Ave is worse than that of the BH and BC procedures.
-Figure 2: FDR and power for Setting S2. The left panel and the right panel correspond to the cases of $n_{a}=50$ and $n_{a}=200$, respectively. Results are based on $500$ Monte Carlo replications.
-In summary, eBH_Ada is capable of improving the worse-performing method within the BH and BC procedures under different scenarios.
-Interestingly, its power can be nearly identical to the better-performing method, which demonstrates the adaptivity of eBH_Ada.
-fast_eBH_Ada delivers almost the same results as eBH_Ada in all settings with a much lower computational burden.
-We therefore recommend the use of fast_eBH_Ada in practice.
-4.2 Fair multiple testing
-In recent years, there has been a growing interest in developing machine learning algorithms that ensure certain notion of fairness for groups categorized by protected attributes, such as gender, race, or ethnicity.
-However, the development of statistical inference procedures that guarantee specific notions of fairness has not received the same level of attention.
-To address this gap, we introduce a new notion of fairness in the context of multiple hypothesis testing, and propose a procedure to ensure this notion of fairness.
-Suppose we have $n$ hypotheses which can be partitioned into $G$ distinct groups, denoted as $\{\mathcal{G}_{g}\colon g=1,\dots,G\}$, according to some protected attributes, where $\cup_{g=1}^{G}\mathcal{G}_{g}=[n]:=\{1,2,\dots,n\}$ and $\mathcal{G}_{g}\cap\mathcal{G}_{g^{\prime}}=\emptyset$ for $g\neq g^{\prime}$.
-Recall that $\theta_{i}\in\{0,1\}$ represents the underlying truth of $H_{i}$ and $\delta_{i}\in\{0,1\}$ denotes the decision rule for $H_{i}$.
-We define the group-wise FDP and FDR based on $\bm{\delta}$ respectively as:
-$\mathrm{FDP}_{g}(\bm{\delta})=\frac{\sum_{i\in\mathcal{G}_{g}}(1-\theta_{i})\delta_{i}}{1\vee\sum_{i\in\mathcal{G}_{g}}\delta_{i}},\quad\text{FDR}_{g}(\bm{\delta})=\mathbb{E}[\text{FDP}_{g}(\bm{\delta})],\quad g=1,2,\dots,G.$
-Definition 1.
-A decision rule $\bm{\delta}$ with the target FDR level $\alpha$ is called a fair decision rule if it controls both the group-wise FDRs uniformly over $1\leq g\leq G$ and the overall FDR at the level $\alpha$, i.e., $\max_{1\leq g\leq G}\text{FDR}_{g}(\bm{\delta})\leq\alpha$ and $\text{FDR}(\bm{\delta})\leq\alpha$.
-Remark 1.
-Among various notions of fairness that have been discussed in the literature (Verma and Rubin, 2018), predictive parity in the classification context (Chouldechova, 2017) is most relevant to our definition of fairness.
-To elaborate, let us consider a binary classification problem where $Y\in\{0,1\}$ represents the true labels and $\widehat{Y}\in\{0,1\}$ denotes the predicted labels.
-In this scenario, the FDR is defined as $P(Y=0|\widehat{Y}=1)$.
-A classifier satisfies predictive parity if it ensures an equal FDR across all groups.
-However, the multiple testing problem differs from the classification problem as the underlying truth of each hypothesis is unobserved and cannot be used to construct the decision rule.
-In our context, it is more appropriate to control the FDRs across different groups at the same desired level rather than forcing them to be the same.
-[…]
-[…]
-Algorithm A 0.039 0.006 0.067
-Algorithm B 0.057 0.048 0.025
-Table 1: FDR results for Setting E2, where the nominal FDR level is $5\%$.
-To overcome the issues above, we consider a multiple testing procedure by assembling the e-values from the BC procedure applied to each group separately.
+In recent work, Li and Zhang (2025) introduced a unified framework for understanding many commonly used multiple testing procedures and demonstrated their equivalence to the e-BH procedure using appropriately defined sets of e-values.
+In this section, we briefly review their results.
+Consider $n$ hypotheses $H_{1},\dots,H_{n}$.
+Let $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$ denote the sets of true null and true alternative hypotheses, respectively.
+Let $\bm{\theta}=(\theta_{1},\dots,\theta_{n})\in\{0,1\}^{n}$ represent the true states of these hypotheses, where $\theta_{i}=0$ indicates that $H_{i}$ is under the null and $\theta_{i}=1$ otherwise.
+We define a decision rule $\bm{\delta}=(\delta_{1},\dots,\delta_{n})\in\{0,1\}^{n}$, where $\delta_{i}=1$ indicates rejection of $H_{i}$, and $\delta_{i}=0$ indicates acceptance.
+The false discovery rate (FDR) associated with a decision rule $\bm{\delta}$ is defined as the expectation of the false discovery proportion (FDP):
+$\mathrm{FDR}(\bm{\delta})=\mathbb{E}[\mathrm{FDP}(\bm{\delta})],\quad\text{where}\quad\mathrm{FDP}(\bm{\delta})=\frac{\sum_{i=1}^{n}(1-\theta_{i})\delta_{i}}{1\vee\sum_{i=1}^{n}\delta_{i}},$
+with $a\vee b=\max\{a,b\}$. An FDR-controlling procedure ensures that the FDR does not exceed a pre-specified threshold $\alpha\in(0,1)$.
+2.1 Multiple Testing Procedures
+Suppose we observe a set of p-values $p_{1},p_{2},\dots,p_{n}$ corresponding to the hypotheses $H_{1},H_{2},\dots,H_{n}$.
+Li and Zhang (2025) summarized several commonly used multiple testing procedures using the following unified form.
+Consider
+$T=\sup\left\{t\in\mathcal{T}:\frac{m(t)}{1\vee\sum_{j=1}^{n}R_{j}(t)}\leq\alpha\right\},$ (1)
+where $\mathcal{T}$ denotes the domain of the threshold, $m(t)$ provides a conservative estimate of the number of false rejections, and $R_{i}(t)$ indicates whether to reject the $i$th hypothesis at threshold $t$.
+The decision rule rejects hypothesis $H_{i}$ if and only if $R_{i}(T)=1$.
+To specify an FDR-controlling procedure within this framework, one must define the functions $m(t)$ and $R_{i}(t)$.
+The most widely used method, the Benjamini-Hochberg (BH) procedure (Benjamini and Hochberg, 1995), sets $m(t)=nt$ and $R_{i}(t)=\mathbbm{1}\{p_{i}\leq t\}$, where $\mathbbm{1}\{A\}$ is the indicator function associated with set $A$.
+Storey’s (ST) procedure (Storey, 2002; Storey et al., 2004) refines the BH procedure by estimating the proportion of true null hypotheses from the observed p-values.
+Specifically, the ST procedure defines $R_{i}(t)=\mathbbm{1}\{p_{i}\leq t\}$ and $m(t)=n\pi_{0}^{\lambda}t$, where $\pi_{0}^{\lambda}\coloneqq\bigl{\{}1+n-R(\lambda)\bigr{\}}/\bigl{\{}(1-\lambda)n\bigr{\}}$ for a fixed $\lambda\in[0,1)$, and $R(\lambda)$ is the number of hypotheses rejected at threshold $\lambda$.
+Barber and Candès (2015) introduced the Barber-Candès (BC) procedure, a model-free approach leveraging symmetry properties of null p-values or test statistics to estimate false rejections.
+The BC procedure defines $m(t)=1+\sum_{i=1}^{n}\mathbbm{1}\{p_{i}\geq 1-t\}$ and $R_{i}(t)=\mathbbm{1}\{p_{i}\leq t\}$.
+The flexible BC (FBC) procedure proposed by Li and Zhang (2025) generalizes the BC approach using hypothesis-specific rejection functions $\varphi_{i}$, given by $m(t)=1+\sum_{i=1}^{n}\mathbbm{1}\{\varphi_{i}(1-p_{i})\leq t\}$ and $R_{i}(t)=\mathbbm{1}\{\varphi_{i}(p_{i})\leq t\}$.
+Table A.1 in Supplement A.1 summarizes the specifications of $m(t)$ and $R_{i}(t)$ for these procedures.
+2.2 E-Values and the e-BH Procedure
+A non-negative random variable $e$ is called an e-value if it satisfies the condition $\mathbb{E}[e]\leq 1$ under the null hypothesis.
+Suppose we have $n$ e-values, denoted as $e_{1},e_{2},\dots,e_{n}$, corresponding to the hypotheses $H_{1},H_{2},\dots,H_{n}$.
+The $\alpha$-level e-BH procedure (Wang and Ramdas, 2022) involves sorting the e-values in decreasing order $e_{(1)}\geq e_{(2)}\geq\dots\geq e_{(n)},$ and rejecting the hypotheses associated with the $\hat{k}$ largest e-values, where $\hat{k}\coloneqq\max\left\{1\leq i\leq n\colon e_{(i)}\geq n/(i\alpha)\right\}$.
+Let $\mathcal{H}_{0}=\{1\leq i\leq n\colon\theta_{i}=0\}$ be the set of true null hypotheses.
+According to Theorem 2 of Wang and Ramdas (2022), a key advantage of the e-BH procedure is that it controls the FDR at level $\alpha$, even when the e-values exhibit arbitrary dependence.
+Proposition 1 (Wang and Ramdas (2022), Theorem 2).
+Suppose the set of e-values $\{e_{i}\}_{1\leq i\leq n}$ satisfies
+$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[e_{i}]\leq n.$ (2)
+Then, the e-BH procedure controls the FDR at level $\alpha$.
+In the context of multiple testing, the requirement that $\mathbb{E}[e]\leq 1$ in the definition of e-values can be relaxed.
+Specifically, throughout the rest of this paper, we refer to $\{e_{i}\}$ as a set of e-values if they satisfy Condition (2).
+2.3 Connection Between Multiple-Testing Procedures and the e-BH Procedure
+Given the threshold $T$ defined in (1), the e-BH procedure, defined based on the e-values $e_{i}=\bigl{(}nR_{i}(T)\bigr{)}/m(T)$ for $1\leq i\leq n$, is equivalent to the multiple testing procedures presented in Table A.1 with the same $m(\cdot)$ and $R_{i}(\cdot)$ functions (Li and Zhang, 2025).
+Here, equivalence means that they produce the same set of rejections.
+Unlike p-values, which are defined in terms of tail probabilities, e-values are defined through expectations, making them easier to aggregate or combine.
+For instance, the arithmetic mean of multiple e-values remains a valid e-value.
+Building upon the insight that the BH and BC procedures, along with their generalized versions, are equivalent to the e-BH procedure when based on specific forms of e-values, we develop novel multiple testing methods by aggregating e-values derived from different procedures or by assembling e-values obtained from the same procedure applied to distinct subsets of data.
+By ensuring that these aggregated or assembled e-values satisfy Condition (2), the resulting e-BH procedures maintain finite-sample FDR control.
+We illustrate these developments with several concrete applications in Sections 3-5.
+3 Assembling E-Values Across Data Subsets
+In this section, we consider a scenario that we have $L$ sets of e-values, $\{e_{i}^{l}:i\in\mathcal{G}_{l},|\mathcal{G}_{l}|=n_{l}\}$, derived from $L$ distinct datasets, where $\bigcup_{l}\mathcal{G}_{l}=[n]$ and $\mathcal{G}_{l_{1}}\cap\mathcal{G}_{l_{2}}=\emptyset$ for $l_{1}\neq l_{2}$.
+Each $e_{i}^{l}$ is associated with the hypothesis $H_{i}$, and $\sum_{i\in\mathcal{G}_{l}\cap\mathcal{H}_{0}}\mathbb{E}[e_{i}^{l}]\leq n_{l}$.
+In this context, our objective is to combine the $L$ sets of e-values into a single e-value vector $(e_{1},\dots,e_{n})$ that satisfies Condition (2).
+3.1 Simultaneous Group-Wise and Overall FDR Control
+Recall that $\theta_{i}\in\{0,1\}$ represents the true state of hypothesis $H_{i}$, and $\delta_{i}\in\{0,1\}$ denotes the decision rule for $H_{i}$.
+We define the group-wise FDP and FDR based on $\bm{\delta}$ as follows:
+$\mathrm{FDP}_{l}(\bm{\delta})=\frac{\sum_{i\in\mathcal{G}_{l}}(1-\theta_{i})\delta_{i}}{1\vee\sum_{i\in\mathcal{G}_{l}}\delta_{i}},\quad\text{FDR}_{l}(\bm{\delta})=\mathbb{E}[\text{FDP}_{l}(\bm{\delta})],\quad l=1,2,\dots,L.$
+A decision rule $\bm{\delta}$ with target FDR level $\alpha$ is said to simultaneously control both group-wise and overall FDR if it uniformly controls the group-wise FDRs for all $1\leq l\leq L$ and maintains the overall FDR at level $\alpha$.
+Specifically, this means that $\max_{1\leq l\leq L}\text{FDR}_{l}(\bm{\delta})\leq\alpha$ and $\text{FDR}(\bm{\delta})\leq\alpha$.
+A decision rule that simultaneously controls both group-wise and overall FDR is relevant to predictive parity within the classification context in the fairness community (Chouldechova, 2017).
+We compare our definitions with predictive parity in Supplement D.
+We propose a multiple testing procedure that simultaneously controls both group-wise and overall FDR by assembling the e-values from the BC procedure applied to each group separately.
 Specifically, we implement the BC procedure at the level $\alpha$ for each individual group and let
-$T_{g}=\sup\left\{0<t<0.5\colon\frac{1+\sum_{i\in\mathcal{G}_{g}}\mathbbm{1}\{p_{i}\geq 1-t\}}{1\vee\sum_{i\in\mathcal{G}_{g}}\mathbbm{1}\{p_{i}\leq t\}}\leq\alpha\right\}$ (15)
-be the rejection threshold for the $g$th group with $1\leq g\leq G$. Define
-$e_{i}=\frac{n_{g}w_{i}\mathbbm{1}\{p_{i}\leq T_{g}\}}{1+\sum_{j\in\mathcal{G}_{g}}\mathbbm{1}\{p_{j}\geq 1-T_{g}\}},$ (16)
-for $i\in\mathcal{G}_{g}$, where $w_{i}>0$ represents the weight for the $i$th hypothesis, which will be specified in Section 4.2.1.
+$T_{l}=\sup\left\{0<t<0.5\colon\frac{1+\sum_{i\in\mathcal{G}_{l}}\mathbbm{1}\{p_{i}\geq 1-t\}}{1\vee\sum_{i\in\mathcal{G}_{l}}\mathbbm{1}\{p_{i}\leq t\}}\leq\alpha\right\}$ (3)
+be the rejection threshold for the $l$th group with $1\leq l\leq L$. Define
+$e_{i}=\frac{n_{l}w_{i}\mathbbm{1}\{p_{i}\leq T_{l}\}}{1+\sum_{j\in\mathcal{G}_{l}}\mathbbm{1}\{p_{j}\geq 1-T_{l}\}},$ (4)
+for $i\in\mathcal{G}_{l}$, where $w_{i}>0$ represents the weight for the $i$th hypothesis, which will be specified in Section 3.1.1.
 After collecting the e-values from each group, we implement the e-BH procedure at level $\alpha$.
 The testing procedure is summarized in Algorithm 1.
-Algorithm 2 Fair multiple testing
-p-values $p_{1},\dots,p_{n}$, group indices $\mathcal{G}_{1},\dots,\mathcal{G}_{G}$, significance level $\alpha$
-for $g=1,\dots,G$ do
-     Implement the BC procedure utilizing the p-values $\{p_{i}\colon i\in\mathcal{G}_{g}\}$ at the level $\alpha$.
-     Calculate the threshold $T_{g}$ using (15).
-     for $i\in\mathcal{G}_{g}$ do
-         Calculate the e-value $e_{i}$ using (16).
-     end for
-end for
-Assemble the e-values from all groups.
-Run the e-BH procedure utilizing the assembled e-values at the level $\alpha$.
-The indices of rejected hypotheses.
+Algorithm 1 Multiple testing procedure that simultaneously controls both group-wise and overall FDR
+1:p-values $p_{1},\dots,p_{n}$, group indices $\mathcal{G}_{1},\dots,\mathcal{G}_{L}$, significance level $\alpha$
+2:for $l=1,\dots,L$ do
+3:  Implement the BC procedure utilizing the p-values $\{p_{i}\colon i\in\mathcal{G}_{l}\}$ at the level $\alpha$.
+4:  Calculate the threshold $T_{l}$ using (3).
+5:  for $i\in\mathcal{G}_{l}$ do
+6:   Calculate the e-value $e_{i}$ using (4).
+7:  end for
+8:end for
+9:Assemble the e-values from all groups.
+10:Run the e-BH procedure utilizing the assembled e-values at the level $\alpha$.
+11:The indices of rejected hypotheses.
 It is important to note that only the nonzero e-values can be rejected in the e-BH procedure.
@@
 In the following section, we will demonstrate that the e-BH procedure can effectively control the overall FDR even when the weights are selected in a data-dependent manner.
-4.2.1 Choice of weights and FDR control
-Controlling the overall FDR requires ensuring that the e-values defined in (16) satisfy Condition (3).
-One strategy is to set $w_{i}=1$ for all $i$.
-Another strategy takes into account the group size by setting $w_{i}=n/(Gn_{g})$ for all $i\in\mathcal{G}_{g}$ and $g=1,\dots,G$, where $n_{g}=|\mathcal{G}_{g}|$ is the cardinality of the $g$th group.
-It can be verified that Condition (3) is satisfied by both strategies.
-The e-BH procedures based on these two weight choices are referred to as eBH_1 and eBH_2, respectively.
-Theorem 6.
-Suppose that the null p-values $\{p_{i}\}_{i\in\mathcal{H}_{0}}$ are mutually independent and satisfy Condition (8), and are independent of the alternative p-values $\{p_{i}\}_{i\notin\mathcal{H}_{0}}$.
-Further suppose the weights $w_{i}$ are independent of the p-values and satisfy
-$\sum_{g=1}^{G}n_{g}\max_{i\in\mathcal{G}_{g}}w_{i}\leq n.$ (17)
-Then, the e-values specified in (16) satisfy Condition (3). Hence, the corresponding e-BH procedure controls the FDR in finite sample.
-Our simulations show that eBH_1 and eBH_2 frequently suffer from low statistical power.
-To enhance efficiency, we suggest utilizing a data-dependent weight selection approach inspired by the leave-one-out technique (Barber et al., 2020).
-Denote the p-values in the $g$th group by $\mathbf{p}_{g}=\{p_{i}\}_{i\in\mathcal{G}_{g}}$.
-Write $\tilde{p}_{i}=\min\{p_{i},1-p_{i}\}$, and let $\mathbf{p}_{g,i}$ for $i\in\mathcal{G}_{g}$ be the collection of p-values obtained by replacing $p_{i}$ with $\tilde{p}_{i}$ in $\mathbf{p}_{g}$.
-By viewing $T_{g}$ as a functions of $\mathbf{p}_{g}$, we define $T_{g,i}=T_{g}(\mathbf{p}_{g,i})$, i.e., the threshold of the BC procedure applied to the set of p-values $\mathbf{p}_{g,i}$.
+3.1.1 Choosing Weights and FDR Control
+Controlling the overall FDR requires that the e-values defined in (4) satisfy Condition (2).
+One approach is to set $w_{i}=1$ for all $i$.
+Alternatively, the group size can be taken into account by setting $w_{i}=n/(Ln_{l})$ for all $i\in\mathcal{G}_{l}$ and $l=1,\dots,L$.
+According to Proposition 6 in Li and Zhang (2025), both strategies satisfy Condition (2).
+The e-BH procedures based on these weight choices are referred to as eBH_1 and eBH_2, respectively.
+However, our simulations indicate that eBH_1 and eBH_2 often suffer from low statistical power.
+To enhance efficiency, we propose using a data-dependent weight approach inspired by the leave-one-out technique (Barber et al., 2020).
+Denote the p-values in the $l$th group by $\mathbf{p}_{l}=\{p_{i}\}_{i\in\mathcal{G}_{l}}$.
+Write $\tilde{p}_{i}=\min\{p_{i},1-p_{i}\}$, and let $\mathbf{p}_{l,i}$ for $i\in\mathcal{G}_{l}$ be the collection of p-values obtained by replacing $p_{i}$ with $\tilde{p}_{i}$ in $\mathbf{p}_{l}$.
+By viewing $T_{l}$ as a functions of $\mathbf{p}_{l}$, we define $T_{l,i}=T_{l}(\mathbf{p}_{l,i})$, i.e., the threshold of the BC procedure applied to the set of p-values $\mathbf{p}_{l,i}$.
 We define the data-dependent weights as
-$w_{i}=\frac{\frac{n}{n_{g}}\left(1+\sum_{j\neq i,j\in\mathcal{G}_{g}}\mathbbm{1}\{p_{j}\geq 1-T_{g}\}\right)}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{g}}\mathbbm{1}\{p_{j}\geq 1-T_{g}\}\right)+\sum_{g^{\prime}\neq g}\sum_{j\in\mathcal{G}_{g^{\prime}}}\mathbbm{1}\{p_{j}\geq 1-T_{g^{\prime},j}\}},$ (18)
-for $i\in\mathcal{G}_{g}$. When $G=2$, we have
-$w_{i}=\frac{\frac{n}{n_{1}}\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1}\}\right)}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1}\}\right)+\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{p_{j}\geq 1-T_{2,j}\}}$
-for $i\in\mathcal{G}_{1}$ and
-$w_{i}=\frac{\frac{n}{n_{2}}\left(1+\sum_{j\neq i,j\in\mathcal{G}_{2}}\mathbbm{1}\{p_{j}\geq 1-T_{2}\}\right)}{\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1,j}\}+\left(1+\sum_{j\neq i,j\in\mathcal{G}_{2}}\mathbbm{1}\{p_{j}\geq 1-T_{2}\}\right)}$
-for $i\in\mathcal{G}_{2}$.
-The e-BH procedure based on the weights specified in (18) is referred to as the eBH_Ada method in the following discussions.
-The theorem below states that eBH_Ada has finite sample FDR control.
+$w_{i}=\frac{\frac{n}{n_{l}}\left(1+\sum_{j\neq i,j\in\mathcal{G}_{l}}\mathbbm{1}\{p_{j}\geq 1-T_{l}\}\right)}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{l}}\mathbbm{1}\{p_{j}\geq 1-T_{l}\}\right)+\sum_{l^{\prime}\neq l}\sum_{j\in\mathcal{G}_{l^{\prime}}}\mathbbm{1}\{p_{j}\geq 1-T_{l^{\prime},j}\}},$ (5)
+for $i\in\mathcal{G}_{l}$.
+The e-BH procedure, based on the weights specified in (5), will henceforth be referred to as the eBH_Ada method in the following discussions.
+If the null $p$-values satisfy the following condition:
+$P(p_{i}\leq a)\leq P(p_{i}\geq 1-a)=P(1-p_{i}\leq a),\quad\text{ for all }0\leq a\leq 0.5,$ (6)
+then eBH_Ada has finite-sample FDR control.
 Theorem 1.
@@
 Then, the e-values specified in (4) with the weights defined via (5) satisfy Condition (2).
-Hence, the corresponding e-BH procedure controls the FDR in finite sample.
+Hence, the corresponding e-BH procedure controls the overall FDR in finite sample.
+3.1.2 Illustrative Example of Data-Dependent Weights
+In this subsection, we present a toy example to illustrate the effectiveness of data-dependent weights.
+Consider a scenario where we have two groups of p-values.
+The first group contains $n_{1}=100$ p-values, and the second group contains $n_{2}=1000$ p-values.
+We apply the BC method with a threshold of $0.05$.
+In each group, we identify $n_{a1}=n_{a2}=20$ significant features, yielding e-values of $e_{1}=100$ and $e_{2}=1000$ for the first and second groups, respectively.
+Next, we combine the e-values from both groups into a single vector and sort them in decreasing order as $e_{(1)}\geq e_{(2)}\geq\cdots\geq e_{(n)}$, where $n=n_{1}+n_{2}=1100$.
+To select $40$ significant features, the combined e-values should satisfy $e_{(40)}\geq\frac{1100}{40\times 0.05}=550$.
+Similarly, to select $20$ significant features, the combined e-values should satisfy $e_{(20)}\geq\frac{1100}{20\times 0.05}=1100$.
+Without weighting, however, we find that $e_{(20)}=1000<1100$ and $e_{(40)}=100<550$.
+Thus, when the e-values are assembled without weights and the e-BH method is applied, no hypotheses are identified as significant.
+Figure 1: Comparison of weighted and unweighted e-values. Circles represent weighted e-values, while triangles denote unweighted e-values. Different colors indicate two groups.
+Figure 1 illustrates the scenario with only the non-zero e-values.
+Circles represent the weighted e-values, while triangles denote the e-values without weights.
+Different colors indicate the two groups.
+When e-values are combined without weights and the e-BH method is applied, no hypotheses are significant.
+In contrast, when the e-values are combined with weights, all hypotheses become significant.
+This demonstrates that data-dependent weights serve to increase the smaller e-values and decrease the larger e-values, thereby enhancing the method’s ability to identify more significant discoveries while maintaining control over the FDR.
 3.1.3 Numerical Studies
@@
 Firstly, it may fail to control the group-wise FDRs, as illustrated in Setting E2.
-Secondly, it fails to ensure comparable power across different groups, resulting in the possibility of one group having high power while the other group has nearly zero power, as illustrated in Setting E1.
+Secondly, it fails to ensure comparable power across different groups, resulting in the possibility of one group having high power.
+In contrast, the other group has nearly zero power, as illustrated in Setting E1.
 The same issue is also encountered by eBH_1.
@@
 Although BC_Sep effectively controls the FDR for individual groups, it does not guarantee the overall FDR control.
-Remark 2.
-We also tried BC_Sep at the level $\alpha/G$, a naive method that controls both the group-wise and overall FDRs.
-Indeed, let $\hat{n}_{ga}$ be the number of rejections for the $g$th group, and denote the number of false rejections for the $g$th group by $\hat{n}_{g0}$.
-Then we have $\mathbb{E}[\hat{n}_{g0}/(1\vee\hat{n}_{ga})]\leq\alpha/G$, which implies that
-$\mathbb{E}\left[\frac{\sum_{g=1}^{G}\hat{n}_{g0}}{1\vee\sum_{g=1}^{G}\hat{n}_{ga}}\right]\leq\mathbb{E}\left[\sum_{g=1}^{G}\frac{\hat{n}_{g0}}{1\vee\hat{n}_{ga}}\right]\leq\alpha.$
-However, this method has nearly zero power in all our simulation settings. Therefore, we have decided not to include its results in the tables below.
-[…]
+We first consider the case of two groups.
+To evaluate each method, we employ the following metrics: POW represents the overall power combining the rejections from both groups; POW1 denotes the power for the first group, while POW2 represents the power for the second group.
+Similarly, we can define FDR, FDR1, and FDR2.
+The empirical power and FDR are computed based on 1,000 independent Monte Carlo simulations.
 In all settings, we assume that the p-values follow the uniform distribution on $[0,1]$ under the null.
-For the first group, the p-value is assumed to follow $\mathrm{Unif}[0,u_{1}]$ under the alternatives, while for the second group, it follows $\mathrm{Unif}[0,u_{2}]$ under the alternatives.
-The parameter values for different settings are detailed in Table 2.
-Setting $n_{1}$ $n_{1a}$ $u_{1}$ $n_{2}$ $n_{2a}$ $u_{2}$
-E1 $100$ $20$ $1.5\times 10^{-2}$ $1000$ $20$ $1.5\times 10^{-3}$
-E2 $100$ $20$ $1.5\times 10^{-3}$ $1000$ $20$ $1.5\times 10^{-3}$
-Table 2: Parameter settings for the case of $G=2$.
-Here, $n_{g}$ represents the number of hypotheses for the $g$th group; $n_{ga}$ denotes the number of non-null hypotheses in the $g$th group.
-The parameter $u_{g}$ is the upper bound of the uniform distribution for the p-values under the alternatives for the $g$th group.
-Setting E1 corresponds to a scenario in which, for instance, the first group consists of ethnic minorities while the second group comprises of majorities.
-The number of non-nulls are the same across the two groups.
-The alternative p-values in the first group are larger than those in the second group on average as $u_{1}>u_{2}$.
-Table 3 summarizes the results for this setting.
-BC_Com exhibits high power for the second group, yet its power in the first group are rather low.
-This is because the non-null p-values from the first group are not sufficiently small, and a combined analysis of the two groups demands a lower threshold and thus fails to reject them.
-Additionally, BC_Sep has a slightly inflated overall FDR in this case.
-[…]
-BC_Com 0.144 0.041 0.248 0.023 0.012 0.025
-BC_Sep 0.296 0.352 0.24 0.053 0.036 0.025
-eBH_1 0.07 0 0.141 0.02 0 0.02
-eBH_2 0.078 0.077 0.078 0.008 0.008 0.008
-eBH_Ada 0.181 0.212 0.149 0.028 0.021 0.015
-Table 3: FDR and power for Setting E1, where the nominal FDR level is 5%.
-Setting E2 is the same as Setting E1 except that $u_{1}=u_{2}=1.5\times 10^{-2}$.
-The results for Setting E2 are presented in Table 4.
+For the first group, the p-value is supposed to follow Beta($\alpha_{1},\beta_{1}$) under the alternatives, while for the second group, it follows Beta($\alpha_{2},\beta_{2}$) under the alternatives.
+The parameter values for different settings are detailed in Table D.1 in Supplement D.1.
+Setting E1 corresponds to a scenario in which, for instance, the first group consists of ethnic minorities while the second group comprises ethnic majorities.
+The number of non-nulls is the same across the two groups.
+The alternative p-values in the first group are larger than those in the second group on average.
+The results for Setting E1 are presented in Table 1.
+BC_Com exhibits high power for the second group, yet its power in the first group is quite low.
+This is because the non-null p-values from the first group are not sufficiently small, and a combined analysis of the two groups demands a lower threshold, which thus fails to reject them.
+Additionally, BC_Sep has an inflated overall FDR in this case.
+Setting E1 Setting E2
+Method POW POW1 POW2 FDR FDR1 FDR2 POW POW1 POW2 FDR FDR1 FDR2
+BC_Com 0.21 0.01 0.41 0.036 0.043 0.034 0.49 0.49 0.49 0.036 0.006 0.062
+BC_Sep 0.378 0.336 0.420 0.060 0.035 0.035 0.416 0.727 0.105 0.056 0.048 0.017
+eBH_1 0.075 0.000 0.149 0.021 0.000 0.021 0.024 0.000 0.049 0.010 0.000 0.010
+eBH_2 0.127 0.128 0.126 0.012 0.013 0.010 0.079 0.082 0.077 0.009 0.005 0.012
+eBH_Ada 0.212 0.185 0.238 0.027 0.019 0.019 0.289 0.499 0.079 0.038 0.034 0.013
+Table 1: False discovery rate (FDR) and power for Settings E1 and E2 (nominal FDR level = 5%).
+The results for Setting E2 are also presented in Table 1.
 We observe that BC_Com fails to control the FDR for the second group, which can be explained as follows.
@@
 In contrast, all three variants of the e-BH procedure control the group-wise and overall FDRs at the desired level.
-eBH_Ada has a much higher power than the other two e-value based methods.
-[…]
-BC_Com 0.605 0.605 0.606 0.039 0.006 0.067
-BC_Sep 0.573 0.906 0.24 0.057 0.048 0.025
-eBH_1 0.07 0 0.141 0.02 0 0.02
-eBH_2 0.216 0.217 0.215 0.017 0.011 0.022
-eBH_Ada 0.384 0.552 0.216 0.036 0.029 0.023
-Table 4: FDR and power for Setting E2, where the nominal FDR level is 5%.
-We present the results for $G=4$ in Appendix C.1, where we consider three different scenarios (Settings F1-F3).
-In particular, we note that BC_Com suffers from a severe FDR inflation with the empirical FDR being 0.312 at the 5% target level in Setting F2.
-In Setting F3, BC_Sep has an empirical overall FDR being 0.346, which is much higher than the 20% level.
+eBH_Ada has a much higher power than the other two e-value-based methods.
+We present the results for $G=4$ in Supplement D.1, where we consider three different scenarios (Settings F1-F3).
+In particular, we note that BC_Com suffers from severe FDR inflation, with the empirical FDR reaching 0.318 at the 5% target level in Setting F2.
+In Setting F3, BC_Sep has an empirical overall FDR of 0.343, which is much higher than the 20% target level.
 To summarize, as seen in Settings E2 and F2, BC_Com has no guarantee in controlling the group-wise FDR.
-On the other hand, BC_Sep fails to control the overall FDR, as observed in all the settings, particularly Settings E2 and F3 in Appendix C.1.
-The e-BH based approaches provide both group-wise and overall FDR control, making them fair according to Definition 1.
+On the other hand, BC_Sep fails to control the overall FDR, as observed in all the settings, particularly Settings E2 and F3.
+The e-BH-based approaches provide both group-wise and overall FDR control.
 However, eBH_1 and eBH_2 may suffer from power loss under certain scenarios.
-In contrast, eBH_Ada demonstrates consistent effectiveness across all settings by achieving FDR fairness and reasonable power.
-4.3 Structure adaptive multiple testing
+In contrast, eBH_Ada demonstrates consistent effectiveness across all settings by achieving both group-wise and overall FDR control, while maintaining reasonable power.
+3.1.4 Real-Data Example
+We illustrate the proposed method by conducting a differential abundance analysis using the microbiome dataset cdi_schubert, obtained from the MicrobiomeHD repository (Duvallet et al., 2017), originally collected in a case-control study comparing individuals with Clostridium difficile infection (CDI) to those without (nonCDI).
+After preprocessing, the feature table contains 2,293 operational taxonomic units (OTUs), representing bacterial taxa annotated at the phylum level.
+We then applied the LinDA method (Zhou et al., 2022) to identify taxa that differ between the CDI and nonCDI groups.
+Additional details about the dataset, the LinDA method, and implementation procedures are provided in Supplement D.2.
+In this study, controlling the overall FDR ensures the reliability of global inference, whereas controlling the FDR within each phylum is essential for accurately interpreting results within biologically meaningful groups.
+The number of rejected taxa for each phylum is summarized in Table D.6 in Supplement D.2.
+The results show that for the phyla Bacteroidetes and Firmicutes, the eBH_Ada method yields fewer rejections compared to the BC_Com method, suggesting that BC_Com might inadequately control FDR within these specific groups.
+Conversely, for the phylum Proteobacteria, the eBH_Ada methods identify a greater number of rejections, mirroring the pattern observed in our simulation scenario E1, where the BC_Com method exhibits reduced power in certain groups.
+Moreover, the two data-independent weighting schemes, eBH_1 and eBH_2, produced no discoveries; consequently, we omit their results from the table.
+This outcome highlights the necessity of using data-dependent weights when aggregating e-values.
+4 Aggregating E-Values From Different Results
+In this section, we consider a scenario where we have $L$ sets of e-values, $\{e_{i}^{l}:i\in[n]\}_{l=1}^{L}$, potentially derived from $L$ distinct multiple testing procedures or the same multiple testing procedure with different test statistics or tuning parameters.
+Here, $\{e_{i}^{l}\}_{l=1}^{L}$ denotes the $L$ e-values associated with hypothesis $H_{i}$, and it satisfies $\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[e_{i}^{l}]\leq n.$ Our objective is to aggregate these $L$ sets of e-values into a single e-value vector $e_{1},\dots,e_{n}]$ that satisfies Condition (2).
+We illustrate the proposed idea with a knockoff example in Supplement E, where we introduce a hybrid knockoff procedure that integrates multiple test statistics to achieve robustness across diverse modeling scenarios.
+Due to space constraints, a comprehensive discussion of the knockoff framework and the hybrid method is deferred to the Supplement.
+Algorithm 2 Hybrid Procedure
+1:p-values $p_{1},\dots,p_{n}$ and significance levels $\alpha_{\mathrm{BH}}$, $\alpha_{\mathrm{BC}}$, and $\alpha_{\mathrm{eBH}}$
+2:Execute the BH procedure at significance level $\alpha_{\mathrm{BH}}$.
+Compute the threshold $T_{\mathrm{BH}}$ using (1) with the BH-specified $m(t)$ and $R_{i}(t)$ defined in Table A.1.
+Calculate the e-value for the BH procedure as
+$e_{\mathrm{BH},i}=\frac{1}{T_{\mathrm{BH}}}\mathbbm{1}\{p_{i}\leq T_{\mathrm{BH}}\}.$
+3:Execute the BC procedure at significance level $\alpha_{\mathrm{BC}}$.
+Compute the threshold $T_{\mathrm{BC}}$ using (1) with the BC-specified $m(t)$ and $R_{i}(t)$ defined in Table A.1.
+Calculate the e-value for the BC procedure as
+$e_{\mathrm{BC},i}=\frac{n\mathbbm{1}\{p_{i}\leq T_{\mathrm{BC}}\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC}}\}}.$
+4:Compute the weighted averaged e-value using (7).
+5:Apply the e-BH procedure using the weighted average e-values at significance level $\alpha_{\mathrm{eBH}}$.
+6:Indices of rejected hypotheses.
+4.1 Hybrid Multiple-Testing Procedure
+The second application of our proposed method leverages e-values to integrate results from both the BC and BH procedures.
+Empirical results in the literature indicate that neither the BH procedure nor the BC procedure consistently outperforms the other Arias-Castro and Chen (2017).
+In real-world applications, it is often impossible to determine which method will perform better.
+Applying both methods and reporting the results of the one that yields more rejections does not guarantee FDR control.
+In this section, we introduce a new multiple testing procedure that ensures finite-sample FDR control and maintains high power across a broader range of signals by leveraging the strengths of both the BH and BC procedures.
+Let $e_{\mathrm{BH},i}$ and $e_{\mathrm{BC},i}$ denote the e-values from the BH and BC procedures (at significance levels $\alpha_{\mathrm{BH}}$ and $\alpha_{\mathrm{BC}}$, respectively) for testing the $i$th hypothesis.
+We define the weighted e-values as
+$e_{i}=w_{\mathrm{BH},i}e_{\mathrm{BH},i}+w_{\mathrm{BC},i}e_{\mathrm{BC},i},$ (7)
+which aggregates information from both the BH and BC procedures, where $w_{\mathrm{BH},i}$ and $w_{\mathrm{BC},i}$ are non-negative weights.
+We then apply the e-BH procedure to these aggregated e-values to obtain our rejection set.
+The detailed implementation is given in Algorithm 2.
+4.1.1 Choosing Significance Levels and Weights
+We now discuss the choices of the significance levels $(\alpha_{\mathrm{BH}},\alpha_{\mathrm{BC}})$ and the weights $(w_{\mathrm{BH},i},w_{\mathrm{BC},i})$ in Algorithm 2, which play important roles in the hybrid procedure.
+Different from the target FDR level $\alpha_{\mathrm{eBH}}$, the choice of $(\alpha_{\mathrm{BH}},\alpha_{\mathrm{BC}})$ does not affect the FDR control level but instead affects the power of the hybrid procedure.
+Following the discussion in Section 3.2 of Ren and Barber (2024), when there are $n_{a}$ non-nulls with extremely strong signals, we expect that $1+\sum_{j=1}^{n}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC}}\}\approx\bigl{(}\alpha_{\mathrm{BC}}n_{a}\bigr{)}/\bigl{(}1-\alpha_{\mathrm{BC}}\bigr{)}$.
+In a similar spirit, we expect the FDR of the BH procedure to be $\tau_{0}\alpha_{\mathrm{BH}}$, where $\tau_{0}=n_{0}/n.$ Let $R_{\mathrm{BH}}$ be the number of rejections in the BH procedure.
+Then we have $R_{\mathrm{BH}}\approx n_{a}/(1-\tau_{0}\alpha_{\mathrm{BH}})$ and $T_{\mathrm{BH}}\approx\bigl{(}\alpha_{\mathrm{BH}}n_{a}\bigr{)}/\bigl{(}n(1-\tau_{0}\alpha_{\mathrm{BH}})\bigr{)}.$ Thus the hybrid procedure will reject $H_{i}$ with $i\notin\mathcal{H}_{0}$ when
+$e_{i}\approx\frac{w_{\mathrm{BC},i}n(1-\alpha_{\mathrm{BC}})}{\alpha_{\mathrm{BC}}n_{a}}+\frac{w_{\mathrm{BH},i}n(1-\tau_{0}\alpha_{\mathrm{BH}})}{\alpha_{\mathrm{BH}}n_{a}}\geq\frac{n}{\alpha_{\mathrm{eBH}}n_{a}}.$
+We found that setting $\alpha_{\mathrm{BC}}=\alpha_{\mathrm{BH}}=\alpha_{\mathrm{eBH}}/(1+\alpha_{\mathrm{eBH}})$ fulfills the above constraint if $w_{\mathrm{BC},i}+w_{\mathrm{BH},i}\leq 1$, and leads to good performance in our numerical studies.
+Next, we consider the selection of $(w_{\mathrm{BH},i},w_{\mathrm{BC},i})$, which balances the contributions from the two methods.
+A natural choice is to set $w_{\mathrm{BH},i}=w_{\mathrm{BC},i}=0.5$ for all $i$.
+According to Proposition 5 in Li and Zhang (2025), the e-values defined in (7) with these data-independent weights satisfy Condition (2).
+Therefore, by Proposition 1, the corresponding e-BH procedure controls the FDR at the desired level.
+Our simulations indicate that the e-BH procedure, which is based on averaged e-values, achieves only a slight improvement in power over the less powerful method between the BH and BC procedures.
+To address this issue, we introduce a data-dependent approach to construct weights.
+Our idea is partly motivated by the leave-one-out technique used in proving the FDR control for the BH (Ferreira and Zwinderman, 2006) and BC procedures (Barber et al., 2020).
+For $i=1,\dots,n$, denote $\tilde{p}_{i}\coloneqq\min\{p_{i},1-p_{i}\}$, $\mathbf{p}_{-i}\coloneqq\{p_{1},\dots,p_{i-1},\,\tilde{p}_{i},\,p_{i+1},\dots,p_{n}\}$ and $\tilde{\mathbf{p}}_{-i}\coloneqq\{\tilde{p}_{1},\dots,\tilde{p}_{i-1},\,0,\,\tilde{p}_{i+1},\dots,\tilde{p}_{n}\}$.
+By viewing $T_{\mathrm{BH}}$ and $T_{\mathrm{BC}}$ as functions of the p-values, we define $T_{\mathrm{BH},i}=T_{\mathrm{BH}}(\tilde{\mathbf{p}}_{-i})$ and $T_{\mathrm{BC},j}=T_{\mathrm{BC}}(\mathbf{p}_{-j})$.
+Further define $T_{\mathrm{BC},j,i}$ in the same way as $T_{\mathrm{BC},j}$ but with $p_{i}$ being replaced by 0 when $j\neq i$.
+We propose the following e-value weights
+$\begin{split}w_{\mathrm{BH},i}&=\frac{T_{\mathrm{BH},i}}{T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j,i}\}\right)},\\ w_{\mathrm{BC},i}&=\frac{\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC}}\}\right)}{\max_{j}T_{\mathrm{BH},j}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC}}\}\right)}.\end{split}$ (8)
+Theorem 2.
+Suppose that the null p-values $\{p_{i}\}_{i\in\mathcal{H}_{0}}$ are mutually independent and super-uniform, and are independent of the alternative p-values $\{p_{i}\}_{i\notin\mathcal{H}_{0}}$.
+Let $w_{\mathrm{BH},i}$ and $w_{\mathrm{BC},i}$ be defined in (8).
+Then, the weighted average e-values specified in (7) with the data-dependent weights given in (8) satisfy Condition (2).
+As a consequence of Theorem 2, the hybrid procedure with the data-dependent weights in (8) provides finite-sample FDR control.
+Furthermore, in view of the proof of Theorem 2, the above conclusion remains true if we replace $T_{\mathrm{BH},i}$ in (8) by any (deterministic) function of $\tilde{\mathbf{p}}_{-i}$.
+4.1.2 Numerical Studies
+We investigate the finite sample performance of the hybrid procedure via several simulation examples.
+We set the significance level at $\alpha_{\mathrm{eBH}}=0.05$.
+For each experimental setting, the average FDP (which estimates the FDR) and the average power based on 500 Monte Carlo replications are reported.
+We consider two different ways to implement the hybrid procedure: (1) eBH_Ada, for which the weights are calculated via (8) and $\alpha_{\mathrm{BH}}=\alpha_{\mathrm{BC}}=\alpha_{\mathrm{eBH}}/(1+\alpha_{\mathrm{eBH}})$; (2) eBH_Ave, for which the weights are set as $w_{\mathrm{BH},i}=w_{\mathrm{BC},i}=0.5$ for all $i=1,\dots,n$ and $\alpha_{\mathrm{BH}}=\alpha_{\mathrm{BC}}=\alpha_{\mathrm{eBH}}/2$.
+Notice that the weights defined in (8) involve the term $T_{\mathrm{BC},j,i}$, which can be computationally expensive.
+To reduce the computational burden, we also consider a fast implementation of eBH_Ada, referred to as fast_eBH_Ada, which uses the weights
+$w_{\mathrm{BH},i}=\frac{T_{\mathrm{BH},i}}{T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j}\}\right)},$
+but otherwise, it is the same as eBH_Ada.
+We generate p-values from two settings, Setting S1 and Setting S2, where the BH procedure outperforms the BC procedure in Setting S1 while the BC procedure provides significantly higher power in Setting S2.
+The simulation setups are deferred to Supplement F.
+The left panel of Figure 2 summarizes the results for Setting S1.
+All methods under consideration control the FDR at the 5% level.
+eBH_Ada demonstrates nearly the same power as the BH procedure and surpasses eBH_Ave, which offers a slight improvement over the BC procedure.
+Notably, fast_eBH_Ada achieves results almost identical to those of eBH_Ada but at a significantly lower computational cost.
+The right panel of Figure 2 summarizes the results for Setting S2.
+The performance of eBH_Ada lies between that of the BH and BC procedures.
+In contrast, eBH_Ave shows little improvement over the BH procedure.
+Again, fast_eBH_Ada produces nearly identical results to eBH_Ada but with a much lower computational cost.
+Figure 2: The left panel and right panel correspond to FDR and power for Setting S1 and Setting S2, respectively. The results are based on $500$ Monte Carlo replications.
+Next, we fix one setting in Setting S1 and calculate the computational cost for each method, as shown in Table F.1 in Supplement F.
+We observe that the data-independent method, eBH_Ave, performs the analysis as quickly as the BH or BC procedures.
+The fast_eBH_Ada method demonstrates acceptable speed and is significantly faster than the eBH_Ada method.
+In summary, eBH_Ada effectively enhances the performance of the weaker method between the BH and BC procedures across different scenarios.
+Notably, its power can be nearly identical to that of the stronger of the two procedures, highlighting the adaptivity of eBH_Ada.
+fast_eBH_Ada achieves almost identical results to eBH_Ada in all settings, while significantly reducing computational cost.
+Therefore, we recommend fast_eBH_Ada for practical applications.
+5 Structure-Adaptive Multiple Testing
 Having access to various types of auxiliary information that reflect the structural relationships among hypotheses is becoming increasingly common.
@@
 This side information represents heterogeneity among the p-values and may affect the prior probabilities of the null hypotheses being true or the signal strength under alternatives.
-Our goal is to develop a multiple testing procedure that can incorporate such external structural information to improve statistical power and has guaranteed FDR control in finite sample.
-The high-level idea behind our approach is to relax the p-value thresholds for hypotheses that are more likely to be non-null and tighten the thresholds for the other hypotheses through the use of hypothesis-specific rejection rule, i.e., $\varphi_{i}(p_{i})\leq t$, so that the FDR can be controlled.
-Our proposed method combines the cross-fitting technique (Ignatiadis and Huber, 2021) (a sample-splitting and fitting approach that enables learning the hypothesis-specific rejection function $\varphi_{i}$ without overfitting as long as the hypotheses can be partitioned into independent folds) with the GBC procedure introduced in Section 3.
+Our goal is to develop a multiple testing procedure that can incorporate such external structural information to improve statistical power and guarantee FDR control in finite sample.
+The high-level idea behind our approach is to relax the p-value thresholds for hypotheses more likely to be non-null and tighten the thresholds for others through the use of a hypothesis-specific rejection rule, i.e., $\varphi_{i}(p_{i})\leq t$, so that the FDR can be controlled.
+Our proposed method combines the cross-fitting technique (Ignatiadis and Huber, 2021) (a sample-splitting and fitting approach that enables learning the hypothesis-specific rejection function $\varphi_{i}$ without overfitting as long as the hypotheses can be partitioned into independent folds) with the FBC procedure introduced in Li and Zhang (2025).
 First, we randomly split the data into $G$ distinct groups, denoted as $\{\mathcal{G}_{g}\colon g=1,\dots,G\}$, where $\bigcup_{g=1}^{G}\mathcal{G}_{g}=[n]$ and $\mathcal{G}_{g}\cap\mathcal{G}_{g^{\prime}}=\emptyset$ for $g\neq g^{\prime}$.
-We estimate the rejection function $\varphi_{i}$ for the hypothesis $i$ in the $g$th group using the data from all other groups, which ensures that the estimated rejection function is independent of the p-values in group $g$.
-We then apply the GBC procedure based on the estimated rejection functions separately to each group and obtain the corresponding e-values.
+We estimate the rejection function $\varphi_{i}$ for hypothesis $i$ in the $g$th group using the data from all other groups, which ensures that the estimated rejection function is independent of the p-values in group $g$.
+We then apply the FBC procedure based on the estimated rejection functions separately to each group and obtain the corresponding e-values.
 Finally, we collect all the e-values and apply the e-BH procedure at the target level to control the FDR.
 To describe the cross-fitting procedure, let us assume that $\varphi_{i}(p)=\varphi(p,x_{i};\beta)$ for some unknown parameter $\beta$ that needs to be estimated from the data.
-We define the cross-fitted estimate as
-$\hat{\beta}_{-g}=\operatorname*{arg\,min}_{\beta}\sum_{i\notin\mathcal{G}_{g}}\mathcal{L}(p_{i},x_{i},\beta).$
-Here $\mathcal{L}$ is some loss function such as negative log-likelihood under the two-group mixture model in Example 3, in which case we assume $\pi_{i}(p)=\pi(p,x_{i};\beta_{1})$, $f_{1,i}(p)=f(p,x_{i};\beta_{2})$ and $\beta=(\beta_{1},\beta_{2}).$ It should be noted that the choice of the functional form of $\varphi_{i}$ is highly flexible and does not impact the FDR control due to cross-fitting that avoids over-fitting.
-Given $\hat{\beta}_{-g}$, we define
-$\hat{\varphi}_{i}(p)=\varphi(p,x_{i};\hat{\beta}_{-g})\quad\text{for}\quad i\in\mathcal{G}_{g}.$
-Next, we apply the GBC procedure using the cross-fitted functions $\{\hat{\varphi}_{i}(\cdot)\}_{i\in\mathcal{G}_{g}}$ at the level $\alpha_{\mathrm{GBC}}$.
+We define the cross-fitting estimate as $\hat{\beta}_{-g}=\operatorname*{arg\,min}_{\beta\in\mathcal{B}}\sum_{i\notin\mathcal{G}_{g}}\mathcal{L}(p_{i},x_{i},\beta)$.
+Here $\mathcal{L}$ is some loss function, such as negative log-likelihood, and $\mathcal{B}$ is a parameter space.
+Given $\hat{\beta}_{-g}$, we define $\hat{\varphi}_{i}(p)=\varphi(p,x_{i};\hat{\beta}_{-g})$ for $i\in\mathcal{G}_{g}$.
+Next, we apply the FBC procedure using the cross-fitted functions $\{\hat{\varphi}_{i}(\cdot)\}_{i\in\mathcal{G}_{g}}$ at the level $\alpha_{\mathrm{FBC}}$.
 The corresponding threshold for the $g$th group is given by
-$T_{g}=\sup\left\{0<t\leq T_{g,\mathrm{up}}\colon\frac{1+\sum_{i\in\mathcal{G}_{g}}\mathbbm{1}\{\hat{\varphi}_{i}(1-p_{i})\leq t\}}{1\vee\sum_{i\in\mathcal{G}_{g}}\mathbbm{1}\{\hat{\varphi}_{i}(p_{i})\leq t\}}\leq\alpha_{\mathrm{GBC}}\right\},$ (19)
-where $T_{g,\mathrm{up}}$ satisfies $T_{g,\mathrm{up}}<\min_{i\in\mathcal{G}_{g}}\hat{\varphi}_{i}(0.5)$. We define the e-value for all $i\in\mathcal{G}_{g}$ as
+$T_{g}=\sup\left\{0<t\leq T_{g,\mathrm{up}}\colon\frac{1+\sum_{i\in\mathcal{G}_{g}}\mathbbm{1}\{\hat{\varphi}_{i}(1-p_{i})\leq t\}}{1\vee\sum_{i\in\mathcal{G}_{g}}\mathbbm{1}\{\hat{\varphi}_{i}(p_{i})\leq t\}}\leq\alpha_{\mathrm{FBC}}\right\},$ (9)
+where $T_{g,\mathrm{up}}<\min_{i\in\mathcal{G}_{g}}\hat{\varphi}_{i}(0.5)$. We define the e-value for all $i\in\mathcal{G}_{g}$ as
 $e_{i}=\frac{n_{g}w_{i}\mathbbm{1}\{\hat{\varphi}_{i}(p_{i})\leq T_{g}\}}{1+\sum_{j\in\mathcal{G}_{g}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{g}\}},$ (10)
@@
 Finally, we aggregate all e-values from each group and implement the e-BH procedure.
-A detailed description of our procedure is given in Algorithm 3.
-Algorithm 3 Cross-fitting based structure adaptive multiple testing
-p-values $p_{1},\dots,p_{n}$, covariates $x_{1},\dots,x_{n}$, group indices $\mathcal{G}_{1},\dots,\mathcal{G}_{G}$, significance levels $\alpha_{\mathrm{GBC}}$ and $\alpha_{\mathrm{eBH}}$
-for $g=1,\dots,G$ do
-     Compute the estimate of the rejection rule $\hat{\varphi}_{i}(\cdot)$ for $i\in\mathcal{G}_{g}$ using $\mathbf{p}_{-g}$ and $\mathbf{x}_{-g}$.
-     Calculate the threshold $T_{g}$ using (19).
-     for $i\in\mathcal{G}_{g}$ do
-         Calculate the e-value $e_{i}$ using (20).
-     end for
-end for
-Assemble the e-values from all groups.
-Run the e-BH procedure utilizing the assembled e-values at the level $\alpha_{\mathrm{eBH}}$.
-The indices of rejected hypotheses.
-4.3.1 Choice of weights and FDR control
+A detailed description of our procedure is given in Algorithm G.1 in Supplement G.1.
+5.1 Weights and FDR Control
 To ensure that the FDR is controlled at the desired level, it is crucial to verify that the e-values defined in equation (10) satisfy Condition (2).
 We shall show that under certain conditions on the weights, the e-values defined by (10) satisfy (2), and as a result, the corresponding e-BH procedure controls the FDR at the desired level.
-Before stating the main theorem, let us first introduce some notation.
+Before stating the main theorem, let us first introduce some notations.
 Define $\tilde{p}_{i}=\min\{p_{i},1-p_{i}\}$, $\mathbf{p}_{g}=\{p_{i}\}_{i\in\mathcal{G}_{g}}$, and $\mathbf{p}_{g,i}$ as the collection of p-values obtained by replacing $p_{i}$ with $\tilde{p}_{i}$ in $\mathbf{p}_{g}$ for $i\in\mathcal{G}_{g}$.
@@
 Let $\{(p_{i},x_{i})\}$ for $1\leq i\leq n$ be the p-value and covariate pairs.
-  1. (A)
-    The null pairs $\{(p_{i},x_{i})\}_{i\in\mathcal{H}_{0}}$ are mutually independent.
-  2. (B)
-    The null pairs $\{(p_{i},x_{i})\}_{i\in\mathcal{H}_{0}}$ are independent of the alternative pairs $\{(p_{i},x_{i})\}_{i\notin\mathcal{H}_{0}}$.
-  3. (C)
-    For $i\in\mathcal{H}_{0}$, $p_{i}$ is independent of $x_{i}$ and satisfies Condition (8).
+(A) The null pairs $\{(p_{i},x_{i})\}_{i\in\mathcal{H}_{0}}$ are mutually independent.
+(B) The null pairs $\{(p_{i},x_{i})\}_{i\in\mathcal{H}_{0}}$ are independent of the alternative pairs $\{(p_{i},x_{i})\}_{i\notin\mathcal{H}_{0}}$.
+(C) For $i\in\mathcal{H}_{0}$, $p_{i}$ is independent of $x_{i}$ and satisfies Condition (6).
 Assumption 2.
@@
 Assumption 2 implies that $P(\varphi_{i}(p_{i},x_{i};\beta)\leq b)\leq P(\varphi_{i}(1-p_{i},x_{i};\beta)\leq b)$ for all $\varphi_{i}(0,x_{i};\beta)\leq b\leq\varphi_{i}(0.5,x_{i};\beta)$, which will be used in our proof.
-We describe a concrete choice of $\varphi_{i}$ in Section 4.3.2 below.
+We shall describe a concrete choice of $\varphi_{i}$ in Section 5.2 below.
 Theorem 3.
-Suppose Assumptions 1 and 2 hold. If the weights $\{w_{i}\}$ are independent of the p-values and covariate information and satisfy (17), then the e-values defined in (20) fulfills Condition (3).
-A naive choice is to set $w_{i}=1$ for all $i=1,2,\cdots,n$, which satisfy (17).
+Suppose Assumptions 1 and 2 hold. If the weights $\{w_{i}\}$ are independent of the p-values and covariate information and satisfy
+$\sum_{g=1}^{G}n_{g}\max_{i\in\mathcal{G}_{g}}w_{i}\leq n,$ (11)
+then the e-values defined in (10) fulfill Condition (2).
+A naive choice is to set $w_{i}=1$ for all $i=1,2,\dots,n$, which satisfies (11).
 However, this choice of weights often leads to low statistical power in simulations.
@@
 $w_{i}=\frac{\frac{n}{n_{g}}\left(1+\sum_{j\neq i,j\in\mathcal{G}_{g}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{g}\}\right)}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{g}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{g}\}\right)+\sup_{p\in[0,1]}\sum_{g^{\prime}\neq g}\sum_{j\in\mathcal{G}_{g^{\prime}}}\mathbbm{1}\{\hat{\varphi}_{j}^{g,i,p}(1-p_{j})\leq T_{g^{\prime},j}^{g,i,p}\}}.$ (12)
-In the case of $G=2$, we have
-$w_{i}=\frac{\frac{n}{n_{1}}\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1}\}\right)}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1}\}\right)+\sup_{p\in[0,1]}\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{\hat{\varphi}_{j}^{1,i,p}(1-p_{j})\leq T_{2,j}^{1,i,p}\}}$
-for $i\in\mathcal{G}_{1}$ and
-$w_{i}=\frac{\frac{n}{n_{2}}\left(1+\sum_{j\neq i,j\in\mathcal{G}_{2}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{2}\}\right)}{\sup_{p\in[0,1]}\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}^{2,i,p}(1-p_{j})\leq T_{1,j}^{2,i,p}\}+\left(1+\sum_{j\neq i,j\in\mathcal{G}_{2}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{2}\}\right)}$
-for $i\in\mathcal{G}_{2}$.
-The construction of $w_{i}$ involves taking supremum over $p$, which is crucial for the proof to go through.
+The construction of $w_{i}$ involves taking the supremum over $p$, which is crucial for the proof to go through.
 On the one hand, it renders $w_{i}$ independent of $p_{i}$, which is a useful fact in the proof.
-On the other hand, it makes the weight sufficiently small in the sense that we can upper bound $w_{i}$ with the term $\sup_{p\in[0,1]}\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}^{2,i,p}(1-p_{j})\leq T_{1,j}^{2,i,p}\}$ in its denominator replaced by $\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1,j}\}$, which is another fact used in our argument.
+On the other hand, it makes the weight sufficiently small in the sense that we can upper bound $w_{i}$ with the term $\sup_{p\in[0,1]}\sum_{j\in\mathcal{G}_{g^{\prime}}}\mathbbm{1}\{\hat{\varphi}_{j}^{g,i,p}(1-p_{j})\leq T_{g^{\prime},j}^{g,i,p}\}$ in its denominator replaced by $\sum_{j\in\mathcal{G}_{g^{\prime}}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{g^{\prime},j}\}$, which is another fact used in our argument.
 Theorem 4.
 Suppose Assumptions 1 and 2 hold. Then, the e-values defined in (10) with weights specified by (12) satisfy Condition (2).
-4.3.2 Simulation studies
+The $\tau$-censored weighted BH procedure proposed by Zhao and Zhou (2024) is a variant of the weighted BH procedure that employs a leave-one-out strategy to construct weights.
+A detailed comparison between our method and theirs is provided in Supplement G.1.
+5.2 Simulation Studies
 We shall compare the finite sample performance of the proposed method with several existing approaches through simulation studies.
 Throughout, we fix the sample size $n=3,000$ and set the target FDR level at $\alpha_{\mathrm{eBH}}=0.1$.
-For each experimental setting, we conduct 100 simulations, and report the average FDP (as an estimate of the FDR) and power over the independent simulation runs.
-We begin by providing the implementation details of the proposed method.
-In the GBC procedure, we utilize the rejection rule based on the local FDR under the two group mixture model in Example 3, where we set $f_{0}(p)=\mathbbm{1}\{p\in[0,1]\}$ and $f_{1,i}(p)=(1-\pi_{i})(1-\kappa_{i})p^{-\kappa_{i}}$ with $\kappa_{i}\in(0,1).$ We consider the working models that link $(\pi_{i},\kappa_{i})$ with the external covariates:
-$\pi_{i}$ $=\pi_{\beta_{\pi}}(x_{i})=\frac{1}{1+\exp(-\beta_{\pi,0}-\beta_{\pi,1}^{\top}x_{i})},$
-$\kappa_{i}$ $=\kappa_{\beta_{\kappa}}(x_{i})=\frac{1}{1+\exp(-\beta_{\kappa,0}-\beta_{\kappa,1}^{\top}x_{i})},$
-where the parameters $\beta_{\pi}=(\beta_{\pi,0},\beta_{\pi,1}^{\top})\in\mathbb{R}^{d+1}$ and $\beta_{\kappa}=(\beta_{\kappa,0},\beta_{\kappa,1}^{\top})\in\mathbb{R}^{d+1}$ can be estimated by maximizing the pseudo-log-likelihood using the EM algorithm.
-Please refer to Zhang and Chen (2022) for more optimization details.
-After obtaining the estimates $\hat{\beta}_{\pi}$ and $\hat{\beta}_{\kappa}$ from the EM algorithm, we define
-$\hat{\pi}_{i}=\begin{cases}\epsilon_{1}&\text{if }1/\bigl{(}1+\exp(-\hat{\beta}_{\pi,0}-\hat{\beta}_{\pi,1}^{\top}x_{i})\bigr{)}\leq\epsilon_{1},\\ 1/\bigl{(}1+\exp(-\hat{\beta}_{\pi,0}-\hat{\beta}_{\pi,1}^{\top}x_{i})\bigr{)}&\text{if }\epsilon_{1}<1/\bigl{(}1+\exp(-\hat{\beta}_{\pi,0}-\hat{\beta}_{\pi,1}^{\top}x_{i})\bigr{)}<1-\epsilon_{2},\\ 1-\epsilon_{2}&\text{otherwise,}\end{cases}$
-where winsorization is used to prevent $\hat{\pi}_{i}$ from being too close to zero or one to stabilize the algorithm. We define the rejection rule
-$\hat{\varphi}_{i}(p)=\frac{1}{\hat{\pi}_{i}+(1-\hat{\pi}_{i})(1-\hat{\kappa}_{i})p^{-\hat{\kappa}_{i}}}\leq t,$
-where $\hat{\kappa}_{i}=1/\bigl{(}1+\exp(-\beta_{\kappa,0}-\hat{\beta}_{\kappa,1}^{\top}x_{i})\bigr{)}$.
-Additionally, to speed up the procedure, we propose the following weight that is computationally less expensive:
-$w_{i}=\frac{\frac{n}{n_{g}}\left(1+\sum_{j\neq i,j\in\mathcal{G}_{g}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{g}\}\right)}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{g}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{g}\}\right)+\sum_{g^{\prime}\neq g}\sum_{j\in\mathcal{G}_{g^{\prime}}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{g,j}\}}.$
-We fix $\epsilon_{1}=0.1$, $\epsilon_{2}=1\times 10^{-5}$, $G=2$, and $\alpha_{\mathrm{GBC}}=\alpha_{\mathrm{eBH}}/(1+\alpha_{\mathrm{eBH}})$.
-We refer to this method as eBH_GBC for future reference.
-We compare the proposed method with the following competing methods:
-  • BH: The BH procedure (Benjamini and Hochberg, 1995). We implement this method using the p.adjust function in R.
-  • IHW_storey: The covariate-powered cross-weighted method with the Storey’s procedure to estimate the null-proportion (Ignatiadis and Huber, 2021).
-We implement this method using the ihw_bh function in the R package IHWStatsPaper.
-  • IHW_betamix: The covariate-powered cross-weighted method with the beta mixture model (Ignatiadis and Huber, 2021).
-We implement this method using the ihw_betamix_censored function in the R package IHWStatsPaper.
-  • AdaPT: The adaptive p-value thresholding procedure (Lei and Fithian, 2018). We implement this method using the adapt_glm function in the R package adaptMT.
-  • SABHA: The structure adaptive BH procedure (Li and Barber, 2019). The code was downloaded from the link provided by the original paper.
-To illustrate the effect of the covariate, we generate a single covariate $x_{i}$ from the standard normal distribution. Given the value of $x_{i}$, we define $\pi_{i}$ as
-$\pi_{i}=\frac{\exp(a_{0}+a_{1}x_{i})}{1+\exp(a_{0}+a_{1}x_{i})},$
-where $a_{0}$ and $a_{1}$ determine the baseline signal density and the informativeness of the covariate, respectively.
+For each experimental setting, we conduct 100 simulations and report the average FDP (as an estimate of the FDR) and power over the independent simulation runs.
+We begin by detailing the implementation of the proposed method.
+In the FBC procedure, we employ a rejection rule based on the local FDR (Sun and Cai, 2007) within the two-group mixture model framework.
+Specifically, we propose to use
+$\varphi_{i}(p)=\mathrm{Lfdr}_{i}(p)=\frac{\pi_{i}f_{0}(p)}{\pi_{i}f_{0}(p)+(1-\pi_{i})f_{1,i}(p)},$
+which represents the posterior probability that the $i$th hypothesis is null given the observed p-value $p$.
+The literature demonstrates that the rejection rule $\mathbbm{1}\{\varphi_{i}(p_{i})=\mathrm{Lfdr}_{i}(p_{i})\leq t\}$ is optimal in maximizing the expected number of true positives among decision rules that control the marginal FDR at level $\alpha$ (see, e.g., Sun and Cai (2007); Lei and Fithian (2018); Cao et al.
+(2022)).
+Additional discussions about local FDR are deferred to Supplement G.1.
+Set $f_{0}(p)=\mathbbm{1}\{p\in[0,1]\}$ and $f_{1,i}(p)=(1-\pi_{i})(1-\kappa_{i})p^{-\kappa_{i}}$ with $\kappa_{i}\in(0,1)$, we consider the working models proposed in Zhang and Chen (2022), please refer to Supplement G.1 for more details.
+After getting $\hat{\pi}_{i}$ and $\hat{\kappa}_{i}$, we define the rejection rule
+$\hat{\varphi}_{i}(p)=\frac{\hat{\pi}_{i}}{\hat{\pi}_{i}+(1-\hat{\pi}_{i})(1-\hat{\kappa}_{i})p^{-\hat{\kappa}_{i}}}\leq t.$
+We then apply the FBC procedure with the estimated $\hat{\varphi}_{i}$ at the target FDR level $\alpha$.
+The corresponding e-values are computed via (10), and the weights are obtained from (12).
+To reduce computational cost, we introduce the following, less expensive weighting scheme:
+$w_{i}=\frac{\frac{n}{n_{g}}\left(1+\sum_{j\neq i,j\in\mathcal{G}_{g}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{g}\}\right)}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{g}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{g}\}\right)+\sum_{g^{\prime}\neq g}\sum_{j\in\mathcal{G}_{g^{\prime}}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{g^{\prime},j}\}}.$
+We refer to this method as eBH_FBC for future reference.
+We compare the proposed method with the following competing methods: BH, IHW_storey, IHW_betamix, AdaPT, and SABHA.
+The implementation details of these methods are deferred to Supplement G.1.
+To illustrate the effect of the covariate, we generate a single covariate $x_{i}$ from the standard normal distribution.
+Given the value of $x_{i}$, we define $\pi_{i}$ as $\pi_{i}=\exp(a_{0}+a_{1}x_{i})/\bigl{(}1+\exp(a_{0}+a_{1}x_{i})\bigr{)}$, where $a_{0}$ and $a_{1}$ determine the baseline signal density and the informativeness of the covariate, respectively.
 The values of $a_{0}$ and $a_{1}$ are fixed for each simulated dataset.
-Specifically, we set $a_{0}$ to take on values from the set $\{3.5,1.5,1.5\}$, achieving signal densities of approximately $3\%$, $8\%$, and $18\%$, respectively, representing sparse, medium, and dense signals.
+Specifically, we set $a_{0}$ to take on values from the set $\{3.5,2.5,1.5\}$, achieving signal densities of approximately $3\%$, $8\%$, and $18\%$, respectively, representing sparse, medium, and dense signals.
 Furthermore, we set $a_{1}$ to take on values from the set $\{1.5,2,2.5\}$, representing a less informative, moderately informative, and strongly informative covariate, respectively.
 The underlying truth $\theta_{i}$ is then simulated based on $\pi_{i}:\theta_{i}\sim\text{Bernoulli}(1-\pi_{i}).$ We next generate the covariate that affects the alternative function $f_{1,i}$.
-Specifically, we sample another covariate $x_{i}^{\prime}\sim\mathcal{N}(0,1)$ and define
-$\eta_{i}=\frac{2\exp(a_{f}x_{i}^{\prime})}{1+\exp(a_{f}x_{i}^{\prime})},$
-where we set $a_{f}\in\{0,0.5,1\}$ for no informativeness, less informativeness, and strong informativeness. Then, the z-scores are sampled from
-$z_{i}\sim\mathcal{N}(\eta_{i}\mu\theta_{i},1),$
-where $\mu$ denotes the signal strength with the values evenly distributed in the interval $[2.5,3.4]$.
+Specifically, we sample another covariate $x_{i}^{\prime}\sim\mathcal{N}(0,1)$ and define $\eta_{i}=2\exp(a_{f}x_{i}^{\prime})/\bigl{(}1+\exp(a_{f}x_{i}^{\prime})\bigr{)}$, where we set $a_{f}\in\{0,0.5,1\}$ for no informativeness, less informativeness, and strong informativeness.
+Then, the z-scores are sampled from $z_{i}\sim\mathcal{N}(\eta_{i}\mu\theta_{i},1)$, where $\mu$ denotes the signal strength with the values evenly distributed in the interval $[2.5,3.4]$.
 These $z$-scores are transformed into p-values using the one-sided formula $1-\Phi(z_{i})$.
-The p-values, along with the corresponding covariates $x_{i}$ and $x_{i}^{\prime}$, serve as the input for the structure adaptive multiple testing methods.
-Figure 3: Empirical FDR and power for $a_{f}=1$.
-Figure 3 shows the results for $a_{f}=1$.
+The p-values, along with the corresponding covariates $x_{i}$ and $x_{i}^{\prime}$, serve as the input for the structure-adaptive multiple testing methods.
+Figure 3: Empirical FDR and power with $a_{f}=1$.
+Signal sparsity is controlled by setting $a_{0}\in\{3.5,2.5,1.5\}$, giving rise to sparse, moderate, and dense alternatives, respectively.
+Covariate informativeness is tuned via $a_{1}\in\{1.5,2,2.5\}$, corresponding to weak, moderate, and strong auxiliary signals.
+Figure 3 presents the results for $a_{f}=1$.
 All methods successfully controlled the FDR at the desired level.
-When the signal is sparse ($a_{0}=3.5$), eBH_GBC is the most powerful method.
-The SABHA method has the second-best performance, while the two versions of IHW show only slight improvement over the BH procedure.
+When the signal is sparse ($a_{0}=3.5$), eBH_FBC is the most powerful method.
+The SABHA method exhibits the second-best performance, while both versions of IHW show only slight improvements over the BH procedure.
 When the covariate is less informative ($a_{1}=1.5$), AdaPT is less powerful than the BH procedure.
-As the covariate becomes strongly informative ($a_{1}=2.5$), all structure adaptive methods outperform the BH procedure in terms of power.
-Our proposed method exhibits the highest power in most cases, with only AdaPT surpassing eBH_GBC in power when the signal is dense ($a_{0}=1.5$) and the covariate is strongly informative ($a_{1}=2.5$).
-The results for the settings with $a_{f}=0$ and $a_{f}=0.5$ are deferred to Appendix C.2.
-4.3.3 Real data examples
+As the covariate becomes strongly informative ($a_{1}=2.5$), all structure-adaptive methods outperform the BH procedure in terms of power.
+Our proposed method demonstrates the highest power in most cases, with AdaPT surpassing eBH_FBC in power when the signal is dense ($a_{0}=1.5$) and the covariate is strongly informative ($a_{1}=2.5$).
+The results for settings with $a_{f}=0$ and $a_{f}=0.5$ are deferred to Supplement G.2.
+To demonstrate the stability of each method in structure adaptive multiple testing, particularly the consistency of results when the data are randomly generated from the same distribution, we plot the variance of the FDP in Figure G.1 in Supplement G.2.
+The figure shows that as the signal becomes dense or the covariate contains more information, the variance of all methods decreases.
+Notably, the proposed method exhibits a smaller variance compared to AdaPT.
+Additionally, Table G.1 in Supplement G.2 displays the running time for each method to compare their computational efficiency.
+We focus on a simulation setting with $a_{0}=1.5$, $a_{1}=2$, $a_{f}=1$, and $\mu=3$.
+We conducted 100 simulation runs for each method and reported the average time taken to complete the analysis.
+The results indicate that our method is approximately ten times faster than AdaPT.
+5.2.1 Real-Data Examples
 We analyzed three omics datasets: Airway (Himes et al., 2014), Bottomly (Bottomly et al., 2011), and MWAS (McDonald et al., 2018).
-The Airway and Bottomly datasets are transcriptomics data obtained from RNA-seq experiments, which have been analyzed previously by Ignatiadis et al.
-(2016); Lei and Fithian (2018); Zhang and Chen (2022).
-These two datasets are available in the airway and IHWStatsPaper packages, respectively.
-For both datasets, we used the logarithm of the “basemean” as the covariate and removed the samples with missing values, leaving us with 18,028 and 11,709 tests, respectively.
+The Airway and Bottomly datasets are transcriptomics data obtained from RNA-seq experiments.
+For both datasets, we used the logarithm of the ‘basemean” as the covariate and removed the samples with missing values, leaving us with 18,028 and 11,709 tests, respectively.
 We obtained the MWAS dataset from the publicly available data of the American Gut project (McDonald et al., 2018).
 We focused on a subset of subjects with ages greater than thirteen and with complete sex and country information.
-We excluded OTUs (clustered sequencing units representing bacterial species) observed in fewer than ten subjects, resulting in 3,394 OTUs tested using the Wilcoxon rank sum test on normalized abundances.
+We excluded OTUs observed in fewer than ten subjects, resulting in 3,394 OTUs tested using the Wilcoxon rank sum test on normalized abundances.
 We used the library size of samples as the external covariate.
-Figure 4: Number of discoveries of various methods with the target FDR level ranging from $0.01$ to $0.1$ in three real datasets.
 The results of different methods for target FDR levels ranging from $0.01$ to $0.1$ are presented in Figure 4.
-AdaPT and eBH_GBC are the two methods that make the most discoveries (except for the MWAS data set with a target FDR level below 0.025).
-For the airway dataset, AdaPT consistently produces the most discoveries when the target FDR level is high.
-This is due to the high signal density of this data set.
+AdaPT and eBH_FBC are the two methods that make the most discoveries (except for the MWAS data set with a target FDR level below 0.025).
+For the airway dataset, AdaPT consistently produces the most discoveries at higher target FDR levels, which is due to the high signal density of this dataset.
 For instance, when the target FDR level is 10%, AdaPT is able to identify 6,053 discoveries out of 18,028 tests.
@@
 We observe a similar phenomenon for the Bottomly dataset.
-For the MWAS dataset, AdaPT and eBH_GBC fail to make any discoveries when the target FDR level is 1%.
-This is a limitation of the BC-based method, which may have reduced power when the signal is very sparse.
-However, as the target FDR level increases, AdaPT and eBH_GBC quickly surpass the other methods in terms of the number of discoveries.
-eBH_GBC outperforms AdaPT with more discoveries when the FDR level is above 5%.
-Overall, eBH_GBC performs comparably to AdaPT.
+For the MWAS dataset, both AdaPT and eBH_FBC fail to make any discoveries when the target FDR level is set to 1%.
+This is a limitation of the BC-type method, which may have reduced power when the signal is very sparse.
+However, as the target FDR level increases, AdaPT and eBH_FBC quickly outperform the other methods in terms of the number of discoveries.
+eBH_FBC outperforms AdaPT with more discoveries when the FDR level is above 5%.
+Overall, eBH_FBC performs comparably to AdaPT.
+Figure 4: Number of discoveries of various methods with the target FDR level ranging from $0.01$ to $0.1$ in three real datasets.
 6 Discussions
-In this paper, we have shown that the BH and BC procedures, as well as their generalized versions, are all equivalent to the e-BH procedure based on some forms of e-values.
-This simple insight opens up new avenues for constructing multiple testing procedures in different contexts.
-Specifically, leveraging this idea, one can translate the testing results from different procedures or the same procedure for different subsets of the data into e-values.
-By aggregating or assembling these e-values, one can obtain a combined set of e-values that contains valuable information about various multiple testing procedures and subsets of the data.
-We have demonstrated the potential of this strategy in three applications, backed by rigorous theoretical analysis and numerical studies.
-An important aspect of our proposed methods is the use of data-dependent weights, which are constructed using the leave-one-out technique to aggregate and assemble e-values.
-Using these weights, we can obtain valid e-values that ensure the corresponding e-BH procedure controls the FDR in finite sample.
-The e-BH procedure based on the data-dependent weight is often more powerful than the unweighted version, but it comes at the cost of more expensive computation.
-Through simulations, we have demonstrated that a more cost-friendly version of these weights exists, which achieves nearly identical performance with a much lower computational burden.
+Motivated by the recent findings in Li and Zhang (2025), we transform testing results from different procedures or data subsets into e-values.
+By aggregating these e-values, we obtain a combined set of e-values that captures information across procedures or partitions.
+A key feature of our method is the use of data-dependent weights, constructed via a leave-one-out approach, to ensure the resulting e-values yield finite-sample FDR control under the e-BH procedure.
+This weighted version is often more powerful than its unweighted counterpart.
+Simulations further reveal that a computationally efficient approximation of the weights achieves comparable performance.
 We envision the idea of aggregating different multiple testing results through e-values to be useful in other contexts, such as meta-analysis or federated learning.
 Other interesting future research problems include finding the optimal way of combining the e-values with respect to certain criteria and investigating the robustness of the proposed methods when the data exhibit dependence.
-Appendix
-Appendix A Proof of the main results
-We first state the following propositions whose proofs are deferred to Appendix  B. These results will be used frequently in the subsequent proofs of the main theorems.
-Proposition 5 (Lemma 6 of Barber et al. [2020]).
+SUPPLEMENTARY MATERIAL
+Supplement:
+Including all the proofs, additional discussions, and numerical results.
+Appendices
+Appendix A Preliminaries
+A.1 Unified Form of Multiple Testing Procedures
+Table A.1 summarizes the specifications of $m(t)$ and $R_{i}(t)$ for different multiple testing procedures.
+Method $m(t)$ $R_{i}(t)$
+BH ntnt $\mathbbm{1}\{p_{i}\leq t\}$
+ST $n\pi_{0}^{\lambda}t$ $\mathbbm{1}\{p_{i}\leq t\}$
+BC $1+\sum_{i=1}^{n}\mathbbm{1}\{p_{i}\geq 1-t\}$ $\mathbbm{1}\{p_{i}\leq t\}$
+FBH $ng(t)$ $\mathbbm{1}\{\varphi_{i}(p_{i})\leq t\}$
+FBC $1+\sum_{i=1}^{n}\mathbbm{1}\{\varphi_{i}(1-p_{i})\leq t\}$ $\mathbbm{1}\{\varphi_{i}(p_{i})\leq t\}$
+Table A.1: Definitions of $m(t)$ and $R_{i}(t)$ for various multiple testing procedures.
+Appendix B Proofs of Main Results
+We begin by stating the following propositions, whose proofs are deferred to Appendix C to ensure self-containment.
+These results will be frequently utilized in the subsequent proofs of the main theorems.
+Proposition B.1 (Lemma 6 of Barber et al. [2020]).
 Let $T_{\mathrm{BC},i}$ be the threshold for the BC methods when $p_{i}$ is replaced with $\min\{p_{i},1-p_{i}\}$.
 For any $i$, $j$, if $\min(p_{i},p_{j})\geq 1-\max\{T_{\mathrm{BC},i},T_{\mathrm{BC},j}\}$, then we have $T_{\mathrm{BC},i}=T_{\mathrm{BC},j}$.
-Proposition 6.
-Suppose the assumptions in Proposition 4 hold.
+Proposition B.2 (Proposition A.2 of Li and Zhang [2025]).
+Suppose that the null p-values are mutually independent, are independent of the alternative p-values, and satisfy Condition (6).
 Let $T_{i}$ be the threshold for the generalized BC methods when $p_{i}$ is replaced with $\min\{p_{i},1-p_{i}\}$.
 For any $i$, $j$, if $\max\{\varphi_{i}(1-p_{i}),\varphi_{j}(1-p_{j})\}\leq\max\{T_{i},T_{j}\}$, then we have $T_{i}=T_{j}$.
-A.1 Proof of Theorem 1
-Proof.
-Theorem 1 is a special case of Theorem 2 by choosing $\varphi_{i}$ as the identity function for all $1\leq i\leq n$. ∎
-A.2 Proof of Theorem 2
-Proof.
-Let $\hat{k}$ be the cardinality of $\mathcal{S}_{\mathrm{GBH}}$.
-If $i\notin\mathcal{S}_{\mathrm{GBH}}$, then $\varphi_{i}(p_{i})>T_{\mathrm{GBH}}$ and thus $e_{i}=0$.
-This implies that the $i$th hypothesis is not rejected by the e-BH procedure, and hence $\mathcal{S}_{\mathrm{eBH}}\subset\mathcal{S}_{\mathrm{GBH}}$.
-Conversely, if $i\in\mathcal{S}_{\mathrm{GBH}}$, we have $\varphi_{i}(p_{i})\leq T_{\mathrm{GBH}}\leq\bar{F}^{-1}\left(\frac{\alpha\hat{k}}{n}\right)$, leading to
-$e_{i}=\frac{1}{\bar{F}(T_{\mathrm{GBH}})}\geq\frac{n}{\alpha\hat{k}}$
-as $\bar{F}$ is strictly increasing.
-Define $e_{(i)}=n\mathbbm{1}\{q_{(i)}\leq T_{\mathrm{GBH}}\}/\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}})$ for $1\leq i\leq n,$ where $q_{(1)}\leq\dots\leq q_{(n)}$ are the order statistics of $\{q_{i}=\varphi_{i}(p_{i})\}^{n}_{i=1}$.
-Let $\hat{k}$ represent the maximum $i$ for which $q_{(i)}\leq T_{\mathrm{GBH}}$.
-We get
-$e_{(\hat{k})}\geq\frac{n}{\alpha\hat{k}},$ (A.1)
-which indicates that $|\mathcal{S}_{\mathrm{eBH}}|\geq\hat{k}$.
-Because $\mathcal{S}_{\mathrm{eBH}}\subset\mathcal{S}_{\mathrm{GBH}}$, it is clear that $\mathcal{S}_{\mathrm{eBH}}=\mathcal{S}_{\mathrm{GBH}}$.
-∎
-A.3 Proof of Theorem 3
-Let $\hat{k}$ be the cardinality of $\mathcal{S}_{\mathrm{GBC}}$.
-If $i\notin\mathcal{S}_{\mathrm{GBC}}$, then $\varphi_{i}(p_{i})>T_{\mathrm{GBC}}$ and thus $e_{i}=0$.
-Hence the $i$th hypothesis is not rejected by the e-BH procedure, which implies that $\mathcal{S}_{\mathrm{eBH}}\subset\mathcal{S}_{\mathrm{GBC}}$.
-For the other direction, note that if $i\in\mathcal{S}_{\mathrm{GBC}}$, then $\varphi_{i}(p_{i})\leq T_{\mathrm{GBC}}$ and $1+\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{\mathrm{GBC}}\}\leq\hat{k}\alpha$.
-Hence, we have
-$e_{i}\geq\frac{n}{\hat{k}\alpha}.$
-We sort the e-values in descending order as $e_{(1)}\geq\dots\geq e_{(n)}$.
-It is clear that $e_{(\hat{k})}\geq n/(\hat{k}\alpha)$.
-Thus, $|\mathcal{S}_{\mathrm{eBH}}|\geq\hat{k}$, which implies that $\mathcal{S}_{\mathrm{eBH}}=\mathcal{S}_{\mathrm{GBC}}$.
-A.4 Proof of Theorem 4
-Proof.
-Consider the BH procedure and observe that for a given number of rejections $R_{\mathrm{BH}}$, $T_{\mathrm{BH}}=T_{\mathrm{BH}}(R_{\mathrm{BH}})$ is a deterministic function of $R_{\mathrm{BH}}$.
-Let $R_{\mathrm{BH}}(p_{i}\rightarrow 0)$ be the number of rejections obtained by replacing the p-value $p_{i}$ with 0.
-Using the above fact and that the weights are independent of the p-values, we have
-$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[w_{\mathrm{BH},i}e_{\mathrm{BH},i}]=$ $\sum_{i\in\mathcal{H}_{0}}w_{\mathrm{BH},i}\mathbb{E}\left[\frac{1}{T_{\mathrm{BH}}}\mathbbm{1}\{p_{i}\leq T_{\mathrm{BH}}\}\right]$
-$=$ $\sum_{i\in\mathcal{H}_{0}}w_{\mathrm{BH},i}\sum_{k=1}^{n}\mathbb{E}\left[\frac{1}{T_{\mathrm{BH}}(k)}\mathbbm{1}\{p_{i}\leq T_{\mathrm{BH}}(k),R_{\mathrm{BH}}(p_{i}\rightarrow 0)=k\}\right]$
-$\leq$ $\sum_{i\in\mathcal{H}_{0}}w_{\mathrm{BH},i}\sum_{k=1}^{n}\mathbb{E}[\mathbbm{1}\{R_{\mathrm{BH}}(p_{i}\rightarrow 0)=k\}]$
-$=$ $\sum_{i\in\mathcal{H}_{0}}w_{\mathrm{BH},i}.$
-For the BC procedure, denote $\tilde{p}_{i}\coloneqq\min\{p_{i},1-p_{i}\}$ and $\mathbf{p}_{-i}\coloneqq\{p_{1},\dots,p_{i-1},\,\tilde{p}_{i},\,p_{i+1},\dots,p_{n}\}$.
-By viewing $T_{\mathrm{BC}}$ as a function of the p-values, we define $T_{\mathrm{BC},i}=T_{\mathrm{BC}}(\mathbf{p}_{-i})$.
-Let $\mathcal{F}_{i}$ be the sigma algebra generated by $\mathbf{p}_{-i}$.
-Then, we have
-$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[w_{\mathrm{BC},i}e_{\mathrm{BC},i}]=$ $\sum_{i\in\mathcal{H}_{0}}w_{\mathrm{BC},i}\mathbb{E}\left[\frac{n\mathbbm{1}\{p_{i}\leq T_{\mathrm{BC}}\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC}}\}}\right]$
-$=$ $\sum_{i\in\mathcal{H}_{0}}w_{\mathrm{BC},i}\mathbb{E}\left[\frac{n\mathbbm{1}\{p_{i}\leq T_{\mathrm{BC},i}\}}{1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}}\right]$
-$=$ $\sum_{i\in\mathcal{H}_{0}}w_{\mathrm{BC},i}\mathbb{E}\left[\frac{n}{1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}}\mathbb{E}[\mathbbm{1}\{p_{i}\leq T_{\mathrm{BC},i}\}\lvert\mathcal{F}_{i}]\right]$
-$\leq$ $\sum_{i\in\mathcal{H}_{0}}w_{\mathrm{BC},i}\mathbb{E}\left[\frac{n}{1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}}\mathbb{E}[\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}\lvert\mathcal{F}_{i}]\right]$
-$=$ $\sum_{i\in\mathcal{H}_{0}}w_{\mathrm{BC},i}\mathbb{E}\left[\frac{n\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}}{1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}}\right],$
-where (i) we have used the fact that $T_{\mathrm{BC}}=T_{\mathrm{BC},i}$ when $p_{i}\leq T_{BC}<0.5$ to get the second equation, (ii) the third equation follows because $T_{\mathrm{BC},i}$ are measurable with respect to $\mathcal{F}_{i}$, and (iii) the inequality is due to the assumption that $p_{i}$ follows the super-uniform distribution on $[0,1]$, and thus it satisfies Condition (8) under the null.
-By Proposition (5), we have
-$\frac{\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}}{1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}}=\frac{\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}}{\sum_{j=1}^{n}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j}\}}.$ (A.2)
-If $p_{i}<1-T_{\mathrm{BC},i}$, both sides are equal to 00.
-If $p_{i}\geq 1-T_{\mathrm{BC},i}$, we claim that $\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}=\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j}\}$.
-Indeed, if $p_{j}\geq 1-T_{\mathrm{BC},i}$ but $p_{j}<1-T_{\mathrm{BC},j}$, we have $T_{\mathrm{BC},i}>T_{\mathrm{BC},j}$.
-This implies that $\min(p_{i},p_{j})\geq 1-\max\{T_{\mathrm{BC},i},T_{\mathrm{BC},j}\}$.
-By Proposition 5, we have $T_{\mathrm{BC},i}=T_{\mathrm{BC},j}$, which contradicts with the fact that $T_{\mathrm{BC},i}>T_{\mathrm{BC},j}$.
-The other direction can be proved similarly.
-Hence we have
-$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[w_{\mathrm{BC},i}e_{\mathrm{BC},i}]\leq n\max_{i\in\mathcal{H}_{0}}w_{\mathrm{BC},i}\mathbb{E}\left[\frac{\sum_{i\in\mathcal{H}_{0}}\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}}{\sum_{j=1}^{n}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j}\}}\right]\leq n\max_{i}w_{\mathrm{BC},i}.$
-By the assumption $w_{\mathrm{BH},i}+\max_{j}w_{\mathrm{BC},j}\leq 1$, we have
-$\mathbb{E}[w_{\mathrm{BH},i}e_{\mathrm{BH},i}+w_{\mathrm{BC},i}e_{\mathrm{BC},i}]\leq\sum_{i=1}^{n}w_{\mathrm{BH},i}+n\max_{j}w_{\mathrm{BC},j}\leq n.$
-A.5 Proof of Theorem 5
+B.1 Proof of Theorem 2
 Proof.
@@
 $=$ $\sum_{i\in\mathcal{H}_{0}}\sum_{k=1}^{n}\mathbb{E}\left[\frac{T_{\mathrm{BH},i}}{T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j,i}\}\right)}\frac{1}{T_{\mathrm{BH}}(k)}\mathbbm{1}\{p_{i}\leq T_{\mathrm{BH}}(k),R_{\mathrm{BH}}(p_{i}\rightarrow 0)=k\}\right],$
-where to get the second equality we have used the fact that when the $i$th hypothesis is rejected (i.e., $p_{i}\leq T_{\mathrm{BH}}$), $R_{\mathrm{BH}}=R_{\mathrm{BH}}(p_{i}\rightarrow 0)$.
+where to get the second equality, we have used the fact that when the $i$th hypothesis is rejected (i.e., $p_{i}\leq T_{\mathrm{BH}}$), $R_{\mathrm{BH}}=R_{\mathrm{BH}}(p_{i}\rightarrow 0)$.
 Let $\mathcal{F}_{i}$ be the sigma algebra generated by $\{p_{1},\cdots,p_{i-1},0,p_{i+1},\cdots,p_{n}\}$.
@@
 $\leq$ $\mathbb{E}\left[\frac{n\max_{i}T_{\mathrm{BH},i}}{\max_{i}T_{\mathrm{BH},i}+\frac{1}{n}\sum_{j=1}^{n}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j}\}}\right].$
-For the BC procedure, let ~$\widetilde{\mathcal{F}}_{i}$ be the sigma algebra generated by $\mathbf{p}_{-i}$. Then, we have
+For the BC procedure, let $\widetilde{\mathcal{F}}_{i}$ be the sigma algebra generated by $\mathbf{p}_{-i}$. Then, we have
 $\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[w_{\mathrm{BC},i}e_{\mathrm{BC},i}]$
@@
 $=$ $\sum_{i\in\mathcal{H}_{0}}\mathbb{E}\left[\frac{\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}\right)}{\max_{i}T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}\right)}\frac{n\mathbbm{1}\{p_{i}\leq T_{\mathrm{BC},i}\}}{1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}}\right]$
-$\leq$ $\sum_{i\in\mathcal{H}_{0}}\mathbb{E}\left[\frac{1}{\max_{i}T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}\right)}\mathbb{E}[\mathbbm{1}\{p_{i}\leq T_{\mathrm{BC},i}\}\lvert\widetilde{\mathcal{F}}_{i}]\right]$
-$=$ $\sum_{i\in\mathcal{H}_{0}}\mathbb{E}\left[\frac{\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}}{\max_{i}T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}\right)}\right],$
-where (i) we have used the fact that $T_{\mathrm{BC}}=T_{\mathrm{BC},i}$ when $p_{i}\leq T_{BC}<0.5$ to get the second equation, (ii) the third equation follows because both $\max_{i}T_{\mathrm{BH},i}$ and $T_{\mathrm{BC},i}$ are measurable with respect to ~$\widetilde{\mathcal{F}}_{i}$, and (iii) the inequality is due to the assumption that $p_{i}$ follows the super-uniform distribution on $[0,1]$ and thus satisfies Condition (8).
-By Proposition 5 and following the same discussion used to prove (A.2), we obtain
-$\frac{\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}}{\max_{i}T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}\right)}=\frac{\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}}{\max_{i}T_{\mathrm{BH},i}+\frac{1}{n}\left(\sum_{j=1}^{n}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j}\}\right)}.$
-Hence we have
+$=$ $\sum_{i\in\mathcal{H}_{0}}\mathbb{E}\left[\frac{1}{\max_{i}T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}\right)}\mathbb{E}[\mathbbm{1}\{p_{i}\leq T_{\mathrm{BC},i}\}\lvert\widetilde{\mathcal{F}}_{i}]\right]$
+$\leq$ $\sum_{i\in\mathcal{H}_{0}}\mathbb{E}\left[\frac{\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}}{\max_{i}T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}\right)}\right],$
+where (i) we have used the fact that $T_{\mathrm{BC}}=T_{\mathrm{BC},i}$ when $p_{i}\leq T_{BC}<0.5$ to get the second equation, (ii) the third equation follows because both $\max_{i}T_{\mathrm{BH},i}$ and $T_{\mathrm{BC},i}$ are measurable with respect to $\widetilde{\mathcal{F}}_{i}$, and (iii) the inequality is due to the assumption that $p_{i}$ follows the super-uniform distribution on $[0,1]$ and thus satisfies Condition (6).
+By Proposition B.1, we have
+$\frac{\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}}{\max_{i}T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}\right)}=\frac{\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}}{\max_{i}T_{\mathrm{BH},i}+\frac{1}{n}\left(\sum_{j=1}^{n}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j}\}\right)}.$ (B.1)
+If $p_{i}<1-T_{\mathrm{BC},i}$, both sides are equal to 0.
+If $p_{i}\geq 1-T_{\mathrm{BC},i}$, we claim that $\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},i}\}=\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j}\}$.
+Indeed, if $p_{j}\geq 1-T_{\mathrm{BC},i}$ but $p_{j}<1-T_{\mathrm{BC},j}$, we have $T_{\mathrm{BC},i}>T_{\mathrm{BC},j}$.
+This implies that $\min(p_{i},p_{j})\geq 1-\max\{T_{\mathrm{BC},i},T_{\mathrm{BC},j}\}$.
+By Proposition B.1, we have $T_{\mathrm{BC},i}=T_{\mathrm{BC},j}$, which contradicts with the fact that $T_{\mathrm{BC},i}>T_{\mathrm{BC},j}$.
+The other direction can be proved similarly.
+Hence, we have
 $\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[w_{\mathrm{BC},i}e_{\mathrm{BC},i}]=$ $\mathbb{E}\left[\frac{\sum_{i\in\mathcal{H}_{0}}\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}}{\max_{i}T_{\mathrm{BH},i}+\frac{1}{n}\left(1+\sum_{j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j}\}\right)}\right]$
@@
 $\sum_{i\in\mathcal{H}_{0}}\{\mathbb{E}[w_{\mathrm{BH},i}e_{\mathrm{BH},i}]+\mathbb{E}[w_{\mathrm{BC},i}e_{\mathrm{BC},i}]\}\leq\mathbb{E}\left[\frac{n\max_{i}T_{\mathrm{BH},i}+\sum_{i=1}^{n}\mathbbm{1}\{p_{i}\geq 1-T_{\mathrm{BC},i}\}}{\max_{i}T_{\mathrm{BH},i}+\frac{1}{n}\sum_{j=1}^{n}\mathbbm{1}\{p_{j}\geq 1-T_{\mathrm{BC},j}\}}\right]=n.$
-A.6 Proof of Theorem 6
+B.2 Proof of Theorem 1
 Proof.
@@
 Let us consider the first group.
-Following the same discussion in the proof of Theorem 4, we have
-$\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}\mathbb{E}[w_{i}e_{i}]=$ $\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}w_{i}\mathbb{E}\left[\frac{n_{1}\mathbbm{1}\{p_{i}\leq T_{1}\}}{1+\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1}\}}\right]$
-$=$ $\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}w_{i}\mathbb{E}\left[\frac{n_{1}\mathbbm{1}\{p_{i}\leq T_{1,i}\}}{1+\sum_{j\in\mathcal{G}_{1},j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{1,i}\}}\right]$
-$\leq$ $\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}w_{i}\mathbb{E}\left[\frac{n_{1}\mathbbm{1}\{p_{i}\geq 1-T_{1,i}\}}{1+\sum_{j\in\mathcal{G}_{1},j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{1,i}\}}\right].$
-By Proposition 5 and the discussion in the proof of Theorem 4, we have
-$\frac{n_{1}\mathbbm{1}\{p_{i}\geq 1-T_{1,i}\}}{1+\sum_{j\in\mathcal{G}_{1},j\neq i}\mathbbm{1}\{p_{j}\geq 1-T_{1,i}\}}=\frac{n_{1}\mathbbm{1}\{p_{i}\geq 1-T_{1,i}\}}{\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1,j}\}}.$
-Thus,
-$\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}\mathbb{E}[w_{i}e_{i}]\leq\max_{i\in\mathcal{G}_{1}}w_{i}\mathbb{E}\left[\frac{n_{1}\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}\mathbbm{1}\{p_{i}\geq 1-T_{1,i}\}}{\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1,j}\}}\right]\leq n_{1}\max_{i\in\mathcal{G}_{1}}w_{i}.$
-Using the same argument for the second group, we obtain
-$\sum_{i\in\mathcal{G}_{2}\cap\mathcal{H}_{0}}\mathbb{E}[w_{i}e_{i}]\leq n_{2}\max_{i\in\mathcal{G}_{2}}w_{i}.$
-Hence, by (17), we deduce that
-$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[w_{i}e_{i}]\leq n,$
-which completes the proof. ∎
-A.7 Proof of Theorem 7
-Proof.
-We only present the proof for the case of $G=2$.
-The arguments can be generalized to the general case without essential difficulty.
-Let us consider the first group.
-Following the same discussion in the proof of Theorem 4, we have
+Let $\mathcal{F}_{i}$ be the sigma algebra generated by $\mathbf{p}_{1,i}$.
+Then, we have
 $\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}\mathbb{E}[w_{i}e_{i}]$
@@
 $=$ $\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}\mathbb{E}\left[\frac{n\mathbbm{1}\{p_{i}\leq T_{1,i}\}}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1,i}\}\right)+\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{p_{j}\geq 1-T_{2,j}\}}\right]$
-$\leq$ $\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}\mathbb{E}\left[\frac{n\mathbbm{1}\{p_{i}\geq 1-T_{1,i}\}}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1,i}\}\right)+\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{p_{j}\geq 1-T_{2,j}\}}\right].$
-By Proposition 5 and the discussion in the proof of Theorem 4, we have
+$=$ $\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}\mathbb{E}\left[\frac{n}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1,i}\}\right)+\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{p_{j}\geq 1-T_{2,j}\}}\mathbb{E}[\mathbbm{1}\{p_{i}\leq T_{1,i}\}|\mathcal{F}_{i}]\right]$
+$\leq$ $\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}\mathbb{E}\left[\frac{n\mathbbm{1}\{p_{i}\geq 1-T_{1,i}\}}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1,i}\}\right)+\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{p_{j}\geq 1-T_{2,j}\}}\right],$
+where (i) we have used the fact that $T_{1}=T_{1,i}$ when $p_{i}\leq T_{1}<0.5$ to get the second equation, (ii) the third equation follows because $T_{1,i}$ are measurable with respect to $\mathcal{F}_{i}$, and (iii) the inequality is due to the assumption that $p_{i}$ satisfies Condition (6) under the null.
+By Proposition B.1 and the proof of (B.1), we have
 $\frac{n\mathbbm{1}\{p_{i}\geq 1-T_{1,i}\}}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1,i}\}\right)+\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{p_{j}\geq 1-T_{2,j}\}}$
@@
 $\sum_{i\in\mathcal{G}_{1}\cap\mathcal{H}_{0}}\mathbb{E}[w_{i}e_{i}]+\sum_{i\in\mathcal{G}_{2}\cap\mathcal{H}_{0}}\mathbb{E}[w_{i}e_{i}]\leq n.$
-A.8 Proof of Theorem 8
+B.3 Proof of Theorem 3
 Proof.
@@
 where we use Assumption 1(C) to get the inequality.
-By Proposition  6 and Assumption 2, we have
+By Proposition  B.2 and Assumption 2, we have
 $\frac{\mathbbm{1}\{\hat{\varphi}_{i}(1-p_{i})\leq T_{1,i}\}}{1+\sum_{j\in\mathcal{G}_{1},j\neq i}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1,i}\}}$ $=\frac{\mathbbm{1}\{\hat{\varphi}_{i}(1-p_{i})\leq T_{1,i}\}}{1+\sum_{j\in\mathcal{G}_{1},j\neq i}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1,j}\}}$
 $=\frac{\mathbbm{1}\{\hat{\varphi}_{i}(1-p_{i})\leq T_{1,i}\}}{\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1,j}\}}.$
-If $\hat{\varphi}_{i}(1-p_{i})>T_{1,i}$, both sides are equal to 00.
+If $\hat{\varphi}_{i}(1-p_{i})>T_{1,i}$, both sides are equal to 0.
 If $\hat{\varphi}_{i}(1-p_{i})\leq T_{1,i}$, we claim that $\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1,i}\}=\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1,j}\}$.
@@
 Hence, $\hat{\varphi}_{i}(1-p_{i})\leq T_{1,i}<T_{1,j}$.
-By proposition 6, we have $T_{1,i}=T_{1,j}$, which contradicts with $T_{1,i}<T_{1,j}$.
+By proposition B.2, we have $T_{1,i}=T_{1,j}$, which contradicts with $T_{1,i}<T_{1,j}$.
 The other direction can be proved similarly.
@@
 $\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[w_{i}]\leq n.$
-A.9 Proof of Theorem 9
+B.4 Proof of Theorem 4
 Proof.
@@
 $\sum_{i\in\mathcal{H}_{0}\cap\mathcal{G}_{1}}\mathbb{E}[w_{i}e_{i}]$
-$=$ $\sum_{i\in\mathcal{H}_{0}\cap\mathcal{G}_{1}}\mathbb{E}\left[\frac{\frac{n}{n_{1}}\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1}\}\right)}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1}\}\right)+\sup_{p\in[0,1]}\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{\hat{\varphi}_{j}^{1,i,p}(1-p_{j})\leq T_{2,j}^{1,i,p}\}}\right.$
-×$\qquad\times\left.\frac{n_{1}\mathbbm{1}\{\hat{\varphi}_{i}(p_{i})\leq T_{1}\}}{1+\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1}\}}\right]$
+$=$ $\sum_{i\in\mathcal{H}_{0}\cap\mathcal{G}_{1}}\mathbb{E}\left[\frac{\frac{n}{n_{1}}\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1}\}\right)}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1}\}\right)+\sup_{p\in[0,1]}\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{\hat{\varphi}_{j}^{1,i,p}(1-p_{j})\leq T_{2,j}^{1,i,p}\}}\right$.\
+$\qquad\times\left.\frac{n_{1}\mathbbm{1}\{\hat{\varphi}_{i}(p_{i})\leq T_{1}\}}{1+\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1}\}}\right]$
 $\leq$ $\sum_{i\in\mathcal{H}_{0}\cap\mathcal{G}_{1}}\mathbb{E}\left[\frac{n\mathbbm{1}\{\hat{\varphi}_{i}(p_{i})\leq T_{1,i}\}}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1,i}\}\right)+\sup_{p\in[0,1]}\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{\hat{\varphi}_{j}^{1,i,p}(1-p_{j})\leq T_{2,j}^{1,i,p}\}}\right].$
@@
 $\leq$ $\sum_{i\in\mathcal{H}_{0}\cap\mathcal{G}_{1}}\mathbb{E}\left[\frac{n\mathbbm{1}\{\hat{\varphi}_{i}(1-p_{i})\leq T_{1,i}\}}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1,i}\}\right)+\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{2,j}\}}\right],$
-where we have used Assumption 1(C) to obtain the first inequality and the second inequality is due to the fact that
+where we have used Assumption 1(C) to obtain the first inequality, and the second inequality is due to the fact that
 $\sum_{j\in G_{2}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{2,j}\}$ $=\sum_{j\in G_{2}}\mathbbm{1}\{\hat{\varphi}_{j}^{1,i,p}(1-p_{j})\leq T_{2,j}^{1,i,p}\}\lvert_{p=p_{i}}$
 $\leq\sup_{p\in[0,1]}\sum_{j\in G_{2}}\mathbbm{1}\{\hat{\varphi}_{j}^{1,i,p}(1-p_{j})\leq T_{2,j}^{1,i,p}\}.$
-By Proposition 6 and the argument in the proof of Theorem 8, we have
+By Proposition B.2 and the argument in the proof of Theorem 3, we have
 $\frac{n\mathbbm{1}\{\hat{\varphi}_{i}(1-p_{i})\leq T_{1,i}\}}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1,i}\}\right)+\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{2,j}\}}$
@@
 $\sum_{i\in\mathcal{H}_{0}\cap\mathcal{G}_{2}}\mathbb{E}[w_{i}e_{i}]\leq\mathbb{E}\left[\frac{n\sum_{i\in\mathcal{G}_{2}}\mathbbm{1}\{\hat{\varphi}_{i}(1-p_{i})\leq T_{1,i}\}}{\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1,j}\}+\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{2,j}\}}\right].$
-Combing the above results leads to
+Combining the above results leads to
 $\sum_{g=1}\sum_{i\in\mathcal{H}_{0}\cap\mathcal{G}_{g}}\mathbb{E}[w_{i}e_{i}]\leq n.$
-Appendix B Additional proofs
-B.1 Proof of Proposition 1
+Appendix C Supplementary Proofs
+C.1 Proof of Proposition 1
 Proof.
 Note that
-[…]
+$\text{FDP}=$ $\sum_{i=1}^{n}\frac{\mathbbm{1}\{ie_{(i)}\geq n/\alpha,H_{(i)}\text{ is under the null}\}}{1\vee\hat{k}}$
 $\leq$ $\sum_{i=1}^{n}\frac{\mathbbm{1}\{ie_{(i)}\geq n/\alpha,H_{(i)}\text{ is under the null}\}}{1\vee i}$
@@
 $\text{FDR}=\mathbb{E}[\text{FDP}]\leq\alpha.$
-B.2 Proof of Proposition 2
+C.2 Proof of Proposition B.1
 Proof.
-Let $V_{i}=\mathbbm{1}\{H_{i}\text{ is rejected}\}$. We have
-$\mathrm{FDP}(T_{\mathrm{GBH}})=\sum_{i\in\mathcal{H}_{0}}\frac{V_{i}}{R(T_{\mathrm{GBH}})\vee 1}=\sum_{i\in\mathcal{H}_{0}}\frac{V_{i}}{\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}})}\frac{\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}})}{R(T_{\mathrm{GBH}})\vee 1}\leq\alpha\sum_{i\in\mathcal{H}_{0}}\frac{V_{i}}{\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}})}.$
-Therefore, it suffices to prove
-$\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}}\frac{V_{i}}{\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}})}\right]\leq 1.$
-Observing that, for a given $R$, $T_{\mathrm{GBH}}=T_{\mathrm{GBH}}(R)$ is a deterministic function of $R$, we have:
-$\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}}\frac{V_{i}}{\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}})}\right]=\sum_{i\in H_{0}}\sum_{k=1}^{n}\mathbb{E}\left[\frac{V_{i}\mathbbm{1}\{R=k\}}{\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}}(k))}\right]=\sum_{i\in H_{0}}\sum_{k=1}^{n}\mathbb{E}\left[\frac{V_{i}\mathbbm{1}\{R(p_{i}\rightarrow 0)=k\}}{\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}}(k))}\right],$
-where $R(p_{i}\rightarrow 0)$ is the number of rejections obtained by replacing the p-value $p_{i}$ with 0.
-To clarify the second equality, note that if $V_{i}=0$, the equation is trivially true.
-When $V_{i}=1$, setting $p_{i}$ to 00 does not change the number of rejections.
-Using the assumption that $F_{j}(x)=c_{j}h(x)$, we obtain:
-$\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}}\frac{V_{i}}{\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}})}\right]$ $=\sum_{i\in\mathcal{H}_{0}}\sum_{k=1}^{n}\frac{1}{\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}}(k))}\mathbb{E}[\mathbbm{1}\{\varphi_{i}(p_{i})\leq T_{\mathrm{GBH}}(k)\}]\mathbb{E}[\mathbbm{1}\{R(p_{i}\rightarrow 0)=k\}]$
-$\leq\sum_{i\in\mathcal{H}_{0}}\sum_{k=1}^{n}\frac{F_{i}(T_{\mathrm{GBH}}(k))}{\sum_{j=1}^{n}F_{j}(T_{\mathrm{GBH}}(k))}\mathbb{E}[\mathbbm{1}\{R(p_{i}\rightarrow 0)=k\}]$
-$=\sum_{i\in\mathcal{H}_{0}}\sum_{k=1}^{n}\frac{c_{i}}{\sum_{j=1}^{n}c_{j}}\mathbb{E}[\mathbbm{1}\{R(p_{i}\rightarrow 0)=k\}]$
-$=\frac{\sum_{i\in\mathcal{H}_{0}}c_{i}}{\sum_{j=1}^{n}c_{j}}\leq 1.$
-B.3 Proof of Proposition 3
+Proposition B.1 is a special case of Proposition B.2 by choosing $\varphi_{i}$ as the identity function for all $1\leq i\leq n$. ∎
+C.3 Proof of Proposition B.2
 Proof.
-Recall from the proof of Proposition 2 that $T_{\mathrm{GBH}}=T_{\mathrm{GBH}}(R_{\mathrm{GBH}})$ is a deterministic function of $R_{\mathrm{GBH}}$.
-Additionally, for any $i\in\mathcal{H}_{0}$ and $\varphi_{i}(p_{i})\leq T_{\mathrm{GBH}}$, replacing the p-value $p_{i}$ with 0 does not change the number of rejections, i.e., $R_{\mathrm{GBH}}(p_{i}\leftarrow 0)=R_{\mathrm{GBH}}$.
-Therefore, we can infer that
-$\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}}e_{i}\right]=$ $\sum_{i\in\mathcal{H}_{0}}\sum^{n}_{k=1}\mathbb{E}\left[\frac{n\mathbbm{1}\{\varphi_{i}(p_{i})\leq T_{\mathrm{GBH}},R_{\mathrm{GBH}}=k\}}{\sum^{n}_{j=1}F_{j}(T_{\mathrm{GBH}})}\right]$
-$=$ $\sum_{i\in\mathcal{H}_{0}}\sum^{n}_{k=1}\frac{n\mathbb{E}[\mathbbm{1}\{\varphi_{i}(p_{i})\leq T_{\mathrm{GBH}}(k),R_{\mathrm{GBH}}=k\}]}{\sum^{n}_{j=1}F_{j}(T_{\mathrm{GBH}}(k))}$
-$=$ $\sum_{i\in\mathcal{H}_{0}}\sum^{n}_{k=1}\frac{n\mathbb{E}[\mathbbm{1}\{\varphi_{i}(p_{i})\leq T_{\mathrm{GBH}}(k),R_{\mathrm{GBH}}(p_{i}\leftarrow 0)=k\}]}{\sum^{n}_{j=1}F_{j}(T_{\mathrm{GBH}}(k))}$
-$\leq$ $\sum_{i\in\mathcal{H}_{0}}\sum^{n}_{k=1}\frac{nF_{i}(T_{\mathrm{GBH}}(k))\mathbb{E}[\mathbbm{1}\{R_{\mathrm{GBH}}(p_{i}\leftarrow 0)=k\}]}{\sum^{n}_{j=1}F_{j}(T_{\mathrm{GBH}}(k))}$
-$=$ $\sum_{i\in\mathcal{H}_{0}}\sum_{k=1}^{n}\frac{nc_{i}h(T_{\mathrm{GBH}}(k))}{\sum_{j=1}^{n}c_{j}h(T_{\mathrm{GBH}}(k))}\mathbb{E}[\mathbbm{1}\{R_{\mathrm{GBH}}(p_{i}\leftarrow 0)=k\}]$
-$=$ $n\frac{\sum_{i\in\mathcal{H}_{0}}c_{i}}{\sum_{j=1}^{n}c_{j}}\leq n,$
-which completes the proof. ∎
-B.4 Proof of Proposition 4
-Write $T=T_{\mathrm{GBC}}$ for the ease of notation. First note that
-$\mathbb{E}\left[\sum_{i\in\mathcal{H}_{0}}\frac{\mathbbm{1}\{\varphi_{i}(p_{i})\leq T\}}{1\vee\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(p_{j})\leq T\}}\right]$
-$=$ $\sum_{i\in\mathcal{H}_{0}}\mathbb{E}\left[\frac{\mathbbm{1}\{\varphi_{i}(p_{i})\leq T\}}{1\vee\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(p_{j})\leq T\}}\frac{1+\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T\}}\right]$
-$\leq$ $\alpha\sum_{i\in\mathcal{H}_{0}}\mathbb{E}\left[\frac{\mathbbm{1}\{\varphi_{i}(p_{i})\leq T\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T\}}\right].$
-Hence, we only need to show that
-$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}\left[\frac{\mathbbm{1}\{\varphi_{i}(p_{i})\leq T\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T\}}\right]\leq 1.$
-One approach to prove FDR control is through the construction of a super-martingale and the use of the optional stopping time theorem.
-Here we employ an alternative argument based on the leave-one-out technique.
-Let $\tilde{p}_{i}=\min\{p_{i},1-p_{i}\}$ and $\mathbf{p}_{-i}=(p_{1},\dots,p_{i-1},\tilde{p}_{i},p_{i+1},\dots,p_{n})$.
-Define $T_{i}=T(\mathbf{p}_{-i})$, where we view $T$ as a function of the p-values.
-Notice that if $\varphi_{i}(p_{i})\leq T$, then we have
-$\varphi_{i}(p_{i})\leq T\leq T_{\mathrm{up}}<\varphi_{i}(0.5),$
-which implies that $p_{i}<0.5$ since $\varphi_{i}$ is increasing.
-Hence, if the $i$th hypothesis is rejected, then $p_{i}=\tilde{p}_{i}$.
-Thus, $\mathbbm{1}\{\varphi_{i}(p_{i})\leq T\}=\mathbbm{1}\{\varphi_{i}(p_{i})\leq T_{i}\}$, which further implies that
-$\mathbb{E}\left[\frac{\mathbbm{1}\{\varphi_{i}(p_{i})\leq T\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T\}}\right]$ $=\mathbb{E}\left[\frac{\mathbbm{1}\{\varphi_{i}(p_{i})\leq T_{i}\}}{1+\sum_{j\neq i}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{i}\}}\right],$
-where we use the fact that if $p_{i}<0.5$, then $\varphi_{i}(1-p_{i})\geq\varphi_{i}(0.5)>T_{\mathrm{up}}\geq T_{i}$.
-Let $\mathcal{F}_{i}$ be the sigma algebra generated by $\mathbf{p}_{-i}$.
-For $i\in\mathcal{H}_{0}$, we have
-$\mathbb{E}\left[\frac{\mathbbm{1}\{\varphi_{i}(p_{i})\leq T\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T\}}\right]=$ $\mathbb{E}\left[\mathbb{E}\left[\frac{\mathbbm{1}\{\varphi_{i}(p_{i})\leq T_{i}\}}{1+\sum_{j\neq i}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{i}\}}\Bigg{|}\mathcal{F}_{i}\right]\right]$
-$=$ $\mathbb{E}\left[\frac{1}{1+\sum_{j\neq i}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{i}\}}\mathbb{E}\left[\mathbbm{1}\{\varphi_{i}(p_{i})\leq T_{i}\}\lvert\mathcal{F}_{i}\right]\lvert\right]$
-$\leq$ $\mathbb{E}\left[\frac{1}{1+\sum_{j\neq i}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{i}\}}\mathbb{E}\left[\mathbbm{1}\{\varphi_{i}(1-p_{i})\leq T_{i}\}\lvert\mathcal{F}_{i}\right]\lvert\right]$
-$=$ $\mathbb{E}\left[\frac{\mathbbm{1}\{\varphi_{i}(1-p_{i})\leq T_{i}\}}{1+\sum_{j\neq i}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{i}\}}\right],$
-where we use the assumption that $p_{i}$ satisfies Condition (8) to get the inequality. By Proposition  6, we have
-$\frac{\mathbbm{1}\{\varphi_{i}(1-p_{i})\leq T_{i}\}}{1+\sum_{j\neq i}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{i}\}}=\frac{\mathbbm{1}\{\varphi_{i}(1-p_{i})\leq T_{i}\}}{1+\sum_{j\neq i}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{j}\}}=\frac{\mathbbm{1}\{\varphi_{i}(1-p_{i})\leq T_{i}\}}{\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{j}\}}.$
-If $\varphi_{i}(1-p_{i})>T_{i}$, both sides are equal to 00.
-If $\varphi_{i}(1-p_{i})\leq T_{i}$, we claim that $\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{i}\}=\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{j}\}$.
-Indeed, if $\varphi_{j}(1-p_{j})>T_{i}$ but $\varphi_{j}(1-p_{j})\leq T_{j}$, then we have $T_{i}<T_{j}$.
-Hence, $\varphi_{i}(1-p_{i})\leq T_{i}<T_{j}$.
-By proposition 6, we have $T_{i}=T_{j}$, which contradicts with the assumption $T_{i}<T_{j}$.
-The other direction can be proved similarly.
-Hence,
-$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}\left[\frac{\mathbbm{1}\{\varphi_{i}(p_{i})\leq T\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T\}}\right]\leq\mathbb{E}\left[\frac{\sum_{i\in\mathcal{H}_{0}}\mathbbm{1}\{\varphi_{i}(1-p_{i})\leq T_{i}\}}{\sum_{j=1}^{n}\mathbbm{1}\{\varphi_{j}(1-p_{j})\leq T_{j}\}}\right]\leq 1,$ (B.1)
-which finishes the proof.
-B.5 Proof of Proposition 5
-Proof.
-Proposition 5 is a special case of Proposition 6 by choosing $\varphi_{i}$ as the identity function for all $1\leq i\leq n$. ∎
-B.6 Proof of Proposition 6
-Proof.
-Write $T=T_{\mathrm{GBC}}$ for the ease of notation. First, given a p-value vector $\mathbf{p}=(p_{1},\cdots,p_{n})$, recall that the threshold $T$ is defined as
+Write $T=T_{\mathrm{FBC}}$ for the ease of notation. First, given a p-value vector $\mathbf{p}=(p_{1},\cdots,p_{n})$, recall that the threshold $T$ is defined as
 $T=\max\left\{0<t\leq T_{\mathrm{up}}\colon\underbrace{\frac{1+\sum_{l=1}^{n}\mathbbm{1}\{\varphi_{l}(1-p_{l})\leq t\}}{\sum_{l=1}^{n}\mathbbm{1}\{\varphi_{l}(p_{l})\leq t\}}}_{\coloneqq g(\mathbf{p},t)}\leq\alpha\right\},$
@@
 By the assumption that $\max\{\varphi_{i}(1-p_{i}),\varphi_{j}(1-p_{j})\}\leq\max\{T_{i},T_{j}\}$, we have $\varphi_{i}(1-p_{i})\leq T_{i}$ and $\varphi_{j}(1-p_{j})\leq T_{i}$.
-Since $\varphi_{i}$ is an increasing function, we have $\varphi_{i}(1-p_{i})\leq T_{\mathrm{up}}<\varphi_{i}(0.5)$, which implies $0.51-p_{i}<$ 0.5.
+Since $\varphi_{i}$ is an increasing function, we have $\varphi_{i}(1-p_{i})\leq T_{\mathrm{up}}<\varphi_{i}(0.5)$, which implies $1-p_{i}<0.5$.
 Thus $\varphi_{i}(p_{i})\geq\varphi_{i}(0.5)>T_{\mathrm{up}}\geq T_{i}$.
@@
 Hence, $g(\mathbf{p}_{-j},T_{i})=g(\mathbf{p}_{-i},T_{i})\leq\alpha$. By the definition of $T_{j}$, we must have $T_{i}\leq T_{j}$. Similarly, we get $T_{j}\leq T_{i}$ and hence $T_{i}=T_{j}$. ∎
-Appendix C Additional numerical results
-C.1 Additional numerical results for fair multiple testing
-[…]
+Appendix D Assembling E-Values from Data Subsets
+We first compare our definition, which simultaneously controls group-wise and overall FDR, with the notion of predictive parity from the classification context in the fairness literature [Chouldechova, 2017].
+To elaborate, consider a binary classification problem where $Y\in{0,1}$ represents the true labels and $\widehat{Y}\in{0,1}$ denotes the predicted labels.
+In this setting, the FDR is defined as $P(Y=0|\widehat{Y}=1)$, and predictive parity requires equal FDR across all groups.
+However, the multiple testing scenario differs fundamentally from the classification problem, as the underlying truth of each hypothesis is unobserved and thus cannot directly inform the decision rule.
+Hence, in our context, it is more appropriate to control the FDRs across different groups at a common target level rather than enforcing strict equality.
+D.1 Additional Numerical Results: Group-Wise and Overall FDR Control
+Example: Data-Dependent Weights
+We begin with a concrete example to illustrate the form of the data-dependent weights. When $L=2$, we have
+$w_{i}=\frac{\frac{n}{n_{1}}\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1}\}\right)}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1}\}\right)+\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{p_{j}\geq 1-T_{2,j}\}}$
+for $i\in\mathcal{G}_{1}$ and
+$w_{i}=\frac{\frac{n}{n_{2}}\left(1+\sum_{j\neq i,j\in\mathcal{G}_{2}}\mathbbm{1}\{p_{j}\geq 1-T_{2}\}\right)}{\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{p_{j}\geq 1-T_{1,j}\}+\left(1+\sum_{j\neq i,j\in\mathcal{G}_{2}}\mathbbm{1}\{p_{j}\geq 1-T_{2}\}\right)}$
+for $i\in\mathcal{G}_{2}$.
+A Naive Method that Controls both Group-wise and Overall FDR
+We tried BC_Sep at the level $\alpha/G$, a naive method that controls both the group-wise and overall FDRs.
+Indeed, let $\hat{n}_{la}$ be the number of rejections for the $l$th group, and denote the number of false rejections for the $l$th group by $\hat{n}_{l0}$.
+Then we have $\mathbb{E}[\hat{n}_{l0}/(1\vee\hat{n}_{la})]\leq\alpha/L$, which implies that
+$\mathbb{E}\left[\frac{\sum_{l=1}^{L}\hat{n}_{l0}}{1\vee\sum_{l=1}^{L}\hat{n}_{la}}\right]\leq\mathbb{E}\left[\sum_{l=1}^{L}\frac{\hat{n}_{l0}}{1\vee\hat{n}_{la}}\right]\leq\alpha.$
+However, this method has nearly zero power in all our simulation settings. Therefore, we have decided not to include its results in Table 1.
+Parameter Values for the Two-Group Setting
+The parameter values for different settings for two groups are detailed in Table D.1.
+Setting $n_{1}$ $n_{1a}$ $\alpha_{1}$ $\beta_{1}$ $n_{2}$ $n_{2a}$ $\alpha_{2}$ $\beta_{2}$
+E1 $100$ 2020 44 $500$ $1000$ 2020 $0.1$ $500$
+E2 $100$ 2020 $0.5$ $500$ $1000$ 2020 $0.5$ $500$
+Table D.1: Parameter settings for the case of $G=2$.
+Here, $n_{g}$ represents the number of hypotheses for the $g$th group; $n_{ga}$ denotes the number of non-null hypotheses in the $g$th group with $g=1,2$.
+$\alpha_{g}$ and $\beta_{g}$ are the parameters of the beta distribution for the p-values under the alternatives for the $g$th group.
+Numerical Results for the Four-Group Setting
+We also consider the case of $G=4$.
+To evaluate the performance of each method, we employ the following metrics: POW represents the overall power combining the rejections from all four groups; POWg denotes the power for the $g$th group with $1\leq g\leq 4$.
+Similarly, we can define FDR and FDRg.
+The empirical power and FDR are computed based on 1,000 independent Monte Carlo simulations.
 In all settings, we assume that the p-values follow the uniform distribution on $[0,1]$ under the null.
-For the $g$th group, the p-value is assumed to follow $\mathrm{Unif}[0,u_{g}]$ under the alternatives.
-The parameter values for different settings are detailed in Table C.2.
-Table C.2 displays the results for Setting F1.
-It can be seen that BC_Com fails to control the FDR for the third and fourth groups with the empirical FDR reaching 0.073 compared to the 5% target level.
-BC_Sep has an empirical FDR being 0.064, higher than the nominal level.
-The results for Setting F2 are presented in Table C.4.
+For the $g$th group, the p-value is assumed to follow Beta($\alpha_{g},\beta_{g}$) under the alternatives.
+The parameter values for different settings are detailed in Table D.3.
+Table D.3 displays the results for Setting F1.
+It can be seen that BC_Com fails to control the FDR for the third and fourth groups, with the empirical FDR reaching 0.073 compared to the 5% target level.
+BC_Sep has an empirical FDR of 0.064, higher than the nominal level.
+The results for Setting F2 are presented in Table D.5.
 We note that BC_Com suffers from a severe FDR inflation with the empirical FDR being 0.312 at the 5% target level.
 In Setting F3, we raise the target FDR level to $20\%$.
-As seen from Table C.4, BC_Sep significantly inflates the overall FDR with the empirical FDR being 0.346.
-Throughout all settings, the e-BH based approach controls both the group-wise FDR and the overall FDR.
+As seen from Table D.5, BC_Sep significantly inflates the overall FDR, with the empirical FDR being 0.346.
+Throughout all settings, the e-BH-based approach controls both the group-wise FDR and the overall FDR.
 Furthermore, eBH_Ada outperforms both eBH_1 and eBH_2 in terms of power.
-C.2 Additional numerical results for structure adaptive multiple testing
-The results for $a_{f}=0.5$ are presented in Figure C.1.
-When the signal is sparse, and the covariate is less informative, slight FDR inflation is observed in SABHA and the two versions of IHW.
-eBH_GBC has the highest power, followed by SABHA and the two versions of IHW.
+Setting Target FDR level $n_{1}$ $n_{1a}$ $\alpha_{1}$ $\beta_{1}$ $n_{2}$ $n_{2a}$ $\alpha_{2}$ $\beta_{2}$ $n_{3}$ $n_{3a}$ $\alpha_{3}$ $\beta_{3}$ $n_{4}$ $n_{4a}$ $\alpha_{4}$ $\beta_{4}$
+Setting F1 0.05 100 20 $0.1$ $500$ 100 20 $0.1$ $500$ 1000 20 $0.1$ $500$ 1000 20 $0.1$ $500$
+Setting F2 0.05 100 1 $0.01$ 5000 100 20 $0.1$ 500 100 20 $0.1$ 500 100 20 $0.1$ $500$
+Setting F3 0.2 50 2 $0.1$ 500 100 2 $0.1$ 500 50 4 $0.2$ 500 100 4 $0.3$ 500
+Table D.2: Parameter settings for the case of $G=4$.
+Here, $n_{g}$ represents the number of hypotheses for the $g$th group; $n_{ga}$ denotes the number of non-null hypotheses in the $g$th group.
+$\alpha_{g}$ and $\beta_{g}$ are parameters of the beta distribution for the p-values under the alternatives for the $g$th group.
+Method POW POW1 POW2 POW3 POW4 FDR FDR1 FDR2 FDR3 FDR4
+BC_Com 0.954 0.955 0.956 0.952 0.953 0.045 0.007 0.006 0.078 0.077
+BC_Sep 0.649 0.895 0.865 0.411 0.428 0.064 0.049 0.042 0.035 0.04
+eBH_1 0.029 0 0 0.057 0.057 0.008 0 0 0.007 0.008
+eBH_2 0.14 0.142 0.142 0.139 0.138 0.01 0.007 0.007 0.011 0.013
+eBH_a 0.222 0.256 0.258 0.179 0.194 0.017 0.013 0.012 0.015 0.018
+Table D.3: FDR and power for Setting F1, where the nominal FDR level is 5%.
+Method POW POW1 POW2 POW3 POW4 FDR FDR1 FDR2 FDR3 FDR4
+BC_Com 0.988 0.999 0.994 0.987 0.983 0.046 0.318 0.034 0.033 0.032
+BC_Sep 0.793 0 0.863 0.784 0.771 0.055 0 0.045 0.043 0.042
+eBH_1 0 0 0 0 0 0 0 0 0 0
+eBH_2 0 0 0 0 0 0 0 0 0 0
+eBH_a 0.53 0 0.544 0.537 0.536 0.031 0 0.028 0.03 0.03
+Table D.4: FDR and power for Setting F2, where the nominal FDR level is 5%.
+Method POW POW1 POW2 POW3 POW4 FDR FDR1 FDR2 FDR3 FDR4
+BC_Com 0.983 0.992 0.991 0.984 0.975 0.181 0.154 0.249 0.089 0.161
+BC_Sep 0.393 0.149 0.141 0.517 0.517 0.343 0.1 0.094 0.164 0.176
+eBH_1 0.002 0 0.003 0 0.003 0.002 0 0.002 0 0.001
+eBH_2 0.004 0.005 0 0.005 0.005 0.003 0.003 0 0.001 0.002
+eBH_a 0.035 0.013 0.006 0.049 0.047 0.019 0.008 0.004 0.016 0.015
+Table D.5: FDR and power for Setting F3, where the nominal FDR level is 20%.
+D.2 Numerical Studies for the Real Dataset
+We illustrate the proposed method by conducting differential abundance analysis using the microbiome dataset cdi_schubert, sourced from the MicrobiomeHD repository Duvallet et al.
+[2017], originally collected in a case-control study comparing individuals with Clostridium difficile infection (CDI) to those without (nonCDI) Schubert et al.
+[2014].
+This dataset comprises 336 microbiome samples.
+Each sample is classified as either CDI (infection case) or non-CDI (healthy control).
+The raw feature table contains a total of 19,314 operational taxonomic units (OTUs), representing bacterial taxa annotated at the phylum level.
+Before analysis, several filtering and preprocessing steps were applied to refine the OTU table.
+Initially, OTUs lacking phylum annotations were excluded.
+Subsequently, entire phylum groups containing fewer than 200 taxa were removed, restricting analysis to well-represented bacterial groups.
+Following this step, features from three major bacterial phyla remained: Bacteroidetes, Firmicutes, and Proteobacteria.
+Additionally, we implemented prevalence-based quality control by filtering out taxa detected in fewer than 10 samples.
+After applying these criteria, the final dataset retained 2293 microbiome features across all 336 samples, ensuring robust and informative taxa for downstream analyses.
+We then performed differential abundance testing to identify taxa differing between the diarrheal case and control groups.
+Specifically, we utilized the LinDA method [Zhou et al., 2022], which fits a log-linear model to compositional microbiome data, adjusting for sequencing depth and compositional bias.
+This method generated a p-value for each taxon, assessing differences in abundance between cases and controls.
+The resulting collection of p-values was subsequently processed using multiple-testing correction procedures.
+In particular, we applied the BC_Com method and three eBH-based methods (eBH_1, eBH_2, and eBH_Ada), as proposed before, with a target FDR of $\alpha=0.2$.
+The number of rejected taxa for each phylum is summarized in Table D.6.
+method/phylum Bacteroidetes Firmicutes Proteobacteria
+BC_Com 354 515 106
+eBH_1 0 0 0
+eBH_2 0 0 0
+eBH_Ada 259 557 175
+Table D.6: Numbers for rejections for each method and phylum.
+In this study, controlling the overall FDR ensures the reliability of global inference, whereas controlling the FDR within each phylum is essential for accurately interpreting results within biologically meaningful groups.
+The results in Table D.6 show that for the phyla Bacteroidetes and Firmicutes, the eBH_Ada method yields fewer rejections compared to the BC_Com method, suggesting that BC_Com might inadequately control FDR within these specific groups.
+Conversely, for the phylum Proteobacteria, the eBH_Ada methods identify a greater number of rejections, mirroring the pattern observed in our simulation scenario E1, where the BC_Com method exhibits reduced power in certain groups.
+Additionally, the two data-independent weighting methods, eBH_1 and eBH_2, have no discovery, underscoring the practical necessity of employing data-dependent weights when combining e-values.
+Appendix E Hybrid Knockoff Procedure
+In this section, we demonstrate how e-values can be used to combine results obtained from different test statistics.
+The knockoff method [Barber and Candès, 2015, Candes et al., 2018] provides a variable selection framework designed to control the FDR at a specified level.
+Typically, knockoff methods utilize the Lasso coefficient-difference statistic; however, when the relationship between response and regressors is non-linear, statistics based on random forests become preferable [Candes et al., 2018].
+Given that the true dependence structure between the response and regressors is usually unknown in practice, we propose a hybrid knockoff approach that combines results from multiple test statistics, enhancing robustness across different modeling scenarios.
+Specifically, motivated by the recent work of Ren and Barber [2024], which establishes the equivalence between the knockoff procedure and the e-BH procedure under certain e-values, our approach first transforms results from multiple knockoff statistics into corresponding e-values.
+These e-values are then aggregated via arithmetic mean, and the e-BH procedure is subsequently applied to determine the final rejection set.
+A detailed exposition of the knockoff framework and the proposed hybrid method is provided in the subsequent sections.
+E.1 Connection between Knockoff and e-BH Procedures
+In Ren and Barber [2024], the authors demonstrated that the knockoff method is equivalent to the e-BH method under a specific form of e-values.
+In this subsection, we briefly review this result for later use.
+In variable selection, the goal is to identify predictors that are significantly associated with the response variable.
+A predictor is considered a null variable if it is conditionally independent of the response given all other predictors.
+Formally, let $Y$ denote the response variable.
+For a predictor $X_{j}$, with the remaining predictors denoted by $X_{-j}=\{X_{i}:1\leq i\leq p,\,i\neq j\}$, $X_{j}$ is a null variable if
+$Y\mathrel{\perp\!\!\!\perp}X_{j}\mid X_{-j}.$
+Now, consider the linear model
+$\mathbf{Y}=\mathbf{X}\bm{\beta}+\bm{\epsilon},$
+[…]
+$\text{FDR}=\mathbb{E}\left[\frac{\#\{i:\beta_{i}=0,\,i\in\hat{S}\}}{1\vee|\hat{S}|}\right$].
+The knockoff method [Barber and Candès, 2015, Candes et al., 2018] provides a variable selection approach that controls the FDR at a desired level when $n>p$.
+The key idea is to construct synthetic predictors, $\tilde{\mathbf{X}}$, called knockoffs, which preserve the correlation structure of the original features $\mathbf{X}$ while remaining conditionally independent of the response $\mathbf{Y}$.
+Specifically, the knockoff procedure generates a knockoff copy $\tilde{\mathbf{X}}$ that satisfies
+$\mathbf{Y}\mathrel{\perp\!\!\!\perp}\tilde{\mathbf{X}}\mid\mathbf{X}$
+and
+$(\mathbf{X}_{j},\tilde{\mathbf{X}}_{j},\mathbf{X}_{-j},\tilde{\mathbf{X}}_{-j})\overset{d}{=}(\tilde{\mathbf{X}}_{j},\mathbf{X}_{j},\tilde{\mathbf{X}}_{-j},\mathbf{X}_{-j}),$
+for each $j$, where $\overset{d}{=}$ denotes equality in distribution. For details on constructing $\tilde{\mathbf{X}}$, see Barber and Candès [2015], Candes et al. [2018], Barber et al. [2020].
+Once the knockoffs $\tilde{\mathbf{X}}$ are constructed, feature importance statistics are computed using the augmented data $([\tilde{\mathbf{X}},\mathbf{X}],Y)$ by
+$\mathbf{W}=\mathcal{W}([\tilde{\mathbf{X}},\mathbf{X}],\mathbf{Y}),$
+where $\mathcal{W}(\cdot)$ is an algorithm that quantifies the importance of each feature.
+A key property of $\mathbf{W}=[W_{1},\dots,W_{p}]^{\top}$ is that swapping $\mathbf{X}_{j}$ and $\tilde{\mathbf{X}}_{j}$ reverses the sign of $W_{j}$, and larger values of $W_{j}$ provide stronger evidence against the null hypothesis for the $j$th predictor.
+For a predetermined FDR level $\alpha$, the knockoff threshold $T$ is defined as
+$T=\min\left\{t\in\mathbb{W}\colon\frac{1+\sum_{j=1}^{p}\mathbbm{1}\{W_{j}\leq-t\}}{1\vee\sum_{j=1}^{p}\mathbbm{1}\{W_{j}\geq t\}}\leq\alpha\right\},$
+where $\mathbb{W}=\{|W_{j}|\colon i=1,2,\ldots,p\}\setminus\{0\}$. The final selected model is given by
+$\hat{S}=\{j:W_{j}\geq T\}.$
+The FDR control properties of the knockoff method are detailed in Barber and Candès [2015], Barber et al. [2020].
+In Ren and Barber [2024], the authors demonstrated that the e-BH procedure, using the e-value defined for the $i$th hypothesis as
+$e_{i}\coloneqq\frac{p\mathbbm{1}\{W_{i}\geq T\}}{1+\sum_{j=1}^{p}\mathbbm{1}\{W_{j}\leq-T\}},$ (E.1)
+is equivalent to the knockoff method. Specifically, the two procedures yield identical rejection sets.
+E.2 Hybrid Knockoff Algorithm
+In knockoff methods, various test statistics can be utilized, and the optimal choice generally depends on the underlying data-generating mechanism.
+Since the true relationship between $Y$ and $X$ is unknown in practice, selecting a suitable test statistic in advance is challenging.
+In this section, we propose a hybrid knockoff approach that combines results from different test statistics, ensuring robust performance under various scenarios.
+One of the most commonly used test statistics is the Lasso coefficient-difference statistic proposed by Candes et al.
+[2018].
+Specifically, one fits a cross-validated Lasso regression to the augmented design matrix $\mathbf{X},\tilde{\mathbf{X}}$] to predict $\mathbf{Y}$.
+Denoting by $\beta_{j}$ and $\tilde{\beta}_{j}$ the fitted coefficients corresponding to the original feature $\mathbf{X}_{j}$ and its knockoff $\tilde{\mathbf{X}}_{j}$, respectively, the test statistic is defined as
+$W_{j}=|\beta_{j}|-|\tilde{\beta}_{j}|,\quad j=1,\dots,p.$
+Intuitively, this statistic performs well when the true relationship between $\mathbf{Y}$ and $\mathbf{X}$ is approximately linear [Candes et al., 2018].
+However, if this linear assumption does not hold, alternative statistics based on nonlinear models, such as random forests, can be considered.
+For example, one can define $W_{j}$ as the difference in feature importance scores between the original feature $\mathbf{X}_{j}$ and its knockoff counterpart $\tilde{\mathbf{X}}_{j}$ computed by a random forest model [Candes et al., 2018].
+Since we do not know the true dependence structure between $\mathbf{Y}$ and $\mathbf{X}$, it is beneficial to develop a unified procedure that performs consistently well regardless of the underlying model form.
+Leveraging the concept of e-values, we propose combining results from multiple knockoff test statistics into a single robust procedure.
+Specifically, we first compute knockoff statistics separately, using both Lasso-based and random-forest-based approaches, and then transform these statistics into e-values.
+These individual e-values are subsequently combined into a single e-value vector, after which the e-BH procedure is applied to determine the final rejection set.
+Formally, let $\mathbf{W}_{1}=[W_{1,1},\dots,W_{1,p}]^{\top}$ and $\mathbf{W}_{2}=[W_{2,1},\dots,W_{2,p}]^{\top}$ denote two distinct sets of knockoff test statistics.
+First, we implement the knockoff methods using these two sets of statistics at the target FDR level $\alpha_{\mathrm{ko}}$, obtaining thresholds $T_{1}$ and $T_{2}$.
+Following the discussion in Ren and Barber [2024], a standard choice is $\alpha_{\mathrm{ko}}=\alpha_{\mathrm{eBH}}/2$, where $\alpha$ is the target FDR level used in the e-BH procedure.
+Subsequently, we apply equation (E.1) to compute individual e-value vectors $\mathbf{e}_{1}$ and $\mathbf{e}_{2}$ based on $(\mathbf{W}_{1},T_{1})$ and $(\mathbf{W}_{2},T_{2})$, respectively.
+By construction, these e-value vectors satisfy
+$\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[\mathbf{e}_{1,i}]\leq p,\quad\sum_{i\in\mathcal{H}_{0}}\mathbb{E}[\mathbf{e}_{2,i}]\leq p.$
+We propose combining the two vectors into a unified e-value vector $\mathbf{e}$, defined component-wise as
+$\mathbf{e}_{i}=w_{1}\mathbf{e}_{1,i}+w_{2}\mathbf{e}_{2,i},\quad\text{with }w_{1}+w_{2}\leq 1.$ (E.2)
+This ensures that the resulting e-value vector $\mathbf{e}$ also satisfies the condition given by (2).
+A natural default choice is $w_{1}=w_{2}=0.5$.
+By Proposition 1, the proposed approach controls the FDR at the desired level.
+Algorithm 3 describes the detailed implementation of this procedure.
+Algorithm 3 Hybrid Knockoff Procedure
+1:Response vector $\mathbf{Y}$, covariate matrix $\mathbf{X}$, and significance level $\alpha_{\mathrm{eBH}}$.
+2:Run the knockoff procedure with Lasso-based test statistics at target FDR level $\alpha_{\mathrm{eBH}}/2$ to obtain test statistic vector $\mathbf{W}_{1}$ and threshold $T_{1}$.
+Compute the corresponding e-values:
+$\mathbf{e}_{1,i}=\frac{n\,\mathbbm{1}\{W_{1,i}\geq T_{1}\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{W_{1,j}\leq-T_{1}\}},\quad i=1,\dots,n.$
+3:Run the knockoff procedure with random-forest-based test statistics at target FDR level $\alpha_{\mathrm{eBH}}/2$ to obtain test statistic vector $\mathbf{W}_{2}$ and threshold $T_{2}$.
+Compute the corresponding e-values:
+$\mathbf{e}_{2,i}=\frac{n\,\mathbbm{1}\{W_{2,i}\geq T_{2}\}}{1+\sum_{j=1}^{n}\mathbbm{1}\{W_{2,j}\leq-T_{2}\}},\quad i=1,\dots,n.$
+4:Compute the combined weighted average e-values according to equation (E.2).
+5:Apply the e-BH procedure to the combined e-values at the significance level $\alpha_{\mathrm{eBH}}$.
+6:Indices of the rejected hypotheses.
+E.3 Numerical Studies for the Hybrid Knockoff Method
+We adopt the simulation settings from Ren and Barber [2024].
+In the first scenario, we generate data from a Gaussian linear model.
+Specifically, we fix the target FDR level $\alpha=0.2$, the sample size at $n=1,000$ and the feature dimension at $p=800$, while varying the signal sparsity $|\mathcal{H}_{1}|\in\{40,80\}$, where $\mathcal{H}_{1}$ is the collection of alternative hypotheses.
+In this first scenario, the covariate matrix $\mathbf{X}$ is drawn from a multivariate normal distribution $\mathcal{N}(\mathbf{0},\Sigma)$, where the covariance matrix $\Sigma$ is defined by $\Sigma_{ij}=0.5^{|i-j|}$.
+The response vector $\mathbf{Y}$ is generated according to the linear model $\mathbf{Y}\sim\mathcal{N}\bigl{(}\mathbf{X}\bm{\beta},1\bigr{)}$, where the nonzero regression coefficients $\beta_{i}$ are sampled independently from $\beta_{i}\sim\mathcal{N}(\mu/\sqrt{n},1)$, with the mean parameter $\mu$ controlling the signal strength and varying within the interval $[2.5,4]$.
+Additionally, half of the nonzero coefficients $\beta_{i}$ are set to be positive, and the other half are set to be negative.
+In the second scenario, we retain the sample size, feature dimension, and signal sparsity.
+We generate the response vector $\mathbf{Y}$ according to a nonlinear model.
+For any matrix $\mathbf{X}$, let $\mathbf{X}_{l_{1}:l_{2}}$ denote the submatrix comprising columns from index $l_{1}$ to index $l_{2}$.
+Similarly, for any vector $\bm{\beta}$, let $\bm{\beta}_{l_{1}:l_{2}}$ denote the subvector containing elements from index $l_{1}$ to index $l_{2}$.
+For any function $f$, $f(\mathbf{X})$ means that $f$ is applied to each component of $\mathbf{X}$.
+We generate the data using the following nonlinear model:
+$\mathbf{Y}\sim\mathcal{N}\left(\sin(\mathbf{X}_{1:p/2})\,\bm{\beta}_{1:p/2}+(\mathbf{X}_{(p/2+1):p})^{2}\,\bm{\beta}_{(p/2+1):p},\,1\right).$
+Thus, the relationship between the response $\mathbf{Y}$ and the first $p/2$ columns of $\mathbf{X}$ involves a sine transformation, while the last $p/2$ columns of $\mathbf{X}$ are squared.
+This creates a nonlinear dependence between $\mathbf{Y}$ and $\mathbf{X}$.
+Similar to the linear scenario, the covariance matrix $\Sigma$ used to generate covariate matrix $\mathbf{X}$ is defined by $\Sigma_{ij}=0.1^{|i-j|}$, the nonzero regression coefficients $\beta_{i}$ are independently sampled from $\beta_{i}\sim\mathcal{N}(\mu/\sqrt{n},1)$, but now with the mean parameter $\mu$ varying within the interval $[5,20]$.
+Again, half of the nonzero coefficients $\beta_{i}$ are set to be positive and half negative.
+The results under both scenarios when $|\mathcal{H}_{1}|=40$ are summarized in Figure E.1.
+The case with $|\mathcal{H}_{1}|=80$ exhibits a similar pattern; we have included these results in the Figure E.2.
+Across all simulation settings, almost every method controls the false discovery rate; only the Lasso-based procedure exhibits slight FDR inflation under the linear-model scenario.
+Regarding the power, under the linear model scenario, the Lasso-based method outperforms the random-forest-based method, whereas in the nonlinear scenario, the random-forest-based method exhibits superior performance.
+Notably, our proposed hybrid approach consistently outperforms the weaker method of the two and, in several instances in the non-linear model case, surpasses both.
+Thus, the proposed method effectively integrates the advantages of both methods and demonstrates robustness across diverse model settings.
+Figure E.1: Simulation results for hybrid knockoff methods under two scenarios: Gaussian linear model (left panel) and nonlinear model (right panel).
+Figure E.2: Simulation results for hybrid knockoff methods under two scenarios: Gaussian linear model (left panel) and nonlinear model (right panel).
+Appendix F Additional Details for the Hybrid Multiple Testing Procedure
+F.1 The Way to Generate P-values
+We generate p-values as follows.
+Under the null hypothesis, we simulate the test statistics $X_{i}$ from the standard normal distribution $\mathcal{N}(0,1)$.
+Under the alternative hypothesis, $X_{i}$ follows $\mathcal{N}(\mu\log(n),\sigma^{2})$, where the parameter $\mu$ governs the signal strength.
+The p-values are then computed as $p_{i}=1-\Phi(X_{i})$, where $\Phi$ denotes the cumulative distribution function of the standard normal distribution.
+In Setting S1, we fix the sample size at $n=1,000$ and set $n_{a}=50$, corresponding to $5\%$ signals.
+The signal strength $\mu$ varies between $0.3$ and $0.5$, and $\sigma=1$.
+In Setting S2, we fix the sample size at $n=3,000$ and set $n_{a}=750$, corresponding to $25\%$ signals.
+The signal strength $\mu$ varies between $0.275$ and $0.295$, and $\sigma=0.4$.
+F.2 The Computational Cost for the Hybrid Approach
+The computational cost for Setting S1 with $\mu=0.4$ for each method are shown in Table F.1.
+method BH BC ST eBH_Ave eBH_Ada fast_eBH_Ada
+time in seconds $1\times 10^{-4}$ $2\times 10^{-4}$ $1\times 10^{-4}$ $2\times 10^{-4}$ 2.9 0.17
+Table F.1: The average running time for each method in the 500 simulations.
+eBH_Ave denotes the hybrid approach with data-independent weights.
+eBH_Ada denotes the hybrid approach with data-dependent weights, and fast_eBH_Ada is the fast version of the eBH_Ada procedure.
+Appendix G Additional Details for Structure-Adaptive Multiple Testing
+G.1 Structure-Adaptive Multiple Testing via Cross-Fitting
+Algorithm 4 summarizes the structure-adaptive multiple testing procedure via cross-fitting.
+Algorithm 4 Cross-fitting based structure adaptive multiple testing
+1:p-values $p_{1},\dots,p_{n}$, covariates $x_{1},\dots,x_{n}$, group indices $\mathcal{G}_{1},\dots,\mathcal{G}_{G}$, significance levels $\alpha_{\mathrm{FBC}}$ and $\alpha_{\mathrm{eBH}}$
+2:for $g=1,\dots,G$ do
+3:  Compute the rejection function estimate $\hat{\varphi}_{i}(\cdot)$ for $i\in\mathcal{G}_{g}$ using $\mathbf{p}_{-g}$ and $\mathbf{x}_{-g}$.
+4:  Calculate the threshold $T_{g}$ using (9).
+5:  for $i\in\mathcal{G}_{g}$ do
+6:   Calculate the e-value $e_{i}$ using (10).
+7:  end for
+8:end for
+9:Assemble the e-values from all groups.
+10:Run the e-BH procedure utilizing the assembled e-values at the level $\alpha_{\mathrm{eBH}}$.
+11:The indices of rejected hypotheses.
+Example: Data-Dependent Weights
+Below is a concrete example illustrating the form of the data-dependent weights. In the case of $G=2$, we have
+$w_{i}=\frac{\frac{n}{n_{1}}\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1}\}\right)}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{1}\}\right)+\sup_{p\in[0,1]}\sum_{j\in\mathcal{G}_{2}}\mathbbm{1}\{\hat{\varphi}_{j}^{1,i,p}(1-p_{j})\leq T_{2,j}^{1,i,p}\}}$
+for $i\in\mathcal{G}_{1}$ and
+$w_{i}=\frac{\frac{n}{n_{2}}\left(1+\sum_{j\neq i,j\in\mathcal{G}_{2}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{2}\}\right)}{\sup_{p\in[0,1]}\sum_{j\in\mathcal{G}_{1}}\mathbbm{1}\{\hat{\varphi}_{j}^{2,i,p}(1-p_{j})\leq T_{1,j}^{2,i,p}\}+\left(1+\sum_{j\neq i,j\in\mathcal{G}_{2}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{2}\}\right)}$
+for $i\in\mathcal{G}_{2}$.
+Introduction of Local FDR
+In the FBC procedure, we employ a rejection rule based on the local FDR [Sun and Cai, 2007] within the two-group mixture model framework.
+Specifically, assume that the p-value $p_{i}$ is independently generated from the two-group mixture model: $\pi_{i}f_{0}+(1-\pi_{i})f_{1,i}$, where $\pi_{i}\in(0,1)$ is the mixing proportion, and $f_{0}$ and $f_{1,i}$ represent the p-value distributions under the null and alternative hypotheses, respectively.
+The local FDR is defined as
+$\mathrm{Lfdr}_{i}(p)=\frac{\pi_{i}f_{0}(p)}{\pi_{i}f_{0}(p)+(1-\pi_{i})f_{1,i}(p)},$
+which represents the posterior probability that the $i$th hypothesis is null given the observed p-value $p$.
+The monotone likelihood ratio assumption [Sun and Cai, 2007] posits that $f_{1,i}(p)/f_{0}(p)$ is decreasing in $p$.
+Under this assumption, $\varphi_{i}(p)=\mathrm{Lfdr}_{i}(p)$ is monotonically increasing in $p$, which satisfies the conditions for the FBC procedure to control FDR at the target level [Li and Zhang, 2025].
+Furthermore, the literature demonstrates that the rejection rule $\mathbbm{1}\{\varphi_{i}(p_{i})=\mathrm{Lfdr}_{i}(p_{i})\leq t\}$ is optimal in maximizing the expected number of true positives among decision rules that control the marginal FDR at level $\alpha$ (see, e.g., Sun and Cai [2007], Lei and Fithian [2018], Cao et al.
+[2022]).
+Parameter Estimation
+In practice, we using Local FDR as the hypothesis specific rejection function.
+Set $f_{0}(p)=\mathbbm{1}\{p\in[0,1]\}$ and $f_{1,i}(p)=(1-\pi_{i})(1-\kappa_{i})p^{-\kappa_{i}}$ with $\kappa_{i}\in(0,1).$ We consider the working models that link $(\pi_{i},\kappa_{i})$ with the external covariates:
+$\pi_{i}$ $=\pi_{\beta_{\pi}}(x_{i})=\frac{1}{1+\exp(-\beta_{\pi,0}-\beta_{\pi,1}^{\top}x_{i})},$
+$\kappa_{i}$ $=\kappa_{\beta_{\kappa}}(x_{i})=\frac{1}{1+\exp(-\beta_{\kappa,0}-\beta_{\kappa,1}^{\top}x_{i})},$
+where the parameters $\beta_{\pi}=(\beta_{\pi,0},\beta_{\pi,1}^{\top})\in\mathbb{R}^{d+1}$ and $\beta_{\kappa}=(\beta_{\kappa,0},\beta_{\kappa,1}^{\top})\in\mathbb{R}^{d+1}$ can be estimated by maximizing the pseudo-log-likelihood using the EM algorithm.
+Please refer to Zhang and Chen [2022] for more optimization details.
+After obtaining the estimates $\hat{\beta}_{\pi}$ and $\hat{\beta}_{\kappa}$ from the EM algorithm, we define
+$\hat{\pi}_{i}=\begin{cases}\epsilon_{1}&\text{if }1/\bigl{(}1+\exp(-\hat{\beta}_{\pi,0}-\hat{\beta}_{\pi,1}^{\top}x_{i})\bigr{)}\leq\epsilon_{1},\\ 1/\bigl{(}1+\exp(-\hat{\beta}_{\pi,0}-\hat{\beta}_{\pi,1}^{\top}x_{i})\bigr{)}&\text{if }\epsilon_{1}<1/\bigl{(}1+\exp(-\hat{\beta}_{\pi,0}-\hat{\beta}_{\pi,1}^{\top}x_{i})\bigr{)}<1-\epsilon_{2},\\ 1-\epsilon_{2}&\text{otherwise,}\end{cases}$
+where winsorization is used to prevent $\hat{\pi}_{i}$ from being too close to zero or one to stabilize the algorithm. We define the rejection rule
+$\hat{\varphi}_{i}(p)=\frac{\hat{\pi}_{i}}{\hat{\pi}_{i}+(1-\hat{\pi}_{i})(1-\hat{\kappa}_{i})p^{-\hat{\kappa}_{i}}}\leq t,$
+where $\hat{\kappa}_{i}=1/\bigl{(}1+\exp(-\hat{\beta}_{\kappa,0}-\hat{\beta}_{\kappa,1}^{\top}x_{i})\bigr{)}$.
+We then apply the FBC procedure with the estimated $\hat{\varphi}_{i}$ at the target FDR level $\alpha$.
+The corresponding e-values are computed via (10), and the weights are obtained from (12).
+To reduce computational cost, we introduce the following, less expensive weighting scheme:
+$w_{i}=\frac{\frac{n}{n_{g}}\left(1+\sum_{j\neq i,j\in\mathcal{G}_{g}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{g}\}\right)}{\left(1+\sum_{j\neq i,j\in\mathcal{G}_{g}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{g}\}\right)+\sum_{g^{\prime}\neq g}\sum_{j\in\mathcal{G}_{g^{\prime}}}\mathbbm{1}\{\hat{\varphi}_{j}(1-p_{j})\leq T_{g^{\prime},j}\}}.$
+In practice, we fix $\epsilon_{1}=0.1$, $\epsilon_{2}=1\times 10^{-5}$, $G=2$ with $|\mathcal{G}_{1}|=|\mathcal{G}_{2}|$, and $\alpha_{\mathrm{FBC}}=\alpha_{\mathrm{eBH}}/(1+\alpha_{\mathrm{eBH}})$.
+Benchmark Methods
+  • BH: The BH procedure [Benjamini and Hochberg, 1995]. We implement this method using the p.adjust function in R.
+  • IHW_storey: The covariate-powered cross-weighted method with Storey’s procedure to estimate the null-proportion [Ignatiadis and Huber, 2021].
+We implement this method using the ihw_bh function in the R package IHWStatsPaper.
+  • IHW_betamix: The covariate-powered cross-weighted method with the beta mixture model [Ignatiadis and Huber, 2021].
+We implement this method using the ihw_betamix_censored function in the R package IHWStatsPaper.
+  • AdaPT: The adaptive p-value thresholding procedure [Lei and Fithian, 2018]. We implement this method using the adapt_glm function in the R package adaptMT.
+  • SABHA: The structure adaptive BH procedure [Li and Barber, 2019]. The code was downloaded from the link provided by the original paper.
+Comparison of Proposed Method with $\tau$-censored Weighted BH Procedure
+The $\tau$-censored weighted BH procedure proposed by Zhao and Zhou [2024] is essentially a variant of the weighted BH procedure that uses a leave-one-out technique to construct weights.
+Our method is built upon the BC procedure, and the weights in our approach are for combining the e-values from different groups (obtained through sample-splitting).
+In other words, the weights serve different goals in the two procedures.
+The exact constructions of the weights in the two procedures are also very different.
+However, the two methods do use a similar trick to construct weights.
+In Zhao and Zhou [2024], the weight involves taking an infimum over $p_{i}\in[0,1]$ in the initial weights (see Section 3.1 therein).
+In our procedure, the construction of the weights involves taking the supremum over $p_{i}\in[0,1]$, which is crucial for the proof to go through.
+G.2 Additional Numerical Results for Structure-Adaptive Multiple Testing
+The results for $a_{f}=0.5$ are presented in Figure G.2.
+When the signal is sparse and the covariate is less informative, slight FDR inflation is observed in SABHA and the two versions of IHW.
+eBH_FBC has the highest power, followed by SABHA and the two versions of IHW.
 AdaPT, on the other hand, shows a power loss when compared to the BH procedure.
-However, as the covariate becomes more informative, all structure adaptive methods outperform the BH procedure, and eBH_GBC has the most true discoveries when the signal is sparse.
-Furthermore, when the signal becomes dense, eBH_GBC, AdaPT, and SABHA have similar performance in power.
-Figure C.2 shows the results for $a_{f}=0$, i.e., the alternative p-value distribution is independent of the covariates.
-In this case, AdaPT performs the best, followed by eBG_GBC and SABHA, which dominate IHW and the BH procedure.
-Table C.1: Parameter settings for the case of $G=4$.
-Here, $n_{g}$ represents the number of hypotheses for the $g$th group; $n_{ga}$ denotes the number of non-null hypotheses in the $g$th group.
-The parameter $u_{g}$ is the upper bound of the uniform distribution for the p-values under the alternatives for the $g$th group.
-Setting Target FDR level $n_{1}$ $n_{1a}$ $p_{1}$ $n_{2}$ $n_{2a}$ $p_{2}$ $n_{3}$ $n_{3a}$ $p_{3}$ $n_{4}$ $n_{4a}$ $p_{4}$
-Setting F1 0.05 100 20 $1.5\times 10^{-3}$ 100 20 $1.5\times 10^{-3}$ 1000 20 $1.5\times 10^{-3}$ 1000 20 $1.5\times 10^{-3}$
-Setting F2 0.05 100 1 $5\times 10^{-4}$ 100 20 $2\times 10^{-3}$ 100 20 $3\times 10^{-3}$ 100 20 $4\times 10^{-3}$
-Setting F3 0.2 50 2 $1\times 10^{-3}$ 100 2 $1\times 10^{-3}$ 50 4 $2\times 10^{-3}$ 100 4 $2\times 10^{-3}$
-[…]
-BC_Com 0.698 0.697 0.698 0.698 0.698 0.042 0.006 0.006 0.072 0.073
-BC_Sep 0.579 0.906 0.893 0.245 0.273 0.064 0.048 0.045 0.025 0.03
-eBH_1 0.022 0 0 0.045 0.045 0.006 0 0 0.005 0.006
-eBH_2 0.051 0.051 0.051 0.051 0.05 0.004 0.003 0.002 0.005 0.006
-eBH_Ada 0.22 0.313 0.313 0.118 0.137 0.021 0.016 0.015 0.012 0.015
-Table C.1: Parameter settings for the case of $G=4$.
-Here, $n_{g}$ represents the number of hypotheses for the $g$th group; $n_{ga}$ denotes the number of non-null hypotheses in the $g$th group.
-The parameter $u_{g}$ is the upper bound of the uniform distribution for the p-values under the alternatives for the $g$th group.
-Table C.2: FDR and power for Setting F1, where the nominal FDR level is 5%.
-Table C.3: FDR and power for Setting F2, where the nominal FDR level is 5%.
-[…]
-BC_Com 0.93 0.947 0.94 0.932 0.919 0.044 0.312 0.032 0.03 0.031
-BC_Sep 0.788 0 0.863 0.794 0.747 0.055 0 0.044 0.043 0.041
-eBH_1 0.001 0 0.001 0.001 0.001 0 0 0 0 0
-eBH_2 0.001 0 0.001 0.001 0.001 0 0 0 0 0
-eBH_Ada 0.541 0 0.56 0.551 0.538 0.032 0 0.029 0.03 0.03
-[…]
-BC_Com 0.955 0.967 0.968 0.948 0.95 0.174 0.147 0.236 0.095 0.151
-BC_Sep 0.393 0.149 0.139 0.528 0.507 0.346 0.1 0.094 0.173 0.171
-eBH_1 0.002 0.001 0.003 0.001 0.003 0.002 0.001 0.002 0 0.002
-eBH_2 0.007 0.005 0.004 0.008 0.008 0.004 0.003 0.003 0.003 0.003
-eBH_Ada 0.038 0.012 0.006 0.053 0.051 0.021 0.008 0.004 0.016 0.018
-Table C.3: FDR and power for Setting F2, where the nominal FDR level is 5%.
-Table C.4: FDR and power for Setting F3, where the nominal FDR level is 20%.
-Figure C.1: Empirical FDR and power for $a_{f}=0.5$.
-Figure C.2: Empirical FDR and power for $a_{f}=0$.
-wzAYj8cwTZPwPH9/sVg8PXweDAauqqr2cDjEer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">
+However, as the covariate becomes more informative, all structure adaptive methods outperform the BH procedure, and eBH_FBC has the most true discoveries when the signal is sparse.
+Furthermore, when the signal becomes dense, eBH_FBC, AdaPT, and SABHA have similar performance in power.
+Figure G.3 shows the results for $a_{f}=0$, i.e., the alternative p-value distribution is independent of the covariates.
+In this case, AdaPT performs the best, followed by eBG_FBC and SABHA, which dominate IHW and the BH procedure.
+Figure G.1: Variance for empirical FDR and power when $a_{f}=1$.
+Signal sparsity is controlled by setting $a_{0}\in\{3.5,2.5,1.5\}$, giving rise to sparse, moderate, and dense alternatives, respectively.
+Covariate informativeness is tuned via $a_{1}\in\{1.5,2,2.5\}$, corresponding to weak, moderate, and strong auxiliary signals.
+Figure G.2: Empirical FDR and power for $a_{f}=0.5$.
+Signal sparsity is controlled by setting $a_{0}\in\{3.5,2.5,1.5\}$, giving rise to sparse, moderate, and dense alternatives, respectively.
+Covariate informativeness is tuned via $a_{1}\in\{1.5,2,2.5\}$, corresponding to weak, moderate, and strong auxiliary signals.
+Figure G.3: Empirical FDR and power for $a_{f}=0$.
+Signal sparsity is controlled by setting $a_{0}\in\{3.5,2.5,1.5\}$, giving rise to sparse, moderate, and dense alternatives, respectively.
+Covariate informativeness is tuned via $a_{1}\in\{1.5,2,2.5\}$, corresponding to weak, moderate, and strong auxiliary signals.
+method BH IHW_Storey IHW_betamix AdaPT SABHA eBH_FBC
+time in seconds $3\times 10^{-4}$ $2.3\times 10^{-3}$ $1.9\times 10^{-3}$ 4.24 $6\times 10^{-4}$ 0.45
+Table G.1: The average running time for each method in 100 simulations.